### v1

Abstract
Maximum likelihood identification of linear time-invariant models is a difficult problem because it is, in general, a nonlinear semidefinite program, with semidefinite covariance matrix arguments and semidefinite filter stability constraints. To enforce filter stability, we establish a general theory of closed constraints on the system eigenvalues using LMI regions. To solve the identification problem, we employ a Cholesky factorization method that reduces the semidefinite program to a standard nonlinear program. Finally, we apply the identification algorithm to a class of linear plant and disturbance models commonly used in offset-free model predictive control applications. Specifically, we consider models that are structured with uncontrollable, integrating disturbance states. We solve this disturbance modeling problem, and validate the resulting controller and estimator performance, in two real-world case studies: first, a low-cost benchmark temperature control laboratory, and second, an industrial-scale chemical reactor at Eastman Chemical’s Kingsport plant.
I Introduction
Linear system identification is an important problem in control applications and theory, with a longstanding history of applied use and a large body of literature on its theory [1, 2, 3]. In particular, stochastic linear time-invariant (LTI) state-space models are used in a variety of control contexts to represent dynamics with process and measurement uncertainty. Maximum likelihood (ML) identification is the preferable method in parametric identification for its desirable statistical properties (consistency, asymptotic efficiency) and ability to handle general parameterizations, constraints, and stochastic noise models [4, 5, 6].
The main computational challenge to ML identification of LTI models is that, in general, the problem is a nonlinear semidefinite program (SDP), with semidefinite matrix arguments and semidefinite filter stability constraints. In the ML identification literature, nonlinear SDPs are avoided either by using the expectation maximization (EM) algorithm [7, 8, 9, 10], by invoking simplified covariance matrix parameterizations (positive diagonal matrices, scaled rotation matrices) [11, 12, 13, 14], or by minimizing the determinant of the sample covariance [4, 5, 6, 15, 16, 17]. Each of these strategies impose a specific structure on the estimates. The chief advantage of the EM algorithm is that, for black-box models, there are closed-form solutions for the iterates. When further structure on the model is imposed, it may invoke an optimization problem within the EM algorithm iterates, significantly slowing down the computation [18]. Diagonal covariances are a highly constrained structure, and the type of simple covariance structures available in the literature do not scale to high-dimensional systems. Finally, the minimum determinant approach requires a fully parameterized covariance matrix for the Kalman filter innovations. Moreover, these last two approaches introduce a filter stability constraint that is not explicitly enforced in the current literature.
EM does not have strong convergence guarantees even in the best case scenario. While it can be shown that the EM iterates produce, almost surely, an increasing sequence of likelihood values [7, 9], slow convergence at low noise levels has been reported on a range of problems [18, 19, 20, 21, 22, 23], with hundreds or thousands of iterates being common for small ML problems in even the latest EM works [18]. Interior point, and even gradient methods [23], are therefore preferable to the standard EM approach.
Linear identification of nonlinear systems
In a wide variety of control applications, including chemical processes [24, 25, 26], aerospace vehicles [17, 27], combustion engines [28], nautical vehicles [11, 29], and speech recognition [8], linear approximations of the nonlinear plant are beneficial for the convenience of linear identification relative to that of nonlinear identification and the ability to meet strict computational constraints, e.g., for on-line optimal control. Linear black-box models are particularly useful when first-principles knowledge of the plant dynamics is not available.
The main difficulty of linear identification of nonlinear systems is plant-model mismatch. With ML identification, properties of the estimates are dependent on the plant’s stochastic behavior [30, 31]. For stationary, input-free models, the solution to the mismatched problem can be interpreted as (asymptotically) minimizing the Kullback-Leibler divergence between the power spectral densities of the model and plant [32]. However, there are still gaps in the treatment of inputs, state-space models, and arbitrary nonlinear plants. Rather than address this theory, we turn to control and estimation applications that specifically address plant-model mismatch.
Application to offset-free control
Model predictive control (MPC) is a widely-used advanced control method in which an optimal control problem is solved on-line, based on the current state or state estimate, and the first input in the solution trajectory is injected into the plant [33, 34]. Model quality is the main contributor to the performance of an MPC implementation [35, 36], and therefore high-quality identification algorithms are of relevance to MPC implementations.
Inherent to the use of identified plant models in MPC is the problem of plant-model mismatch. Moreover, many applications require the rejection of a stochastic, possibly nonstationary, disturbance process with unknown or un-modeled components (e.g., environmental or upstream disturbances, demand changes). In linear offset-free MPC, the stochastic LTI state-space model is augmented with uncontrollable integrating disturbance modes to achieve offset-free control in the presence of plant-model mismatch and persistent disturbances [37, 38]. We refer to these models as linear augmented disturbance models.
Linear augmented disturbance models (or the resulting observer) can either be tuned or identified. Tuning of disturbance models can be roughly divided into three categories: pole placement [39, 40, 41, 42], diagonal covariance matrix tuning [25, 43, 44], and direct filter gain tuning [45, 46, 47]. On the other hand, disturbance models have been identified only via autocovariance least squares (ALS) estimation [48] and (approximate) ML estimation [49, 50, 51]. Only our prior work in [50, 51] integrates the plant and disturbance identification in a single step. However, the method in [50, 51] uses a nested ML estimation approach in which unstructured stochastic LTI models are augmented with integrating disturbance modes. The nested ML method can be improved by consolidation to a single step ML step.
Direct data-driven control
The approach discussed so far is an indirect data-driven control design of offset-free MPC. A potential alternative is the direct data-driven control approach, where the control law is designed according to data [52, 53, 54, 55]. The drawback of this approach is its reliance on Willem’s Fundamental Lemma [56], which does not admit the required linear augmented disturbance model structure.
Contributions and outline
The main contributions of this work are (i) a method for directly solving constrained ML identification problems as NLPs, more efficiently and on a wider class of systems than the state-of-the-art, and (ii) real-world case studies of the application of this algorithm to offset-free control. The ML problem is stated in Section II. In Section III, stability and other eigenvalue constraints are formulated. Inspired by the smoothed spectral radius and abscissa formulations of [57, 58], and we present a novel barrier function theory of constraints on the system eigenvalues via the linear matrix inequality (LMI) regions of [59]. In Section IV, we present a novel modification of the Burer-Monteiro-Zhang (BMZ) method [60], reformulating a class of nonlinear SDPs as NLPs via Cholesky factorization. In identification problems, Cholesky factor subsitution has only been used for replacing semidefinite covariance matrices [49], not for general matrix inequalities. In Section V, we used the ML identification algorithm for two real-world applications of offset-free MPC: first, a benchmark temperature microcontroller [61], and second, an industrial-scale chemical reactor at Eastman Chemical’s Kingsport plant [51]. The advantage of ML identification over other disturbance modeling techniques is demonstrated. Finally, in Section VI, we conclude with a discussion of future work on applying the identification problem to other parts of the control architecture (e.g., performance monitoring, steady-state optimization, and automated MPC upkeep) and extensions to direct data-driven control.
This report is an extended version of a submitted work, and contains proofs of minor results and details on the industrial-scale reactor case study that were omitted from the journal version due to page limitations. Compared to the journal version, this report additional contains the following additions:
Notation
[…]
II Maximum likelihood identification
In this section, we formulate a maximum likelihood problem for identifying models of the following form:
$x_{k+1}$ $=A(\theta)x_{k}+B(\theta)u_{k}+w_{k}$ (1a)
$y_{k}$ $=C(\theta)x_{k}+D(\theta)u_{k}+v_{k}$ (1b)
$x_{0}$ $\sim\mathcal{N}(\hat{x}_{0}(\theta),\hat{P}_{0}(\theta))$ (1c)
[…]
where $x\in{}^{n}$ are the model states, $u\in{}^{m}$ are the inputs, $y\in{}^{p}$ are the outputs, $w\in{}^{n}$ and $v\in{}^{p}$ are the process and measurement noises, and $\mathcal{M}\coloneqq(A,B,C,D,\hat{x}_{0},\hat{P}_{0},S)$ are functions, to be defined, that map the model parameters $\theta\in\Theta$ to system matrices or vectors of appropriate dimensions. The ML estimate $\hat{\theta}_{N}$ is defined a maximizer of $p(\mathbf{y}_{N-1}|\mathbf{u}_{N-1},\theta)$, or equivalently, a solution to
$\min_{\theta\in\Theta}f_{N}(\theta)\coloneqq-\sum_{k=0}^{N-1}\ln p(y_{k}|\mathbf{u}_{k-1},\mathbf{y}_{k-1},\theta).$ (2)
The noise covariance matrix $S(\theta)$ may be partitioned as
$S(\theta)=\begin{bmatrix}Q_{w}(\theta)&S_{wv}(\theta)\\ [S_{wv}(\theta)]^{\top}&R_{v}(\theta)\end{bmatrix}$ (3)
where $Q_{w}(\theta)\in\mathbb{S}^{n}_{+}$ is the process noise covariance, $S_{wv}(\theta)$ is the cross-covariance, and $R_{v}(\theta)\in\mathbb{S}^{p}_{+}$ is the measurement noise covariance. Throughout, we impose the stronger requirement $R_{v}(\theta)\succ 0$ on the measurement noise covariance.
II-A The constraint set
The main difficulty of the ML problem eq. 2 is that the parameter constraint set $\Theta$ must necessarily contain matrix inequalities that make eq. 2 a nonlinear SDP. For example, we expect consistency (i.e., positive semidefiniteness) of the covariance matrices and nondegeneracy of the measurement noise distribution.
Assumption 1.
For all $\theta\in\Theta$, we have $\hat{P}_{0}(\theta)\succeq 0$, $S(\theta)\succeq 0$, and $R_{v}(\theta)\succ 0$.
Other matrix inequalities may arise as stability, eigenvalue, or other system-level constraints. Stability and other eigenvalue constraints are explicitly covered in Section III.
We structure the constraint set to facilitate the Cholesky factor-based reformulation in Section IV. To define this structure, we first need to define some additional notation. Consider the index sets $\mathcal{L}^{n}\coloneqq\set{(i,j)\in\mathbb{N}^{2}}{1\leq i\leq j\leq n}$ and $\mathcal{D}^{n}\coloneqq\set{(i,i)\in\mathbb{N}^{2}}$ corresponding to the sparsity patterns of $n\times n$ lower triangular and diagonal matrices, respectively. With a slight abuse of notation, we define the direct sum of index sets $\mathcal{I}\subseteq\mathcal{L}^{n}$ and $\mathcal{J}\subseteq\mathcal{L}^{m}$ by
$\mathcal{I}\oplus\mathcal{J}\coloneqq\mathcal{I}\cup\set{(i+n,j+n)}{(i,j)\in\mathcal{L}^{m}}\subseteq\mathcal{L}^{n+m}.$
For each $\mathcal{I}\subseteq\mathcal{L}^{n}$, define the sets
$\mathbb{S}^{n}[\mathcal{I}]$$\coloneqq\set{S\in\mathbb{S}^{n}}{S_{ij}=0\;\forall\;(i,j)\in\mathcal{L}^{n}\setminus\mathcal{I}}$
$\mathbb{L}^{n}[\mathcal{I}]$$\coloneqq\set{L\in\mathbb{L}^{n}}{L_{ij}=0\;\forall\;(i,j)\not\in\mathcal{I}}$
$\mathbb{L}^{n}_{++}[\mathcal{I}]$$\coloneqq\set{L\in\mathbb{L}^{n}_{++}}{L_{ij}=0\;\forall\;(i,j)\not\in\mathcal{I}}.$
Using this notation, we make the following assumptions about the structure of the parameter constraint set $\Theta$.
Assumption 2.
There exist index sets $\mathcal{D}^{n_{\Sigma}}\subseteq\mathcal{I}_{\Sigma}\subseteq\mathcal{L}^{n_{\Sigma}}$ and $\mathcal{D}^{n_{\mathcal{A}}}\subseteq\mathcal{I}_{\mathcal{A}}\subseteq\mathcal{L}^{n_{\mathcal{A}}}$ and differentiable functions $g:{}^{n_{\beta}}\times\mathbb{S}^{n_{\Sigma}}\rightarrow{}^{n_{g}}$, $h:{}^{n_{\beta}}\times\mathbb{S}^{n_{\Sigma}}\rightarrow{}^{n_{h}}$, $H:{}^{n_{\beta}}\rightarrow\mathbb{S}^{n_{\Sigma}}$, and $\mathcal{A}:{}^{n_{\beta}}\times\mathbb{S}^{n_{\Sigma}}\rightarrow\mathbb{S}^{n_{\mathcal{A}}}[\mathcal{I}_{\mathcal{A}}]$ such that
$\Theta=\{\;(\beta,\Sigma)\in{}^{n_{\beta}}\times\mathbb{S}^{n_{\Sigma}}[\mathcal{I}_{\Sigma}]\;|\;g(\beta,\Sigma)=0,\\ h(\beta,\Sigma)\leq 0,\;\Sigma\succeq H(\beta),\;\mathcal{A}(\beta,\Sigma)\succeq 0\;\}.$
Moreover, $\textnormal{cl}(\Theta_{++})=\Theta$ where
$\Theta_{++}\coloneqq\{\;(\beta,\Sigma)\in{}^{n_{\beta}}\times\mathbb{S}^{n_{\Sigma}}[\mathcal{I}_{\Sigma}]\;|\;g(\beta,\Sigma)=0,\\ h(\beta,\Sigma)\leq 0,\;\Sigma\succ H(\beta),\;\mathcal{A}(\beta,\Sigma)\succ 0\;\}.$
Under 2, the parameter set $\Theta$ contains four constraints: two standard vector equality and inequality constraints and two matrix inequalities. The purpose of the elimination algorithm is to transform the matrix inequalities into vector equalities while introducing as few new variables as possible. The first matrix inequality enforces a lower bound $H(\beta)$ on the sparse symmetric matrix argument $\Sigma$. In many cases this lower bound is simply zero or a small diagonal matrix, but we have left the bound general for illustrative purposes. This inequality is in the form used by [60], where variables in the Cholesky factorization
$\Sigma$ $=L_{\Sigma}L_{\Sigma}^{\top}+H(\beta),$ $L_{\Sigma}$ $\in\mathbb{L}^{n_{\Sigma}}_{++}$ (4)
are algorithmically eliminated to write $\Sigma$ in terms of just $H(\beta)$ and a sparse lower triangular matrix $L^{\mathcal{I}_{\Sigma}}\in\mathbb{L}^{n_{\Sigma}}_{++}[\mathcal{I}_{\Sigma}]$. Similarly, in Section IV, we use the sparsity structure in the second matrix inequality $\mathcal{A}(\beta,\Sigma)$ to algorithmically eliminate variables in the squared slack variable transformation
$\mathcal{A}(\beta,\Sigma)$ $=L_{\mathcal{A}}L_{\mathcal{A}}^{\top},$ $L_{\mathcal{A}}$ $\in\mathbb{L}^{n_{\mathcal{A}}}_{++}$ (5)
by writing $L_{\mathcal{A}}$ in terms of a sparse lower triangular matrix $L^{\mathcal{I}_{\mathcal{A}}}\in\mathbb{L}^{n_{\mathcal{A}}}_{++}[\mathcal{I}_{\mathcal{A}}]$. The second part of 2 guarantees the existence and uniqueness of these Cholesky factors $(L_{\Sigma},L_{\mathcal{A}})$, helps to avoid divisions by zero during the variable elimination procedure, and allows taking limits.
Remark 3.
2 rules out direct use of the strict inequality $R_{v}(\theta)\succ 0$. To satisfy 1, we use the closed constraint $R_{v}(\theta)\succeq\delta I_{p}$ with a small backoff $\delta>0$.
Remark 4.
The index set $\mathcal{I}_{\Sigma}$ provides enough sparsity in $\Sigma$ to reduce the problem to $n_{\beta}+|\mathcal{I}_{\Sigma}|$ variables (excluding slack terms). This is important when $\Sigma$ has a particular structure. The most common example is the block diagonal structure $\Sigma=\hat{P}_{0}\oplus Q_{w}\oplus R_{v}\in\mathbb{S}^{2n+p}[\mathcal{I}_{\Sigma}]$ where $\mathcal{I}_{\Sigma}\coloneqq\mathcal{L}^{n}\oplus\mathcal{L}^{n}\oplus\mathcal{L}^{p}$. We may further restrict $Q_{w}$ and $R_{v}$ to take block tridiagonal and diagonal structures, e.g.,
$Q_{w}$ $=\begin{bmatrix}Q_{1,1}&Q_{1,2}\\ Q_{1,2}^{\top}&Q_{2,2}&\ddots\\ &\ddots&\ddots&Q_{\tilde{n}-1,\tilde{n}}\\ &&Q_{\tilde{n}-1,\tilde{n}}^{\top}&Q_{\tilde{n},\tilde{n}}\end{bmatrix}$
$R_{v}$ $=R_{1}\oplus\ldots\oplus R_{\tilde{n}}$
that arise in sequentially interconnected processes such as chemical plants. Adding a $Q_{1,\tilde{n}}$ block can account for an overall recycle loop. Note that if we parameterize the block tridiagonal $Q_{w}$ via a sparse shaping matrix (i.e., $Q_{w}=G_{w}G_{w}^{\top}$), then there are more parameters than if the sparsity of $Q_{w}$ is known unless the rank of $G_{w}$ is known to be low.
Remark 5.
The index set $\mathcal{I}_{\mathcal{A}}$ provides sparsity on the range of $\mathcal{A}$ to eliminate all but $|\mathcal{I}_{\mathcal{A}}|$ entries of the squared slack term $L_{\mathcal{A}}$. Typically $\mathcal{A}$ is a block diagonal matrix of eigenvalue constraints, so the dense matrix $L_{\mathcal{A}}$ is reduced to a sparse matrix $L_{\mathcal{A}}=L_{\mathcal{A},1}\oplus\ldots\oplus L_{\mathcal{A},{\tilde{n}_{\mathcal{A}}}}$.
II-B Kalman filtering and the log-likelihood
In this subsection, we derive exact and approximate expressions for the negative log-likelihood. For brevity, we drop the dependence on $\theta$ where appropriate and write
$x_{k+1}$ $=Ax_{k}+Bu_{k}+w_{k}$
$y_{k}$ $=Cx_{k}+Du_{k}+v_{k}$
$x_{0}$ $\sim\mathcal{N}(\hat{x}_{0},\hat{P}_{0})$
[…]
and $\mathcal{M}=(A,B,C,D,\hat{x}_{0},\hat{P}_{0},S)$.
II-B1 Time-varying Kalman filter formulation
Consider the Kalman filter in innovations form
$\hat{x}_{k+1}$ $=A\hat{x}_{k}+Bu_{k}+\mathcal{K}_{k}e_{k}$ (6a)
$y_{k}$ $=C\hat{x}_{k}+Du_{k}+e_{k}$ (6b)
[…]
where
$\hat{P}_{k+1}$ $\coloneqq A\hat{P}_{k}A^{\top}+Q_{w}-\mathcal{K}_{k}\mathcal{R}_{k}\mathcal{K}_{k}^{\top}$ (6d)
$\mathcal{K}_{k}$ $\coloneqq(A\hat{P}_{k}C^{\top}+S_{wv})\mathcal{R}_{k}^{-1}$ (6e)
$\mathcal{R}_{k}$ $\coloneqq C\hat{P}_{k}C^{\top}+R_{v}.$ (6f)
Since the $e_{k}$ in eq. 6 are mutually independent, we have the negative log-likelihood
$f_{N}(\theta)\propto\frac{1}{2}\sum_{k=0}^{N-1}\ln\det\mathcal{R}_{k}(\theta)+|e_{k}(\theta)|_{[\mathcal{R}_{k}(\theta)]^{-1}}^{2}$
and we can write eq. 2 equivalently as
$\min_{\theta\in\Theta}\frac{1}{2}\sum_{k=0}^{N-1}\ln\det\mathcal{R}_{k}(\theta)+|e_{k}(\theta)|_{[\mathcal{R}_{k}(\theta)]^{-1}}^{2}$ (7)
where the $e_{k}(\theta)$ and $\mathcal{R}_{k}(\theta)$ are given by the recursion eq. 6. For black-box covariance models
$\mathcal{M}(\beta,\hat{P}_{0}\oplus S)=(A(\beta),B(\beta),C(\beta),D(\beta),\hat{x}_{0}(\beta),\hat{P}_{0},S)$
the covariance consistency constraint
$\Theta\coloneqq\set{(\beta,\Sigma)\in{}^{n_{\beta}}\times\mathbb{S}^{n+p}[\mathcal{L}^{n}\oplus\mathcal{L}^{n+p}]}{\Sigma\succeq\begin{bmatrix}0\\ &\varepsilon I_{p}\end{bmatrix}}$
suffices, without any vector constraints or general matrix inequalities. These constraints say nothing about stability of the filter eq. 6, so unstable filters may be realized during optimization, producing numerically infinite or undefined values.
II-B2 Steady-state Kalman filter formulation
In most situations, the state error covariance matrix converges exponentially fast to a steady-state solution $\hat{P}_{k}\rightarrow\hat{P}$, so it suffices to consider the following steady-state filter:
$\hat{x}_{k+1}$ $=A\hat{x}_{k}+Bu_{k}+Ke_{k}$ (8a)
$y_{k}$ $=C\hat{x}_{k}+Du_{k}+e_{k}$ (8b)
[…]
where $K\coloneqq(A\hat{P}C^{\top}+S_{wv})R_{e}^{-1}$, $R_{e}\coloneqq C\hat{P}C^{\top}+R_{v}$, and $\hat{P}$ is the unique, stabilizing solution to the discrete algebraic Riccati equation (DARE),
$\hat{P}=A\hat{P}A^{\top}+Q_{w}-(A\hat{P}C^{\top}+S_{wv})\\ \times(C\hat{P}C^{\top}+R_{v})^{-1}(A\hat{P}C^{\top}+S_{wv})^{\top}.$ (9)
Recall a solution to the DARE eq. 9 is stabilizing if the resulting $A_{K}\coloneqq A-KC$ is stable.
Convergence of $\hat{P}_{k}$ to $\hat{P}$ is equivalent to the solution to the DARE eq. 9 being unique and stabilizing. We generally assume such a solution exists, but for completeness, we state the following proposition, adapted from [64, Thm. 18(iii)] (see Appendix A for proof).
Proposition 6.
Assume $R_{v}\succ 0$ and consider the full rank factorization
$\begin{bmatrix}Q_{w}&S_{wv}\\ S_{wv}^{\top}&R_{v}\end{bmatrix}=\begin{bmatrix}\tilde{B}\\ \tilde{D}\end{bmatrix}\begin{bmatrix}\tilde{B}^{\top}&\tilde{D}^{\top}\end{bmatrix}$
Then the following statements are equivalent:
  1. 1.
    The DARE eq. 9 has a unique, stabilizing solution $\hat{P}\succeq 0$.
  2. 2.
    The error covariance converges exponentially fast $\hat{P}_{k}\rightarrow\hat{P}$ for any $\hat{P}_{0}\succeq 0$.
  3. 3.
    $(A,C)$ is detectable and $(A-FC,\tilde{B}-F\tilde{D})$ is stabilizable for all $F\in{}^{n\times p}$.
Under the steady-state approximation, the likelihood is
$\tilde{f}_{N}(\theta)\coloneqq\frac{N}{2}\ln\det R_{e}(\theta)+\frac{1}{2}\sum_{k=0}^{N-1}|e_{k}(\theta)|_{[R_{e}(\theta)]^{-1}}^{2}$
and we can approximate solutions to eq. 2 by solving
$\min_{\theta\in\Theta}\frac{N}{2}\ln\det R_{e}(\theta)+\frac{1}{2}\sum_{k=0}^{N-1}|e_{k}(\theta)|_{[R_{e}(\theta)]^{-1}}^{2}$ (10)
where the $e_{k}(\theta)$ are given by the recursion eq. 8 and $R_{e}(\theta)$ is found by solving the DARE eq. 9.
While $R_{e}(\theta)$ and $K(\theta)$ could be defined via $\hat{P}(\theta)$, taken as the function that returns solutions to the DARE eq. 9 and therefore enforcing filters stability, it is more convenient to directly parameterize these matrices. In fact, it is equivalent to consider the model structure
$\mathcal{M}=\left(A,B,C,D,\hat{x}_{0},0,\begin{bmatrix}KR_{e}K^{\top}&KR_{e}\\ R_{e}K^{\top}&R_{e}\end{bmatrix}\right)$
where $K:\Theta\rightarrow{}^{n\times p}$ and $R_{e}:\Theta\rightarrow\mathbb{S}^{p}_{++}$ are now given functions that are not explicitly related to the DARE eq. 9. We write this model parameterization as
$\mathcal{M}_{\textnormal{KF}}=(A,B,C,D,\hat{x}_{0},K,R_{e})$ (11)
using the subscript KF to denote that $\mathcal{M}_{\textnormal{KF}}$ represents a Kalman filter in innovation form eq. 11. Since the model structure eq. 11 no longer enforces filter stability, the following stability assumption is required.
Assumption 7.
Given the model structure eq. 11, we have $\rho(A_{K}(\theta))<1$ for all $\theta\in\Theta$.
Since $\rho$ is continuous but not differentiable, 7 cannot be directly implemented in a form satisfying 2, even with a backoff term to make the inequality nonstrict, i.e., $\rho(A_{K}(\theta))\leq 1-\delta$ where $\delta>0$. We deal with enforcing the stability constraint in Section III.
II-B3 Minimum determinant formulation
Suppose, in the model structure eq. 11, that $R_{e}$ is parameterized fully, and separately from the other terms, i.e.,
$\mathcal{M}_{\textnormal{KF}}(\beta,\tilde{\Sigma}\oplus R_{e})=(A(\beta,\tilde{\Sigma}),B(\beta,\tilde{\Sigma}),C(\beta,\tilde{\Sigma}),\\ D(\beta,\tilde{\Sigma}),\hat{x}_{0}(\beta,\tilde{\Sigma}),K(\beta,\tilde{\Sigma}),R_{e}).$
Moreover, assume $R_{e}$ is constrained separately as well, i.e.,
$\Theta=\{\;(\beta,\tilde{\Sigma}\oplus R_{e})\in{}^{n_{\beta}}\times\mathbb{S}^{n_{\Sigma}}[\mathcal{I}_{\Sigma}]\;|\;\tilde{g}(\beta,\tilde{\Sigma})=0,\\ \tilde{h}(\beta,\tilde{\Sigma})\leq 0,\;\tilde{\Sigma}\succeq\tilde{H}(\beta),\;\;R_{e}\succeq\varepsilon I_{p},\;\tilde{\mathcal{A}}(\beta,\tilde{\Sigma})\succeq 0\;\}$
for some sufficiently small $\varepsilon>0$, differentiable functions $(\tilde{g},\tilde{h},\tilde{H},\tilde{\mathcal{A}})$, and index set $\mathcal{I}_{\Sigma}\coloneqq\mathcal{I}_{\tilde{\Sigma}}\oplus\mathcal{L}^{p}$ where $\mathcal{D}^{\tilde{n}_{\Sigma}}\subseteq\mathcal{I}_{\tilde{\Sigma}}\subseteq\mathcal{L}^{\tilde{n}_{\Sigma}}$ and $n_{\tilde{\Sigma}}\coloneqq n_{\Sigma}-p$. Then we can always solve eq. 10 stagewise, first in $R_{e}$, and then in the remaining variables $(\beta,\tilde{\Sigma})$. Solving the inner problem gives the solution
$\hat{R}_{e}(\beta,\tilde{\Sigma})\coloneqq\frac{1}{N}\sum_{k=0}^{N-1}e_{k}(\beta,\tilde{\Sigma})[e_{k}(\beta,\tilde{\Sigma})]^{\top}$
where we use the fact that $e_{k}$ is only dependent on $(\beta,\tilde{\Sigma})$, and we assume $\hat{R}_{e}(\beta,\tilde{\Sigma})\succeq\varepsilon I_{p}$ for all $(\beta,\tilde{\Sigma})\in\tilde{\Theta}$ where
$\tilde{\Theta}\coloneqq\{\;(\beta,\tilde{\Sigma})\in{}^{n_{\beta}}\times\mathbb{S}^{\tilde{n}_{\Sigma}}[\mathcal{I}_{\tilde{\Sigma}}]\;|\;\tilde{g}(\beta,\tilde{\Sigma})=0,\\ \tilde{h}(\beta,\tilde{\Sigma})\leq 0,\;\tilde{\Sigma}\succeq\tilde{H}(\beta),\;\tilde{\mathcal{A}}(\beta,\tilde{\Sigma})\succeq 0\;\}.$
The outer problem can be written
$\min_{(\beta,\tilde{\Sigma})\in\tilde{\Theta}}\det\hat{R}_{e}(\beta,\tilde{\Sigma}).$ (12)
The problem eq. 12 is of relevance for avoiding the nonlinear SDP formulation of eq. 10, both in the early ML identification literature [4, 5, 6] and in recent works [15, 16, 17]. None of these works consider filter stability constraints. To the best of our knowledge, only [18] consider the ML problem eq. 7 with stability constraints, but they consider open-loop stability (i.e., $\rho(A)<1$) and use the EM algorithm. To satisfy 7, we must consider filter stability constraints (i.e., $\rho(A_{K})<1$).
Remark 8.
For real-world data, $\det\hat{R}_{e}(\beta,\tilde{\Sigma})=0$ is not attainable because that would imply some direction of $y_{k}$ were perfectly modeled. Therefore, a constant $\varepsilon>0$ exists such that the lower bound $\hat{R}_{e}(\beta,\tilde{\Sigma})\succeq\varepsilon I_{p}$ is satisfied for all $(\beta,\tilde{\Sigma})\in\tilde{\Theta}$. Moreover, we need not explicitly choose $\varepsilon>0$ to satisfy the constraint $R_{e}\succ 0$ of 1.
III Eigenvalue constraints
In this section, we describe stability and other eigenvalue constraints conforming to 2. Notably, we consider the smoothed spectral radius and abscissa formulations of [57, 58], the linear matrix inequality (LMI) region approach of [59], and a novel barrier function method that combines the two approaches. While we consider the matrix $A$ in this section, constraints can be added to any square matrix of interest, such as $A$, $A_{K}$, or some submatrix thereof.
III-A Spectral radius and abscissa bounds
Recall that, for any $Q\in\mathbb{S}^{n}_{++}$, a matrix $A$ is Schur stable if and only if there exists $P\in\mathbb{S}^{n}_{++}$ such that
$P-APA^{\top}=Q.$ (13)
Likewise, for any $Q\in\mathbb{S}^{n}_{++}$, a matrix $A$ is Hurwitz stable if and only if there eixsts $P\in\mathbb{S}^{n}_{++}$ such that
$AP+PA^{\top}=-Q.$ (14)
[…]
Lemma 9 ([57, Lem. 5.1] and [58, Lem. 2.1]).
Let $Q\in\mathbb{S}^{n}_{++}$ and $\|\cdot\|:{}^{n\times n}\rightarrow\mathbb{R}_{\geq 0}$ be a submultiplicative norm. A matrix $A\in{}^{n\times n}$ is Schur stable (Hurwitz stable) if and only if $\|P_{d}(A,Q)\|$ ($\|P_{c}(A,Q)\|$) is finite.
Inspired by this observation, [57, 58] impose an upper bound on the norm of $P$ satisfying either eq. 13 or eq. 14.
III-A1 Smoothed spectral radius
Let $W,V\in\mathbb{S}^{n}_{++}$ and consider the function $\phi_{d}(A,s)\coloneqq\textnormal{tr}(VP_{d}(A/s,W))$. In [57], the implicit function theorem is used to show the existence of a smoothed spectral radius $\rho_{\varepsilon}(A)$ satisfying
$\phi_{d}(A,\rho_{\varepsilon}(A))=\varepsilon^{-1}.$ (15)
Properties of $\rho_{\varepsilon}(A)$ are reiterated in the following theorem.
Theorem 10 ([57, Thms. 5.4, 5.6]).
There exists a function $\rho_{(\cdot)}(\cdot):\mathbb{R}_{>0}\times{}^{n\times n}\rightarrow\real$ such that, for each $A\neq 0$ and $\varepsilon>0$, $\rho_{\varepsilon}(A)$ uniquely solves eq. 15, and $\rho_{\varepsilon}(0)=0$. Moreover, $\rho_{(\cdot)}(\cdot)$ has the following properties:
  1. 1.
    $\rho_{(\cdot)}(\cdot)$ is analytic on $\mathbb{R}_{>0}\times{}^{n\times n}\setminus\set{0}$ and continuous on $\mathbb{R}_{>0}\times\set{0}$;
  2. 2.
    $\rho_{\varepsilon}(A)>\rho(A)$ for all $A\neq 0$ and $\varepsilon>0$;
  3. 3.
    $\rho_{\varepsilon}(A)\searrow\rho(A)$ as $\varepsilon\searrow 0$ for all $A\in{}^{n\times n}$;
  4. 4.
    for each $\varepsilon,s>0$, $\rho_{\varepsilon}(A)\leq s$ if and only if there exists $P\succeq 0$ such that $s^{2}P-APA^{\top}=W$ and $\textnormal{tr}(VP)\leq\varepsilon^{-1}$.
The first property of Theorem 10 establishes the smoothness of $\rho_{\varepsilon}$. The second and fourth properties of Theorem 10 let us construct constraint sets satisfying 2 and 7. Finally, the third property demonstrates that $\rho_{\varepsilon}$ is an approximation of $\rho$ in the sense that $\varepsilon$ can be made arbitrarily small to keep $\rho_{\varepsilon}(A)-\rho(A)$ arbitrarily small.
III-A2 Smoothed spectral abscissa
Let $W,V\in\mathbb{S}^{n}_{++}$ and consider the function $\phi_{c}(A,s)\coloneqq\textnormal{tr}(VP_{c}(A-sI,W))$. In [58], the implicit function theorem is used to show the existence of a smoothed spectral abscissa $\alpha_{\varepsilon}(A)$ satisfying
$\phi_{c}(A,\alpha_{\varepsilon}(A))=\varepsilon^{-1}.$ (16)
Properties of $\alpha_{\varepsilon}(A)$ are reiterated in the following theorem.
Theorem 11 ([58, Thms. 2.5, 2.6]).
There exists a function $\alpha_{(\cdot)}(\cdot):\mathbb{R}_{>0}\times{}^{n\times n}\rightarrow\real$ such that, for each $A\in{}^{n\times n}$ and $\varepsilon>0$, $\alpha_{\varepsilon}(A)$ uniquely solves eq. 15. Moreover, $\alpha_{(\cdot)}(\cdot)$ has the following properties:
  1. 1.
    $\alpha_{(\cdot)}(\cdot)$ is analytic on $\mathbb{R}_{>0}\times{}^{n\times n}$;
  2. 2.
    $\alpha_{\varepsilon}(A)>\alpha(A)$ for all $A\in{}^{n\times n}$ and $\varepsilon>0$;
  3. 3.
    $\alpha_{\varepsilon}(A)\searrow\alpha(A)$ as $\varepsilon\searrow 0$ for all $A\in{}^{n\times n}$; and
  4. 4.
    for each $\varepsilon,s>0$, $\alpha_{\varepsilon}(A)\leq s$ if and only if there exists $P\succeq 0$ such that $(A-sI)P+P(A-sI)^{\top}=-W$ and $\textnormal{tr}(VP)\leq\varepsilon^{-1}$.
Similarly to the properties of Theorem 10, the first property of Theorem 11 gives $\alpha_{\varepsilon}$ its smoothness property, the second and fourth properties allow us to construct constraints that satisfy 2, and the third property demonstrates that $\alpha_{\varepsilon}$ is an approximation of $\alpha$.
III-B $\mathcal{D}$-stability constraints
We may wish to place the eigenvalues in a specified region. In [59] the problem of placing eigenvalues in any convex, open region of the complex plane is posed in terms of solving a linear matrix inequality (LMI). These so-called “LMI regions” are defined as follows.
Definition 12.
A subset $\mathcal{D}\subseteq\mathbb{C}$ that takes the form
$\mathcal{D}=\set{z\in\mathbb{C}}{f_{\mathcal{D}}(z)\succ 0}$
is called an LMI region with the characteristic function $f_{\mathcal{D}}:\mathbb{C}\rightarrow\mathbb{C}^{m\times m}$, defined as
$f_{\mathcal{D}}(z)\coloneqq M_{0}+M_{1}z+M_{1}^{\top}\overline{z}$
and generating matrices $(M_{0},M_{1})\in\mathbb{S}^{n}\times{}^{n\times n}$.
The following lemma defines the four basic LMI regions: shifted half-planes, circles centered on the real axis, conic sections, and horizontal bands.
Lemma 13.
For each $s,x_{0}\in\real$, the subsets
$\mathcal{D}_{1}$$\coloneqq\set{z\in\mathbb{C}}{\textnormal{Re}(z)<s}$
$\mathcal{D}_{2}$$\coloneqq\set{z\in\mathbb{C}}{|z-x_{0}|<s}$
$\mathcal{D}_{3}$$\coloneqq\set{z\in\mathbb{C}}{|\textnormal{Im}(z)|<s(\textnormal{Re}(z)-x_{0})}$
$\mathcal{D}_{4}$$\coloneqq\set{z\in\mathbb{C}}{|\textnormal{Im}(z)|<s}$
are LMI regions with characteristic functions
$f_{\mathcal{D}_{1}}(z)$ $\coloneqq 2s-z-\overline{z}$
$f_{\mathcal{D}_{2}}(z)$ $\coloneqq\begin{bmatrix}s&-x_{0}\\ -x_{0}&s\end{bmatrix}+\begin{bmatrix}0&1\\ 0&0\end{bmatrix}z+\begin{bmatrix}0&0\\ 1&0\end{bmatrix}\overline{z}$
$f_{\mathcal{D}_{3}}(z)$ $\coloneqq-2sx_{0}I_{2}+\begin{bmatrix}s&1\\ -1&s\end{bmatrix}z+\begin{bmatrix}s&-1\\ 1&s\end{bmatrix}\overline{z}$
$f_{\mathcal{D}_{4}}(z)$ $\coloneqq-2sI_{2}+\begin{bmatrix}0&1\\ -1&0\end{bmatrix}z+\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}\overline{z}.$
Proof.
The first identity follows from the formula $2\textnormal{Re}(z)=z+\overline{z}$. For the second identity, we have $f_{\mathcal{D}_{2}}(z)=\begin{bmatrix}s&z-x_{0}\\ \overline{z}-x_{0}&s\end{bmatrix}\succ 0$ if and only if $s>0$ and $s^{2}>|z-x_{0}|^{2}$, or equivalently, $|z-x_{0}|<s$. For the third identity, we have $f_{\mathcal{D}_{3}}(z)=\begin{bmatrix}2s(\textnormal{Re}(z)-x_{0})&2\iota\textnormal{Im}(z)\\ -2\iota\textnormal{Im}(z)&2s(\textnormal{Re}(z)-x_{0})\end{bmatrix}\succ 0$ if and only if $02s(\textnormal{Re}(z)-x_{0})>$ 0 and $4s^{2}(\textnormal{Re}(z)-x_{0})^{2}>4|\textnormal{Im}(z)|^{2}$, or equivalently, $|\textnormal{Im}(z)|<s(\textnormal{Re}(z)-x_{0})$. For the fourth identity, we have $f_{\mathcal{D}_{4}}(z)=\begin{bmatrix}2s&2\iota\textnormal{Im}(z)\\ -2\iota\textnormal{Im}(z)&2s\end{bmatrix}\succ 0$ if and only if $02s>$ 0 and $4s^{2}>4|\textnormal{Im}(z)|^{2}$, or equivalently, $|\textnormal{Im}(z)|<s$. ∎
Remark 14.
For continuous-time systems, $\mathcal{D}_{1}$ corresponds to a minimum decay rate of $s>0$, $\mathcal{D}_{3}$ corresponds to a minimum damping ratio $-\cos(\theta)$, and $\mathcal{D}_{2}\cap\mathcal{D}_{3}$ implies to a maximum natural frequency $r\sin(\theta)$, where $\theta=\tan^{-1}(s)$ and $s<0$ [59]. For discrete-time systems, $\mathcal{D}_{2}$ corresponds to a minimum decay rate of $-\ln r$, and $\mathcal{D}_{2}\cap\mathcal{D}_{3}$ implies a minimum damping ratio $-\cos(\tan^{-1}(\theta/\ln r))$ and maximum natural frequency $(\ln(r)^{2}+\theta^{2})/\Delta$, where $\theta=\tan^{-1}(s)$, $s>0$, and $\Delta$ is the sample time.
Remark 15.
An LMI region $\mathcal{D}$ is convex, open, and symmetric about the imaginary axis. The intersection of two LMI regions $\mathcal{D}\coloneqq\mathcal{D}_{1}\cap\mathcal{D}_{2}$ is an LMI region with the characteristic function $f_{\mathcal{D}}(z)=f_{\mathcal{D}_{1}}(z)\oplus f_{\mathcal{D}_{2}}(z)$. An LMI region $\mathcal{D}$ with characteristic function $f_{\mathcal{D}}$ also has characteristic function $Mf_{\mathcal{D}}(\cdot)M^{\top}$ for any nonsingular $M\in{}^{m\times m}$. By this property, we can construct any convex polyhedron that is symmetric about the real axis by intersecting left and right half-planes, horizontal strips, and conic sections. Moreover, since any convex region can be approximated, to any desired accuracy, by a convex polyhedron, the set of LMI regions is dense in the space of convex subsets of $\mathbb{C}$ that are symmetric about the real axis. For an in-depth discussion of LMI region geometry and other properties, see [65].
Throughout, assume the LMI region $\mathcal{D}$ is nonempty, not equal to $\mathbb{C}$, and its characteristic function $f_{\mathcal{D}}$ and generating matrices $(M_{0},M_{1})$ are fixed. We seek LMI conditions under which the eigenvalues of $A\in{}^{n\times n}$ lie in $\mathcal{D}$ or $\textnormal{cl}(\mathcal{D})$.
III-B1 $\mathcal{D}$-stability
In the following definition, we generalize the notion of asymptotic stability to include pole placement within a given LMI region $\mathcal{D}$.
Definition 16.
We say the matrix $A\in{}^{n\times n}$ is $\mathcal{D}$-stable if $\lambda(A)\subset\mathcal{D}$.
In [59], $\mathcal{D}$-stability of a matrix $A\in{}^{n\times n}$ is shown to be equivalent to the strict feasibility of the following system of matrix inequalities:
$M_{\mathcal{D}}(A,P)$ $\succ 0,$ $P$ $\succ 0$ (17)
where $M_{\mathcal{D}}:{}^{n\times n}\times\mathbb{S}^{n}\rightarrow\mathbb{S}^{nm}$ is defined by
$M_{\mathcal{D}}(A,P)\coloneqq M_{0}\otimes P+M_{1}\otimes(AP)+M_{1}^{\top}\otimes(AP)^{\top}.$ (18)
This fact is restated in the following theorem.
Theorem 17 ([59, Thm. 2.2]).
The matrix $A\in{}^{n\times n}$ is $\mathcal{D}$-stable if and only if eq. 17 holds for some $P\in\mathbb{S}^{n}$.
III-B2 Marginal $\mathcal{D}$-stability
The drawback of Theorem 17 is strictness of the matrix inequalities eq. 17, making them inadmissible in constraint sets satisfying 2. Suppose we relax the first inequality,
$M_{\mathcal{D}}(A,P)$ $\succeq 0,$ $P$ $\succ 0.$ (19)
Since $M_{\mathcal{D}}(A,P)$ is linear in $P$, feasibility of eq. 19 is equivalent to feasibility of
$M_{\mathcal{D}}(A,P)$ $\succeq 0,$ $P$ $\succeq P_{0}$ (20)
for some fixed $P_{0}\in\mathbb{S}^{n}_{++}$. [^1] For any $P_{0}\succ 0$ and $P$ satisfying eq. 19, define the scaling factor $\gamma\coloneqq\|P_{0}\|_{2}\|P^{-1}\|_{2}$ and a rescaled solution $P^{*}\coloneqq\gamma P$. Then $P^{*}\succeq P_{0}$ and $M_{\mathcal{D}}(A,P^{*})=\gamma M_{\mathcal{D}}(A,P)\succeq 0$. Therefore the system eq. 19 is admissible to constraint sets satisfying 2 with minor alteration.
An attempt was made in [62, Thm. 1] to characterize the class of matrices $A\in{}^{n\times n}$ for which eq. 19, but this theorem does not correctly treat eigenvalues on the boundary $\partial\mathcal{D}$. We restate [62, Thm. 1] below as a conjecture and disprove it with a simple counterexample.
Conjecture 18 ([62, Thm. 1]).
The matrix $A\in{}^{n\times n}$ satisfies $\lambda(A)\subset\textnormal{cl}(\mathcal{D})$ if and only if eq. 19 holds for some $P\in\mathbb{S}^{n}$.
Counterexample.
Let $\mathcal{D}$ be the left half-plane, consider the Jordan block $A=\begin{bmatrix}0&1\\ 0&0\end{bmatrix}$, and suppose $P=\begin{bmatrix}p_{11}&p_{12}\\ p_{12}&p_{22}\end{bmatrix}\in\mathbb{S}^{2}$ such that eq. 19 holds. Then $\lambda(A)\subset\textnormal{cl}(\mathcal{D})$ and
$0\preceq M_{\mathcal{D}}(A,P)=-\begin{bmatrix}2p_{12}&p_{22}\\ p_{22}&0\end{bmatrix}$
which implies $p_{12}=p_{22}=0$, a contradiction of eq. 19. ∎
The correction to 18 requires a more careful treatment of eigenvalues lying on the the LMI region’s boundary $\partial\mathcal{D}$. Recall that a matrix $A$ is marginally stable (in continuous- or discrete-time) if all its eigenvalues lie in the closure of the stability region (left half-plane or unit disc) and the non-simple eigenvalues lie strictly in the interior of that region. Replacing the stability region with $\mathcal{D}$, we propose the following definition of marginal $\mathcal{D}$-stability.
Definition 19.
We say a matrix $A\in{}^{n\times n}$ is marginally $\mathcal{D}$-stable if $\lambda(A)\subseteq\textnormal{cl}(\mathcal{D})$ and $\lambda\in\mathcal{D}$ for all non-simple eigenvalues $\lambda\in\lambda(A)$.
In the following proposition, we show $\mathcal{D}$-stability of $A\in{}^{n\times n}$ is equivalent to feasibility of eq. 19 (see Appendix B for proof).
Proposition 20.
The matrix $A\in{}^{n\times n}$ is marginally $\mathcal{D}$-stable if and only if eq. 19 holds for some $P\in\mathbb{S}^{n}$.
III-B3 $\mathcal{D}$-stability barrier functions
Thus far, we have proposed LMI equivalences without regard to the topology of the set of $\mathcal{D}$-stable and marginally $\mathcal{D}$-stable matrices
$\mathbb{A}_{\mathcal{D}}^{n}$$\coloneqq\set{A\in{}^{n\times n}}{\textnormal{$A$ is $\mathcal{D}$-stable}}$
$\tilde{\mathbb{A}}_{\mathcal{D}}^{n}$$\coloneqq\set{A\in{}^{n\times n}}{\textnormal{$A$ is marginally $\mathcal{D}$-stable}}.$
The following proposition characterizes the topology of $\mathbb{A}_{\mathcal{D}}^{n}$ and $\mathbb{A}_{\mathcal{D}}^{n}$ (see Appendix C for proof).
Proposition 21.
  1. (a)
    $\mathbb{A}_{\mathcal{D}}^{n}$ is open.
  2. (b)
    $\tilde{\mathbb{A}}_{\mathcal{D}}^{n}$ is not open if (i) $n\geq 2$ or (ii) $\partial\mathcal{D}\cap\real$ is nonempty.
  3. (c)
    $\tilde{\mathbb{A}}_{\mathcal{D}}^{n}$ is not closed if (i) $n\geq 4$ or (ii) $\partial\mathcal{D}\cap\real$ is nonempty and $n\geq 2$.
  4. (d)
    $\textnormal{cl}(\mathbb{A}_{\mathcal{D}}^{n})=\set{A\in{}^{n\times n}}{\lambda(A)\subset\textnormal{cl}(\mathcal{D})}$.
Proposition 21 reveals a weakness of the marginal $\mathcal{D}$-stability constraints eqs. 19 and 20. Since $\tilde{\mathbb{A}}_{\mathcal{D}}^{n}$ is not closed and $P\in\mathbb{S}^{n}_{+}$ is neither bounded nor regularized, the optimizer may fail to converge in $P$ as it seeks an $A$ that is not a limit point in $\tilde{\mathbb{A}}_{\mathcal{D}}^{n}$. While the optimizer can approach $A$ in this case, $P$ grows unbounded along the path of iterates, and the optimizer itself does not converge.
Motivated by the smoothed spectral radius and abscissa, we upper bound $\textnormal{tr}(VP)$ for some $V\in\mathbb{S}^{n}_{++}$. This still allows rescaling of $P$, however, so we also lower bound $M_{\mathcal{D}}(A,P)$,
$M_{\mathcal{D}}(A,P)$ $\succeq M,$ $P$ $\succeq 0,$ $\textnormal{tr}(VP)$ $\leq\varepsilon^{-1}$ (21)
and assume $M\in\mathbb{S}^{nm}_{+}$ is chosen in a way that implies eq. 17. Consider the parameterized linear SDP,
$\phi_{\mathcal{D}}(A)\coloneqq\inf_{P\in\mathbb{S}^{n}_{+}}\textnormal{tr}(VP)\;\textnormal{subject to}\;M_{\mathcal{D}}(A,P)\succeq M.$ (22)
The optimal value function ${}\phi_{\mathcal{D}}:{}^{n\times n}\rightarrow\mathbb{R}_{\geq 0}\cup\set{\infty}$ is a barrier function for the constraint $A\in\mathbb{A}_{\mathcal{D}}^{n}$. Proposition 22 establishes properties of $\phi_{\mathcal{D}}$ and its $\varepsilon^{-1}$-sublevel sets (see Appendix D for proof).
Proposition 22.
Let $V\in\mathbb{S}^{n}_{++}$ and $M\in\mathbb{S}^{n}_{+}$ such that $M_{\mathcal{D}}(A,P)\succeq M$ implies $M_{\mathcal{D}}(A,P)\succ 0$. Then
  1. (a)
    $\phi_{\mathcal{D}}$ is continuous on $\mathbb{A}_{\mathcal{D}}$;
  2. (b)
    for each $\varepsilon>0$, the $\varepsilon^{-1}$-sublevel set of $\phi_{\mathcal{D}}$,
    $\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon)\coloneqq$$\set{A\in{}^{n\times n}}{\phi_{\mathcal{D}}(A)\leq\varepsilon^{-1}}$ (23)
[…]
    is closed; and
  3. (c)
    $\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon)\nearrow\mathbb{A}_{\mathcal{D}}^{n}$ as $\varepsilon\searrow 0$.
Remark 23.
It is sufficient, but not necessary, to choose a positive definite lower bound $M\succ 0$.
Remark 24.
To reconstruct the $s$-sublevel sets of the smoothed spectral radius via Proposition 22, we set $M=sW\oplus 0_{n\times n}$ for any $W,V\succ 0$ and $s>0$ and apply the Schur complement lemma to $M_{\mathcal{D}_{2}}(A,P)/s-M/s$, where $\mathcal{D}_{2}$ is the circle defined in Lemma 13 with $x_{0}=0$, and $M_{\mathcal{D}_{2}}$ is defined by the generating matrices used in Lemma 13. Then the $\varepsilon^{-1}$-sublevel set of $\phi_{\mathcal{D}_{2}}$ equals the $s$-sublevel set of $\rho_{\varepsilon}$.
Remark 25.
The $\varepsilon^{-1}$-sublevel sets of $\phi_{\mathcal{D}_{1}}$ equal the $s$-sublevel sets of $\alpha_{\varepsilon}$, where $\mathcal{D}_{1}$ is the shifted half-plane defined in Lemma 13, and $M=W$ for any $W,V\succ 0$.
III-C Discussion
We conclude this section with a discussion on implementing eigenvalue constraints. Discrete LTI models with eigenvalues having positive real parts have a one-to-one correspondence with continuous LTI models [66]. As such, it is important to satisfy not only the filter stability constraint $\rho(A_{K})<1$ (and, if desired, open-loop stability constraint $\rho(A)<1$), but it is also desirable to satisfy differentiability constraints $\alpha(-A)<0$ and $\alpha(-A_{K})<0$. In practice, we find $\rho_{\varepsilon}(A_{K})\leq 1-\delta$ and $\alpha_{\varepsilon}(-A_{K})\leq 0$ to be important constraints in ML identification of linear augmented disturbance models. Without the former constraint, the optimizer will frequently pick unstable filters that evaluate to infinite objective values, and without the latter constraint, the optimal filter equations degenerate to aphysical solutions that incorrectly estimate $u\mapsto y$ gain matrices. Last, we refer the reader to [67] for other examples of matrix inequalities (e.g., detectability, minimum phase) that may be useful in system identification.
IV Cholesky reparameterizations
In this section, we seek to transform the ML problems eqs. 2, 7 and 10 from nonlinear SDPs to standard NLPs, while introducing as few new variables as possible. The main idea is to combine a squared slack variable substitution eqs. 4 and 5 with an elimination scheme that adds only $|\mathcal{I}_{\mathcal{A}}|$ variables to the optimization problem.
For this section, we define the following notation. For each $\mathcal{I}\subseteq\mathcal{L}^{n}$, let $\pi_{\mathcal{I}}^{L}:{}^{n\times n}\rightarrow\mathbb{L}^{n}[\mathcal{I}]$ and $\pi_{\mathcal{I}}:{}^{n\times n}\rightarrow\mathbb{S}^{n}[\mathcal{I}]$ denote the orthogonal projections (in the Frobenius norm) from n×n onto the subspaces $\mathbb{L}^{n}[\mathcal{I}]$ and $\mathbb{S}^{n}[\mathcal{I}]$, respectively. Let $\textnormal{chol}:\mathbb{S}^{n}_{++}\rightarrow\mathbb{L}^{n}_{++}$ denote the invertible function that maps a positive definite matrix to its Cholesky factor.
IV-A Burer-Monteiro-Zhang method
In the Burer-Monteiro-Zhang (BMZ) method [60], an invertible map is constructed between the interior of
$\Theta_{\textnormal{BMZ}}\coloneqq\set{(\beta,\Sigma)\in{}^{n_{\beta}}\times\mathbb{S}^{n_{\Sigma}}[\mathcal{I}_{\Sigma}]}{\Sigma\succeq H(\beta)}$ (25)
and
$\Theta_{\textnormal{Chol}}\coloneqq{}^{n_{\beta}}\times\mathbb{L}^{n_{\Sigma}}_{++}[\mathcal{I}_{\Sigma}]$ (26)
where $H:{}^{n_{\beta}}\rightarrow\mathbb{S}^{n_{\Sigma}}$ and $\mathcal{I}_{\Sigma}$ is some index set satisfying $\mathcal{D}^{n_{\Sigma}}\subseteq\mathcal{I}_{\Sigma}\subseteq\mathcal{L}^{n_{\Sigma}}$. The set $\Theta_{\textnormal{BMZ}}$ satisfies 2 so long as $H$ is differentiable since
$\textnormal{int}(\Theta_{\textnormal{BMZ}})=\set{(\beta,\Sigma)\in{}^{n_{\beta}}\times\mathbb{S}^{n_{\Sigma}}[\mathcal{I}_{\Sigma}]}{\Sigma\succ H(\beta)}.$
Recall $\Sigma\succ H(\beta)$ if and only if $\Sigma=L_{\Sigma}L_{\Sigma}^{\top}+H(\beta)$ for some $L_{\Sigma}\in\mathbb{L}^{n_{\Sigma}}_{++}$. With $\mathcal{J}_{\Sigma}\coloneqq\mathcal{L}^{n_{\Sigma}}\setminus\mathcal{I}_{\Sigma}$, we split $L_{\Sigma}$ into the sum of $L^{\mathcal{I}_{\Sigma}}\in\mathbb{L}^{n_{\Sigma}}_{++}[\mathcal{I}_{\Sigma}]$ and $L^{\mathcal{J}_{\Sigma}}\in\mathbb{L}^{n_{\Sigma}}[\mathcal{J}_{\Sigma}]$,
$\Sigma=(L^{\mathcal{I}_{\Sigma}}+L^{\mathcal{J}_{\Sigma}})(L^{\mathcal{I}_{\Sigma}}+L^{\mathcal{J}_{\Sigma}})^{\top}+H.$ (27)
But $\Sigma\in\mathbb{S}^{n_{\Sigma}}[\mathcal{I}_{\Sigma}]$, so we can take
$\pi_{\mathcal{J}_{\Sigma}}[(L^{\mathcal{I}_{\Sigma}}+L^{\mathcal{J}_{\Sigma}})(L^{\mathcal{I}_{\Sigma}}+L^{\mathcal{J}_{\Sigma}})^{\top}+H]=0$ (28)
to produce $|\mathcal{J}_{\Sigma}|$ equations to eliminate the $|\mathcal{J}_{\Sigma}|$ variables of $L^{\mathcal{J}_{\Sigma}}$. For each $(H,L^{\mathcal{I}_{\Sigma}})\in\mathbb{S}^{n_{\Sigma}}\times\mathbb{L}^{n_{\Sigma}}[\mathcal{I}_{\Sigma}]$, define
$L^{\mathcal{J}_{\Sigma}}_{ij}\coloneqq-\frac{1}{L_{jj}^{\mathcal{I}_{\Sigma}}}\left(H_{ij}+\sum_{k=1}^{j-1}(L^{\mathcal{I}_{\Sigma}}_{ik}+L^{\mathcal{J}_{\Sigma}}_{ik})(L^{\mathcal{I}_{\Sigma}}_{jk}+L^{\mathcal{J}_{\Sigma}}_{jk})\right)$ (29)
for each $(i,j)\in\mathcal{J}_{\Sigma}$, in a top-to-bottom and left-to-right order. So long as we never divide by zero, each $L^{\mathcal{J}_{\Sigma}}$ is fully defined by $H$ and $L^{\mathcal{I}_{\Sigma}}$ via eq. 29, and we have the following lemma.
Lemma 26 ([60, Lem. 1]).
For each $(\beta,L^{\mathcal{I}_{\Sigma}})\in{}^{n_{\beta}}\times\mathbb{L}^{n_{\Sigma}}[\mathcal{I}_{\Sigma}]$ such that $L^{\mathcal{I}_{\Sigma}}_{ii}\neq 0$ for each $i\in\mathbb{I}_{1:n_{\Sigma}}$, there is a unique $L^{\mathcal{J}_{\Sigma}}\in\mathbb{L}^{n_{\Sigma}}[\mathcal{J}_{\Sigma}]$ satisfying eq. 28.
Let $L^{\mathcal{J}_{\Sigma}}:{}^{n_{\beta}}\times\mathbb{L}^{n_{\Sigma}}_{++}[\mathcal{I}_{\Sigma}]\rightarrow\mathbb{L}^{n_{\Sigma}}[\mathcal{J}_{\Sigma}]$ be the function that maps each $(\beta,L^{\mathcal{I}_{\Sigma}})\in\Theta_{\textnormal{Chol}}$ to the $L^{\mathcal{J}_{\Sigma}}\in\mathbb{L}^{n_{\Sigma}}[\mathcal{J}_{\Sigma}]$ defined by eq. 29 with $H=H(\beta)$. Similarly, let $\Sigma:{}^{n_{\beta}}\times\mathbb{L}^{n_{\Sigma}}_{++}[\mathcal{I}_{\Sigma}]\rightarrow\mathbb{S}^{n_{\Sigma}}[\mathcal{I}_{\Sigma}]$ be the map defined by eq. 27 for each $(\beta,L^{\mathcal{I}_{\Sigma}})\in\Theta_{\textnormal{Chol}}$ with $H=H(\beta)$ and $L^{\mathcal{J}_{\Sigma}}=L^{\mathcal{J}_{\Sigma}}(\beta,L^{\mathcal{I}_{\Sigma}})$. Finally, we let
$\Psi(\beta,L^{\mathcal{I}_{\Sigma}})\coloneqq\left(\beta,\Sigma(\beta,L^{\mathcal{I}_{\Sigma}})\right)$ (30)
which has the inverse
$\Psi^{-1}(\beta,\Sigma)\coloneqq\left(\beta,\pi_{\mathcal{I}_{\Sigma}}^{L}[\textnormal{chol}(\Sigma-H(\beta))]\right)$ (31)
and we have the following lemma.
Lemma 27 ([60, Lem. 2]).
The function $\Psi$ defined by eq. 30 is a bijection between $\Theta_{\textnormal{Chol}}$ and $\textnormal{int}(\Theta_{\textnormal{BMZ}})$.
Differentiability of $\Psi$ and $\Psi^{-1}$ follow from differentiability of $H$ and the algorithm eq. 29. In fact, these functions are as smooth as $H$ is, so if $H$ is analytic, so are $\Psi$ and $\Psi^{-1}$. More importantly, the bijection $\Psi$ allows us to transform the minimum of a continuous function in $\Theta_{\textnormal{BMZ}}$ to an infimum of a function in $\Theta_{\textnormal{Chol}}$, given by the following theorem.
Theorem 28 ([60, Thm. 1]).
For any continuous function $f:{}^{n_{\beta}}\times\mathbb{S}^{n_{\Sigma}}\rightarrow\real$ that attains a minimum in $\Theta_{\textnormal{BMZ}}$,
$\min_{(\beta,\Sigma)\in\Theta_{\textnormal{BMZ}}}f(\beta,\Sigma)=\inf_{(\beta,L^{\mathcal{I}_{\Sigma}})\in\Theta_{\textnormal{Chol}}}f_{\Psi}(\beta,L^{\mathcal{I}_{\Sigma}})$ (32)
where $f_{\Psi}\coloneqq f\circ\Psi$.
We reiterate the proof of Theorem 28 for illustrative purposes.
Proof.
Continuity of $f$ implies its minimum over $\Theta_{\textnormal{BMZ}}$ equals its infimum over $\textnormal{int}(\Theta_{\textnormal{BMZ}})$, i.e.,
$\min_{(\beta,\Sigma)\in\Theta_{\textnormal{BMZ}}}f(\beta,\Sigma)=\inf_{(\beta,\Sigma)\in\textnormal{int}(\Theta_{\textnormal{BMZ}})}f(\beta,\Sigma)$
Since $\Psi$ is a bijection, we can transform the optimization variables as follows:
$\inf_{(\beta,\Sigma)\in\textnormal{int}(\Theta_{\textnormal{BMZ}})}f(\beta,\Sigma)=\inf_{(\beta,L^{\mathcal{I}_{\Sigma}})\in\Psi^{-1}(\textnormal{int}(\Theta_{\textnormal{BMZ}}))}f(\Psi(\beta,L^{\mathcal{I}_{\Sigma}}))$
and $\Theta_{\textnormal{Chol}}=\Psi^{-1}(\textnormal{int}(\Theta_{\textnormal{BMZ}}))$ and $f_{\Psi}=f\circ\Psi$ imply eq. 32. ∎
IV-B Modified Burer-Monteiro-Zhang method
For constraint sets $\Theta$ satisfying 2, the BMZ method only suffices to eliminate the variables $\Sigma$ and the matrix inequality $\Sigma\succeq H(\beta)$. The addition of the general matrix inequality $\mathcal{A}(\beta,\Sigma)\succeq 0$ requires a similar procedure to introduce squared slack variables, although the elimination procedure will no longer eliminate as many variables as are introduced. We further complicate the generalization of Lemma 27 and Theorem 28 by requiring general vector equality and inequality constraints $g(\beta,\Sigma)=0$ and $h(\beta,\Sigma)\leq 0$.
Lemma 27 gives the invertible map $\Psi:\Theta_{\textnormal{Chol}}\rightarrow\Theta_{\textnormal{BMZ}}$ with which we can define
$g_{\Psi}$ $\coloneqq g\circ\Psi,$ $h_{\Psi}$ $\coloneqq h\circ\Psi,$ $\mathcal{A}_{\Psi}$ $\coloneqq\mathcal{A}\circ\Psi.$
In this notation, consider the set
$\Theta_{\textnormal{NLP}}\coloneqq\{\,(\beta,L^{\mathcal{I}_{\Sigma}})\in{}^{n_{\beta}}\times\mathbb{L}^{n_{\Sigma}}_{++}[\mathcal{I}_{\Sigma}]\;|\;g_{\Psi}(\beta,L^{\mathcal{I}_{\Sigma}})=0,\\ h_{\Psi}(\beta,L^{\mathcal{I}_{\Sigma}})=0,\;\mathcal{A}_{\Psi}(\beta,L^{\mathcal{I}_{\Sigma}})\succ 0\,\}.$ (33)
Recall $\Sigma(\beta,L^{\mathcal{I}_{\Sigma}})\succ H(\beta)$ holds automatically for all $(\beta,L^{\mathcal{I}_{\Sigma}})\in\Theta_{\textnormal{Chol}}\coloneqq{}^{n_{\beta}}\times\mathbb{L}^{n_{\Sigma}}_{++}[\mathcal{I}_{\Sigma}]$. Then by a simple change of variables we have $\Psi^{-1}(\Theta_{++})=\Theta_{\textnormal{NLP}}$. Since $\Theta_{\textnormal{NLP}}$ and $\Theta_{++}$ are subsets of $\Theta_{\textnormal{Chol}}$ and $\textnormal{int}(\Theta_{\textnormal{BMZ}})$, respectively, we have proven the following lemma.
Lemma 29.
The function $\Psi$ defined by eq. 30 is a bijection between $\Theta_{\textnormal{NLP}}$ and $\Theta_{++}$.
Next, we remove the matrix inequality $\mathcal{A}_{\Psi}(\beta,L^{\mathcal{I}_{\Sigma}})\succ 0$ from our representation of the constraint set $\Theta_{\textnormal{NLP}}$. Recall $\mathcal{A}_{\Psi}(\beta,L^{\mathcal{I}_{\Sigma}})\succ 0$ if and only if $\mathcal{A}_{\Psi}(\beta,L^{\mathcal{I}_{\Sigma}})=L_{\mathcal{A}}L_{\mathcal{A}}^{\top}$ for some $L_{\mathcal{A}}\in\mathbb{L}^{n_{\mathcal{A}}}_{++}$. With $\mathcal{J}_{\mathcal{A}}\coloneqq\mathcal{L}^{n_{\mathcal{A}}}\setminus\mathcal{I}_{\mathcal{A}}$, we split $L_{\mathcal{A}}$ into the sum of $L^{\mathcal{I}_{\mathcal{A}}}\in\mathbb{L}^{n_{\mathcal{A}}}_{++}[\mathcal{I}_{\mathcal{A}}]$ and $L^{\mathcal{J}_{\mathcal{A}}}\in\mathbb{L}^{n_{\mathcal{A}}}[\mathcal{J}_{\mathcal{A}}]$,
$\mathcal{A}_{\Psi}(\beta,L^{\mathcal{I}_{\Sigma}})=(L^{\mathcal{I}_{\mathcal{A}}}+L^{\mathcal{J}_{\mathcal{A}}})(L^{\mathcal{I}_{\mathcal{A}}}+L^{\mathcal{J}_{\mathcal{A}}})^{\top}.$ (34)
But $\mathcal{A}_{\Psi}(\beta,L^{\mathcal{I}_{\Sigma}})\in\mathbb{S}^{n_{\mathcal{A}}}[\mathcal{I}_{\mathcal{A}}]$, so we can take
$\pi_{\mathcal{J}_{\mathcal{A}}}[(L^{\mathcal{I}_{\mathcal{A}}}+L^{\mathcal{J}_{\mathcal{A}}})(L^{\mathcal{I}_{\mathcal{A}}}+L^{\mathcal{J}_{\mathcal{A}}})^{\top}]=0$
to produce $|\mathcal{J}_{\mathcal{A}}|$ equalities with which to eliminate the $|\mathcal{J}_{\mathcal{A}}|$ free variables in $L^{\mathcal{J}_{\mathcal{A}}}$. For each $L^{\mathcal{I}_{\mathcal{A}}}\in\mathbb{L}^{n_{\mathcal{A}}}_{++}[\mathcal{I}_{\mathcal{A}}]$, we define
$L^{\mathcal{J}_{\mathcal{A}}}_{ij}\coloneqq-\frac{1}{L_{jj}^{\mathcal{I}_{\mathcal{A}}}}\sum_{k=1}^{j-1}(L^{\mathcal{I}_{\mathcal{A}}}_{ik}+L^{\mathcal{J}_{\mathcal{A}}}_{ik})(L^{\mathcal{I}_{\mathcal{A}}}_{jk}+L^{\mathcal{J}_{\mathcal{A}}}_{jk})$ (35)
for each $(i,j)\in\mathcal{J}_{\mathcal{A}}$, in a top-to-bottom and left-to-right order. Each $L^{\mathcal{J}_{\mathcal{A}}}$ satisfying eq. 34 is thus fully defined by $L^{\mathcal{I}_{\mathcal{A}}}$ via eq. 35, and we have the following corollary of Lemma 26.
Corollary 30.
For each $L^{\mathcal{I}_{\mathcal{A}}}\in\mathbb{L}^{n_{\mathcal{A}}}[\mathcal{I}_{\mathcal{A}}]$ such that $L^{\mathcal{I}_{\mathcal{A}}}_{ii}\neq 0$ for each $i\in\mathbb{I}_{1:n_{\mathcal{A}}}$, there exists a unique $L^{\mathcal{J}_{\mathcal{A}}}\in\mathbb{L}^{n_{\mathcal{A}}}[\mathcal{J}_{\mathcal{A}}]$ satisfying eq. 34.
Let $L^{\mathcal{J}_{\mathcal{A}}}:\mathbb{L}^{n_{\mathcal{A}}}_{++}\rightarrow\mathbb{L}^{n_{\mathcal{A}}}[\mathcal{I}_{\mathcal{A}}]$ denote the map described by Corollary 30, and let $\tilde{\mathcal{A}}(L^{\mathcal{I}_{\mathcal{A}}})\coloneqq(L^{\mathcal{I}_{\mathcal{A}}}+L^{\mathcal{J}_{\mathcal{A}}}(L^{\mathcal{I}_{\mathcal{A}}}))(L^{\mathcal{I}_{\mathcal{A}}}+L^{\mathcal{J}_{\mathcal{A}}}(L^{\mathcal{I}_{\mathcal{A}}}))^{\top}$. Then $\mathcal{A}_{\Psi}(\beta,L^{\mathcal{I}_{\Sigma}})\succ 0$ if and only if $\mathcal{A}_{\Psi}(\beta,L^{\mathcal{I}_{\Sigma}})=\tilde{\mathcal{A}}(L^{\mathcal{I}_{\mathcal{A}}})$ for some $L^{\mathcal{I}_{\mathcal{A}}}\in\mathbb{L}^{n_{\mathcal{A}}}_{++}[\mathcal{I}_{\mathcal{A}}]$. Subsuming this constraint $g_{\Psi}$, we have
$\tilde{g}(\beta,L^{\mathcal{I}_{\Sigma}},L^{\mathcal{I}_{\mathcal{A}}})\coloneqq\begin{bmatrix}g_{\Psi}(\beta,L^{\mathcal{I}_{\Sigma}})\\ \textnormal{vecs}_{\mathcal{I}_{\mathcal{A}}}\left(\mathcal{A}_{\Psi}(\beta,L^{\mathcal{I}_{\Sigma}})-\tilde{\mathcal{A}}(L^{\mathcal{I}_{\mathcal{A}}})\right)\end{bmatrix}$
for all $(\beta,L^{\mathcal{I}_{\Sigma}},L^{\mathcal{I}_{\mathcal{A}}})\in{}^{n_{\beta}}\times\mathbb{L}^{n_{\Sigma}}_{++}[\mathcal{I}_{\Sigma}]\times\mathbb{L}^{n_{\mathcal{A}}}_{++}[\mathcal{I}_{\mathcal{A}}]$, where $\textnormal{vecs}_{\mathcal{I}_{\mathcal{A}}}:\mathbb{S}^{n_{\mathcal{A}}}\rightarrow{}^{|\mathcal{I}_{\mathcal{A}}|}$ vectorizes the $|\mathcal{I}_{\mathcal{A}}|$ entries of the argument corresponding to the index set $\mathcal{I}_{\mathcal{A}}$. We have proven the following lemma.
Lemma 31.
The set $\Theta_{\textnormal{NLP}}$, defined by eq. 33, equals
$\{\,(\beta,L^{\mathcal{I}_{\Sigma}})\in{}^{n_{\beta}}\times\mathbb{L}^{n_{\Sigma}}_{++}[\mathcal{I}_{\Sigma}]\;|\;\exists L^{\mathcal{I}_{\mathcal{A}}}\in\mathbb{L}^{n_{\mathcal{A}}}_{++}[\mathcal{I}_{\mathcal{A}}]:\\ \tilde{g}(\beta,L^{\mathcal{I}_{\Sigma}},L^{\mathcal{I}_{\mathcal{A}}})=0,\;h_{\Psi}(\beta,L^{\mathcal{I}_{\Sigma}}))\leq 0\,\}.$
Finally, we have the following equivalence between minimization problems over the parameter sets $\Theta$ and $\Theta_{\textnormal{NLP}}$.
Proposition 32.
For any continuous function $f:{}^{n_{\beta}}\times\mathbb{S}^{n_{\Sigma}}\rightarrow\real$ that attains a minimum in $\Theta$,
$\min_{(\beta,\Sigma)\in\Theta}f(\beta,\Sigma)=\inf_{(\beta,L^{\mathcal{I}_{\Sigma}})\in\Theta_{\textnormal{NLP}}}f_{\Psi}(\beta,L^{\mathcal{I}_{\Sigma}})$
where $f_{\Psi}\coloneqq f\circ\Psi$.
Proof.
The proof follows that of Theorem 28, noting that 2 gives $\textnormal{cl}(\Theta_{++})=\Theta$ and therefore the minimum of $f$ over $\Theta$ equals the infimum of $f$ over $\Theta_{++}$. ∎
IV-C $\varepsilon$-approximate solutions
The strict inequalities implied by constraints $\mathbb{L}^{n_{\Sigma}}_{\varepsilon}[\mathcal{I}_{\Sigma}]$ and $\mathbb{L}^{n_{\mathcal{A}}}_{\varepsilon}[\mathcal{I}_{\mathcal{A}}]$ are not amenable to implementation in standard NLP software. In [60], a log-barrier approach is used to achieve global convergence for a class of linear SDPs. Instead, we consider a constant backoff on the inequalities, providing a small but nonzero lower bound on the diagonal elements. For each $\mathcal{D}^{n}\subseteq\mathcal{I}\subseteq\mathcal{L}^{n}$, we define
$\mathbb{L}^{n}_{\varepsilon}[\mathcal{I}]\coloneqq\set{L\in\mathbb{L}^{n}_{++}[\mathcal{I}]}{L_{ii}\geq\varepsilon\;\forall i\in\mathbb{I}_{1:n}}.$
These sets imply an additional $n$ inequality constraints, but do not have any ill-posed strictness requirements. Therefore we can optimize over the restricted set
$\Theta_{\varepsilon}\coloneqq\{\,(\beta,L^{\mathcal{I}_{\Sigma}})\in{}^{n_{\beta}}\times\mathbb{L}^{n_{\Sigma}}_{\varepsilon}[\mathcal{I}_{\Sigma}]\;|\;\exists L^{\mathcal{I}_{\mathcal{A}}}\in\mathbb{L}^{n_{\mathcal{A}}}_{\varepsilon}[\mathcal{I}_{\mathcal{A}}]:\\ \tilde{g}(\beta,L^{\mathcal{I}_{\Sigma}},L^{\mathcal{I}_{\mathcal{A}}})=0,\;h_{\Psi}(\beta,L^{\mathcal{I}_{\Sigma}})\leq 0\,\}.$ (36)
In the following proposition we show, for any continuous function $f$, the infimum of $f$ over $\Theta_{\varepsilon}$ converges to the minimum of $f$ over $\Theta$, so long as 2 is satisfied (see Appendix E for proof).
Proposition 33.
For any continuous $f:{}^{n_{\beta}}\times\mathbb{S}^{n_{\Sigma}}\rightarrow\real$ that attains a minimum in $\Theta$, let $\mu_{0}\coloneqq\min_{(\beta,\Sigma)\in\Theta}f(\beta,\Sigma)$ and
$\mu_{\varepsilon}\coloneqq\inf_{(\beta,L^{\mathcal{I}_{\Sigma}})\in\Theta_{\varepsilon}}f(\beta,L^{\mathcal{I}_{\Sigma}})$ (37)
where $f(\beta,L^{\mathcal{I}_{\Sigma}})\coloneqq f(\Psi(\beta,L^{\mathcal{I}_{\Sigma}}))$. If 2 is satisfied, then $\mu_{\varepsilon}\searrow\mu$ as $\varepsilon\searrow 0$.
With requirements on the objective $f$, convergence of $\varepsilon$-approximate solutions to the exact solution is guaranteed by the following proposition (see Appendix E for proof).
Proposition 34.
For any continuous $f:{}^{n_{\beta}}\times\mathbb{S}^{n_{\Sigma}}\rightarrow\real$, consider the set-valued function $\hat{\theta}:\mathbb{R}_{\geq 0}\rightarrow\mathcal{P}(\Theta)$, defined as
$\hat{\theta}_{\varepsilon}\coloneqq\operatorname*{argmin}_{(\beta,\Sigma)\in\Psi(\Theta_{\varepsilon})}f(\beta,\Sigma)$
for all $\varepsilon>0$, and
$\hat{\theta}_{0}\coloneqq\operatorname*{argmin}_{(\beta,\Sigma)\in\Theta}f(\beta,\Sigma).$
If there exists $\alpha\in\real$ and compact $C\subseteq\Theta$ such that
$\Theta_{f\leq\alpha}\coloneqq\set{(\beta,\Sigma)\in\Theta}{f(\beta,\Sigma)\leq\alpha}$
is contained in $C$ and $\Theta_{f\leq\alpha}\cap\Theta_{++}$ is nonempty, then there exists $\overline{\varepsilon}>0$ such that, for all $\varepsilon_{0}\in[0,\overline{\varepsilon})$,
  1. (a)
    $f$ achieves a minimum in $\Theta$ and $\hat{\theta}_{0}$ is nonempty;
  2. (b)
    if $\varepsilon_{0}>0$, then $f$ achieves a minimum in $\Psi(\Theta_{\varepsilon_{0}})$ and $\hat{\theta}_{\varepsilon_{0}}$ is nonempty;
  3. (c)
    $\mu_{\varepsilon}$ is continuous and $\hat{\theta}_{\varepsilon}$ is outer semicontinuous at $\varepsilon=\varepsilon_{0}$; and
  4. (d)
    if $\hat{\theta}_{0}$ is a singleton, then $\limsup_{\varepsilon\searrow 0}\hat{\theta}_{\varepsilon}=\hat{\theta}_{0}$.
V Case Studies
In this section, we apply the identification methods outlined in the previous sections to design Kalman filters for the linear augmented disturbance models used in offset-free MPC:
$\begin{bmatrix}\hat{x}_{k+1}\\ \hat{d}_{k+1}\end{bmatrix}$ $=\begin{bmatrix}A&B_{d}\\ 0&I_{p}\end{bmatrix}\begin{bmatrix}\hat{x}_{k}\\ \hat{d}_{k}\end{bmatrix}+\begin{bmatrix}B\\ 0\end{bmatrix}u_{k}+\begin{bmatrix}K_{x}\\ K_{d}\end{bmatrix}e_{k}$ (38a)
$y_{k}$ $=\begin{bmatrix}C&C_{d}\end{bmatrix}\begin{bmatrix}\hat{x}_{k}\\ \hat{d}_{k}\end{bmatrix}+e_{k}$ (38b)
[…]
where $\hat{x}\in{}^{n}$ denote plant state estimates, $\hat{d}\in{}^{n_{d}}$ denote disturbance state estimates, $(B_{d},C_{d})\in{}^{n\times n_{d}}\times{}^{p\times n_{d}}$ denote disturbance shaping matrices, and $(K_{x},K_{d})\in{}^{n\times p}\times{}^{n_{d}\times p}$ are plant and disturbance state filter gains.
Figure 1: Benchmark temperature Control Laboratory (TCLab) [61].
In the first case study, we consider the TCLab (Figure 1), an Arduino-based temperature control laboratory that serves as a low-cost [^2] The TCLab is available for under \$40 from https://apmonitor.com/heat.htm and https://www.amazon.com/gp/product/B07GMFWMRY. benchmark for linear MIMO control [61]. We identify the TCLab from open-loop data and use the resulting model to design an offset-free MPC. We compare closed-loop control and estimation performance of these models to that of offset-free MPCs designed with the identification methods from [50, 51]. In the second case study, data from an industrial-scale chemical reactor is used to design Kalman filters for the linear augmented disturbance model, and the closed-loop estimation performance is compared to that of the designs proposed in [51].
To aid in optimizer convergence and guarantee uniqueness of the solution, to any likelihood function $f_{N}(\theta)$, we add a regularization term or prior distribution,
$\min_{\theta\in\Theta}f_{N}(\theta)-\ln p(\theta)$
where $p(\theta)$ is the prior density. Throughout, we use the following prior,
$-\ln p(\beta,\Sigma)\propto\frac{\rho}{2}\left(\|\beta-\beta_{0}\|_{2}^{2}+\textnormal{tr}(\Sigma-\Sigma_{0})\right)$ (39)
where $\rho>0$ and $(\beta_{0},\Sigma_{0})\in\Theta_{++}$ is the initial guess given to the optimizer. The prior eq. 39 is equivalent to a Frobenius norm regularizer in the Cholesky factor space,
$-\ln p(\Psi(\beta,L^{\mathcal{I}_{\Sigma}}))\propto\frac{\rho}{2}\left(\|\beta-\beta_{0}\|_{2}^{2}+\|L^{\mathcal{I}_{\Sigma}}-L^{\mathcal{I}_{\Sigma}}_{0}\|_{\mathrm{F}}^{2}\right.\\ \left.+\|L^{\mathcal{J}_{\Sigma}}(\beta,L^{\mathcal{I}_{\Sigma}})-L^{\mathcal{J}_{\Sigma}}_{0}\|_{\mathrm{F}}^{2}\right)$ (40)
where $L^{\mathcal{I}_{\Sigma}}_{0}\coloneqq\pi_{\mathcal{I}_{\Sigma}}^{L}[\textnormal{chol}(\Sigma_{0}-H(\beta_{0}))]$ and $L^{\mathcal{J}_{\Sigma}}_{0}\coloneqq L^{\mathcal{J}_{\Sigma}}(\beta_{0},L^{\mathcal{I}_{\Sigma}}_{0})$. When $\Sigma$ is block diagonal, $L^{\mathcal{J}_{\Sigma}}(\beta,L^{\mathcal{I}_{\Sigma}})=0$ and the last term of eq. 40 vanishes. The squared slack variables $L_{\mathcal{A}}=L^{\mathcal{I}_{\mathcal{A}}}+L^{\mathcal{J}_{\mathcal{A}}}$ are not regularized.
Throughout these experiments, we use the steady-state filter likelihood eq. 10 and a Cholesky factor diagonal backoff $\varepsilon$ of $10^{-6}$. When constraints are considered, they either take the form of smoothed spectral radius and abscissa constraints,
$P_{d}-\frac{A_{K}P_{d}A_{K}^{\top}}{(1-\delta)^{2}}$ $=\varepsilon_{d}I_{n+p},$ $P_{d}$ $\succeq 0,$ $\textnormal{tr}(P_{d})$ $\leq\varepsilon_{d}^{-1}$ (41a)
$A_{K}P_{c}+P_{c}A_{K}^{\top}$ $=\varepsilon_{c}I_{n+p},$ $P_{c}$ $\succeq 0,$ $\textnormal{tr}(P_{c})$ $\leq\varepsilon_{c}^{-1}$ (41b)
or the LMI region barrier function constraints,
$P_{d}-\frac{A_{K}P_{d}A_{K}^{\top}}{(1-\delta)^{2}}$ $\succeq\varepsilon_{d}I,$ $P_{d}$ $\succeq 0,$ $\textnormal{tr}(P_{d})$ $\leq\varepsilon_{d}^{-1}$ (42a)
$A_{K}P_{c}+P_{c}A_{K}^{\top}$ $\succeq\varepsilon_{c}I,$ $P_{c}$ $\succeq 0,$ $\textnormal{tr}(P_{c})$ $\leq\varepsilon_{c}^{-1}.$ (42b)
Each optimization problem is formulated in CasADi as a $\varepsilon$-approximate problem in Cholesky factor form eqs. 36 and 37, and solved with IPOPT. Wall times for a single-thread run on an Intel Core i9-10850K processor are reported. The initial guesses for these models are based on a nested ML identification approach described in [50, 51]. This approach effectively augments a standard identification method (e.g., PCA, Ho-Kalman, canonical correlation analysis algorithms), so we refer to the initial guess models as “augmented” versions of the standard method being used.
V-A TCLab
Figure 2: TCLab identification data and noise-free responses $\hat{y}_{k}=\sum_{j=1}^{k}\hat{C}\hat{A}^{j-1}\hat{B}u_{k-j}$ of a few selected models.
TABLE I: TCLab model fitting results. The augmented PCA/CCA identification methods are not iterative. ∗∗ The maximum number of iterations was set at 500.
Model Time (s) Iterations Log-likelihood
Augmented PCA 0.01 N/A 3823.4
Augmented CCA 0.04 N/A 2415.8
Unregularized ML 121.9∗∗ 500∗∗ -9431.4
Regularized ML 7.5 17 -9411.7
Constrained ML 1 36.3 61 -9407.5
Constrained ML 2 45.1 78 -9412.5
Constrained ML 3 14.2 23 -9407.5
Constrained ML 4 25.1 54 -9412.5
Figure 3: TCLab models open-loop and closed-loop (filter) eigenvalues.
Unless otherwise specified, the TCLab is modeled as a two-state system of the form eq. 1, with internal temperatures as plant states $x=\begin{bmatrix}T_{1}&T_{2}\end{bmatrix}^{\top}$, heater voltages as inputs $u=\begin{bmatrix}V_{1}&V_{2}\end{bmatrix}^{\top}$, and measured temperatures $y=\begin{bmatrix}T_{m,1}&T_{m_{2}}\end{bmatrix}^{\top}$ as outputs. Throughout, we choose $n_{d}=p$ to satisfy the offset-free necessary conditions in [37, 38], and we consider output disturbance models $(B_{d},C_{d})=(0_{2\times 2},I_{2})$. We use an observability canonical form [68] with $(A,B)$ fully parameterized and $C=I_{2}$ to guarantee model identifiability and make the states interpretable as internal temperatures. The remaining model terms $(K_{x},K_{d},R_{e})$ are fully parameterized. As in eq. 38, the models do not include a passthrough term, i.e., $D=0$.
Eight TCLab models were considered:
  1. 1.
    Augmented PCA: the 6-state TCLab model used in [50], where principle component analysis on a $400\times 5100$ data Hankel matrix is used to determine the states in the disturbance-free model.
  2. 2.
    Augmented CCA: a 2-state augmented canonical correlation analysis (CCA) model, based on the canonical method of [69].
  3. 3.
    Unregularized ML: a 2-state model, fit directly to eq. 10 without regularization or constraints, using Augmented CCA as the initial guess.
  4. 4.
    Regularized ML: the same as Unregularized ML but with an added regularizer eq. 39 with $\rho=10^{-2}N$.
  5. 5.
    Constrained ML 1: the same as Unregularized ML but with the smoothed LMI region constraints eq. 42 with $\delta=\varepsilon_{d}=10^{-3}$ and $\varepsilon_{c}=10$.
  6. 6.
    Constrained ML 2: the same as Constrained ML 1 but with $\delta=2\times 10^{-3}$.
  7. 7.
    Constrained ML 3: the same as Constrained ML 1, but using the smoothed spectral radius and abscissa constraints eq. 41 (and the same constants).
  8. 8.
    Constrained ML 4: the same as Constrained ML 3 (i.e., using eq. 41) but with $\delta=2\times 10^{-3}$.
Each ML model uses Augmented CCA as the initial guess as it has the smallest number of states. The augmented PCA model is, in effect, an unsupervised learner of the state estimates, and therefore does not produce a parsimonious state description.
In Figure 2, the identification data is presented along with the noise-free responses $\hat{y}_{k}=\sum_{j=1}^{k}\hat{C}\hat{A}^{j-1}\hat{B}u_{k-j}$ of a few selected models. Computation time, number of IPOPT iterations, and the unregularized log-likelihood value are reported in Table I. The open-loop $A$ and closed-loop $A_{K}$ eigenvalues of each model are plotted in Figure 3.
Except for the augmented PCA model, all of the open-loop eigenvalues cluster around the same region of the complex plane (fig. 3). The closed-loop filter eigenvalues are placed similarly, with the exception of the unregularized ML model, which has a slightly unstable filter. Despite these differences, the ML models all have about the same unregularized log-likelihood value (Table I) and appear to have identical noise-free responses (Figure 2). Here, the constraints eqs. 41 and 42 appear to enforce filter stability and aid in convergence with only a small computational penalty.
As reported in Table I, the unregularized ML model fitting did not converge. It is our experience that encountering iterates with unstable filters and filters with eigenvalues having negative real parts can cause convergence issues. This is because the likelihood function becomes sensitive to small changes in the parameter values due to filter instability or rapid oscillations in the filter predictions.
Figure 4: TCLab setpoint tracking tests.
Figure 5: TCLab disturbance rejection tests.
Figure 6: TCLab setpoint tracking test performance.
Figure 7: TCLab disturbance rejection test performance.
To test offset-free control performance, we performed two sets of closed-loop experiments on offset-free MPCs designed with the models. In Figure 4, identical setpoint changes were applied to a TCLab running at a steady-state power output of 50%. The setpoint changes were tracked with the offset-free MPC design described in [50]. In Figure 5, step disturbances in the output $p_{i}$ and the input $m_{i}$ are injected into a plant trying to maintain a given steady-state temperature. The setpoints are tracked with the offset-free MPC design described in [50].
Control performance indexed by squared distance from the setpoint $\ell_{k}\coloneqq\|y_{k}-y_{\textnormal{sp},k}\|_{2}^{2}$. Estimation performance is indexed by squared filter errors $e_{k}^{\top}e_{k}$. For any signal $a_{k}$, we define a $T$-sample moving average by $\langle a_{k}\rangle_{T}\coloneqq T^{-1}\sum_{j=0}^{T-1}a_{k-j}$. Setpoint tracking performance is reported in Figure 6, and disturbance rejection performance is reported in Figure 7. From Figures 6 and 7, it is clear that the regularized ML model delivers the best overall performance.
Figure 8: TCLab identification index data for (left) setpoint tracking and (right) disturbance rejection tests.
[…]
$\hat{R}_{e}^{\textnormal{aug-PCA}}$ $=\begin{bmatrix}0.5871&0.3365\\ 0.3365&0.2878\end{bmatrix},$ $\hat{R}_{e}^{\textnormal{aug-CCA}}$ $=\begin{bmatrix}0.5889&0.0918\\ 0.1791&0.3152\end{bmatrix},$
$\hat{R}_{e}^{\textnormal{ML}}$ $=\begin{bmatrix}0.0107&0.0006\\ 0.0006&0.008\end{bmatrix}.$
The reference distribution and the ML model’s $\langle q\rangle_{T}$ distribution diverge at large $T$ since, due to plant-model mismatch, the filter’s innovation errors are slightly autocorrelated.
V-B Eastman reactor
$\overline{F}_{2}$
$\overline{T}_{H}$
$\overline{F}_{1}$
$F_{1,\textrm{sp}}$
$T_{\textrm{sp}}$
$F_{2,\textrm{sp}}$
MeOH flow rate, $F_{1}$
Reactor temp, $T$
Slurry flow rate, $F_{2}$
Hot oil TC.op, $T_{H}$
Quality param 1, $r_{1}$
Quality param 2, $r_{2}$
MPC
Hot oil return
Hot oil supply
Recycle
Sludge
Xylene
Crude DMT, H2O, MeOH
Slurry
MeOH
Figure 9: Schematic of the DMT reactor and MPC control strategy.
TABLE II: Eastman reactor model fitting results. The augmented HK/CCA identification methods are not iterative.
Model Time (s) Iterations Log-likelihood
Augmented HK 0.08 N/A -7143.1
Augmented CCA 0.07 N/A -11288.4
Unregularized ML 1 23.7 93 -14345.5
Unregularized ML 2 19.4 73 -14345.5
Regularized ML 1 10.0 28 -13053.6
Regularized ML 2 8.8 22 -13587.7
Constrained ML 1 44.6 115 -13034.1
Constrained ML 2 37.1 91 -13587.7
Figure 10: Training data and noise-free responses for the Eastman reactor models (Augmented HK and ML models using Augmented HK as the initial guess).
Figure 11: Training data and noise-free responses for the Eastman reactor models (Augmented CCA and ML models using Augmented CCA as the initial guess).
Figure 12: Eastman reactor models open-loop and closed-loop (filter) eigenvalues.
A schematic of the chemical reactor considered in the next case study is presented in Figure 9.The control objective of the chemical reactor considered in the next case study is to pick three inputs (the reactant flow rates and utility temperatures $u=\begin{bmatrix}F_{1}&T_{H}&F_{2}\end{bmatrix}^{\top}$) that steer the system to three setpoints (the output, a specified reactor temperature $y=T$, and the flowrates $\begin{bmatrix}u_{1}&u_{2}\end{bmatrix}^{\top}=\begin{bmatrix}F_{1}&F_{2}\end{bmatrix}^{\top}$) without offset. See [51] for more details about the reactor operation. As in Section V-A, we choose $n_{d}=p$, consider output disturbance models $(B_{d},C_{d})=(0_{3\times 1},1)$, and use an observability canonical form [68] this time with the parameterization $A=\begin{bmatrix}0&1&0\\ 0&0&0\\ a_{1}&a_{2}&a_{3}\end{bmatrix}$ and $C=\begin{bmatrix}1&0&0\end{bmatrix}$. Again, the remaining model terms $(B,K_{x},K_{d},R_{e})$ are fully parameterized. As in eq. 38, the models do not include a passthrough term, i.e., $D=0$. We fit eight models (two augmented, six ML) of this form to closed-loop data of this core 3-input, 1-output system:
  1. 1.
    Augmented HK: a Ho-Kalman-based subspace model is augmented with a disturbance model, as detailed in [50].
  2. 2.
    Augmented CCA: a CCA subspace model, based on the method of [69], is augmented with a disturbance model, as detailed in [51].
  3. 3,4)
    Unregularized ML 1 and 2: a model is fit directly to eq. 10 without regularization or constraints, using Augmented HK and CCA (resp.) as the initial guesses.
  4. 5,6)
    Regularized ML 1 and 2: the same as Unregularized ML 1 and 2, but with an added regularizer eq. 39 with $\rho=4N$.
  5. 7,8)
    Constrained ML 1 and 2: the same as Regularized ML 1 and 2, but with the smoothed spectral radius and abscissa constraints eq. 41 with $\delta=0$, $\varepsilon_{d}=10^{-3}$, and $\varepsilon_{c}=10$.
Computation time, number of IPOPT iterations, and the unregularized log-likelihood value are reported in Table I.In Figures 10 and 11, the identification data is presented along with the noise-free responses $\hat{y}_{k}=\sum_{j=1}^{k}\hat{C}\hat{A}^{j-1}\hat{B}u_{k-j}$ of the Ho-Kalman-based and CCA-based models, respectively. The open-loop $A$ and closed-loop $A_{K}$ eigenvalues of each model are plotted in Figure 12.
The eigenvalue clustering of the reactor models is more variable than that of the TCLab models. Unregularized models converge to the same solution despite different initial guesses, and regularization pulls the eigenvalues towards the subspace model clustering around $z=0.98$. A notable feature of the unregularized models is their strong oscillating modes at $\lambda=0.7246\pm 0.3141\iota$. While these oscillations are substantially dampened by regularization and constraints, they persist for most models. These oscillations are likely true plant dynamics caused by oscillations in the lower-level PID loops (i.e., level or flow controls). This effect could be seen as an advantage or disadvantage, with the model either correctly fitting desired oscillations so they can be corrected in a supervisory MPC layer, or the model incorrectly overfitting to unwanted process dynamics. Removing unwanted process dynamics may be desirable when experimentation is costly, as it allows the user to “fix” the data rather than run another experiment.
Figure 13: Test performance for the Eastman reactor models (Augmented HK and ML models using Augmented HK as the initial guess).
Figure 14: Test performance for the Eastman reactor models (Augmented CCA and ML models using Augmented CCA as the initial guess).
The models’ estimation performances are compared against each other in Figures 13 and 14 using one of the test datasets from in [51]. Control performance could not be compared without a costly redesign of the existing MPC strategy. The unregularized models perform the best here, providing over 70% reduction in average filter error for the HK-based models. While the regularized models still capture much of this improvement, they do not have the same level of filter accuracy. This is again evidence of the oscillating modes being a real part of the reactor dynamics rather than a feature of model overfitting.
VI Conclusion
We conclude with a discussion of possible future directions of research for applying our ML identification scheme. An advantage of ML identification is the large body of literature on the asymptotic distribution and statistical efficiency of the parameter estimates. From the asymptotic distribution, decision functions can be constructed to map the on-line MPC performance to a re-identification signal. These decision functions can be constructed so as to not alarm unless sufficiently exciting data is available. Decision-theoretic re-identification therefore has lower cost and risk compared to classic adaptive control or online reinforcement learning methods that require a persistently exciting identification signal. This approach could bring statistical data efficiency to the fields of adaptive control and online reinforcement learning. Recent work on direct data-driven control has incorporated likelihood functions with measurement noise models into the control design [70]. To the best of our knowledge, no current work has considered process noise, Kalman filter forms, or structuring the model with uncontrollable integrators for offset-free MPC. There is a future possibility of direct data-driven offset-free MPC design with both optimal control and estimation performance.
Appendix A Proof of Proposition 6
Silverman [64] contains a more complete characterization of the DARE solutions for regulation problems with cross terms. However, this admits additional nullspace terms into the gain matrix which the Kalman filtering problem does not allow. We avoid nullspace terms through the assumption $R_{v}\succ 0$ and therefore streamline the proof of Proposition 6.
For the following definitions and lemmas, consider the system matrices $\mathcal{W}\coloneqq(A,B,C,D)$ corresponding to a noise-free system.
Definition 35.
The system $\mathcal{W}$ is left invertible on $\mathbb{I}_{0:k-1}$ if
$0=\begin{bmatrix}D\\ CB&D\\ \vdots&\ddots&\ddots\\ CA^{k-2}B&\ldots&CB&D\end{bmatrix}\begin{bmatrix}u_{0}\\ \vdots\\ u_{k-1}\end{bmatrix}$
implies $u_{0}=0$. The system $\mathcal{W}$ is left invertible if there is some $j\in\mathbb{N}$ such that $\mathcal{W}$ is left invertible on $\mathbb{I}_{0:k-1}$ for all $k\geq j$.
Definition 36.
The system $\mathcal{W}$ is strongly detectable if $y_{k}\rightarrow 0$ implies $x_{k}\rightarrow 0$.
The following lemmas are taken directly from [64, Thms. 8, 18(iii)], but the proofs are omitted for the sake of brevity.
Lemma 37 ([64, Thm. 8]).
If $\mathcal{W}$ is left invertible, then $\mathcal{W}$ is strongly detectable if and only if $(A-BF,C-DF)$ is detectable for all $F$ of appropriate dimension.
Lemma 38 ([64, Thm. 18(iii)]).
If $\mathcal{W}$ is left invertible, then the DARE
$P=A^{\top}PA-(A^{\top}PB+C^{\top}D)(B^{\top}PB+D^{\top}D)^{-1}\\ \times(B^{\top}PA+D^{\top}C)$
has a unique, stabilizing solution [^3] Contrary to in Section II, here we mean the solution $P$ is stabilizing when $A-BK(P)$ is stable, where $K(P)\coloneqq(B^{\top}PB+D^{\top}D)^{-1}B^{\top}P$. if and only if $\mathcal{W}$ is stabilizable and semistrongly detectable.
For the remainder of this section, we consider the full rank factorization
$\begin{bmatrix}Q_{w}&S_{wv}\\ S_{wv}^{\top}&R_{v}\end{bmatrix}=\begin{bmatrix}\tilde{B}\\ \tilde{D}\end{bmatrix}\begin{bmatrix}\tilde{B}^{\top}&\tilde{D}^{\top}\end{bmatrix}$
and the dual system $\tilde{\mathcal{W}}\coloneqq(A^{\top},C^{\top},\tilde{B}^{\top},\tilde{D}^{\top})$ to analyze the properties of the original system eq. 1. The following lemma relates the properties $R_{v}\succ 0$ and left invertability of $\tilde{\mathcal{W}}$.
Lemma 39.
If $R_{v}\succ 0$ then $\tilde{\mathcal{W}}$ is left invertible.
Proof.
Left invertability on $\mathbb{I}_{0:k-1}$ is equivalent to
$0=\begin{bmatrix}\tilde{D}^{\top}\\ \tilde{B}^{\top}C^{\top}&\tilde{D}^{\top}\\ \vdots&\ddots&\ddots\\ \tilde{B}^{\top}(A^{\top})^{k-2}C^{\top}&\ldots&\tilde{B}^{\top}C^{\top}&\tilde{D}^{\top}\end{bmatrix}\begin{bmatrix}u_{0}\\ \vdots\\ u_{k-1}\end{bmatrix}$ (43)
implying $u_{0}=0$. But $R_{v}=\tilde{D}\tilde{D}^{\top}\succ 0$, so $\tilde{D}^{\top}$ has a zero nullspace. For each $k\in\mathbb{N}$, the coefficient matrix of eq. 43 has a zero nullspace. Thus, $u_{0}=0$ and $\tilde{\mathcal{W}}$ is left invertible. ∎
Finally, we can prove Proposition 6.
Proof of Proposition 6.
By Lemma 39, we have that $\tilde{\mathcal{W}}$ is left invertible. Therefore, by Lemma 38, the DARE eq. 9 has a unique, stabilizing solution if and only if $\tilde{\mathcal{W}}$ is stabilizable and strongly detectable. But by Lemma 37 and duality, the latter statement is true if and only if $(A,C)$ is detectable and $(A-FC,\tilde{B}-F\tilde{D})$ is stabilizable for all $F\in{}^{n\times p}$. ∎
Appendix B Proof of Proposition 20
Throughout this appendix, we define the set of $n\times n$ Hermitian, Hermitian positive definite, and Hermitian positive semidefinite matrices as $\mathbb{H}^{n}$, ++$\mathbb{H}^{n}_{++}$, and +$\mathbb{H}^{n}_{+}$. Notice that $f_{\mathcal{D}}$ maps to Hermitian matrices so we can write it as $f:\mathbb{C}\rightarrow\mathbb{H}^{m}$. We define the extension of $M_{\mathcal{D}}$ to complex arguments $M_{\mathcal{D}}:\mathbb{C}^{n\times n}\times\mathbb{H}^{n}_{+}\rightarrow\mathbb{H}^{nm}$ as
$M_{\mathcal{D}}(A,P)\coloneqq M_{0}\otimes P+M_{1}\otimes(AP)+M_{1}^{\top}\otimes(AP)^{\mathrm{H}}.$
To show Proposition 20, we need a preliminary result about Hermitian positive semidefinite matrices, generalized from Lemma A.1 in [59].
Lemma 40.
For any $M\in\mathbb{H}^{n}$, if $M\succeq 0$ ($M\succ 0$) then $\textnormal{Re}(M)\succeq 0$ ($\textnormal{Re}(M)\succ 0$).
Proof.
With $M=\textnormal{Re}(M)+\iota\textnormal{Im}(M)$, it is clear $M$ Hermitian implies $\textnormal{Re}(M)$ is symmetric and $\textnormal{Im}(M)$ is skew-symmetric. Thus $v^{\top}Mv=v^{\top}\textnormal{Re}(M)v$ for all $v\in{}^{n}$, and positive (semi)definiteness of $M$ implies positive (semi)definiteness of $\textnormal{Re}(M)$. ∎
In proving Proposition 20, we take the approach of [59] but are careful to distinguish eigenvalues on the interior $\mathcal{D}$ from those on the boundary $\partial\mathcal{D}$.
Proof of Proposition 20.
($\Leftarrow$) Suppose that $M_{\mathcal{D}}(A,P)\succeq 0$ for some $P\succ 0$ and let $\lambda\in\lambda(A)$. Then there exists a nonzero $v\in\mathbb{C}^{n}$ for which $v^{\mathrm{H}}A=\lambda v^{\mathrm{H}}$. Consider the identity
$(I_{m}$ $\otimes v)^{\mathrm{H}}M_{\mathcal{D}}(A,P)(I_{m}\otimes v)$
$=M_{0}\otimes v^{\mathrm{H}}Pv+M_{1}\otimes(v^{\mathrm{H}}APv)+M_{1}^{\top}\otimes(v^{\mathrm{H}}PA^{\top}v)$
$=M_{0}\otimes v^{\mathrm{H}}Pv+M_{1}\otimes(\overline{\lambda}v^{\mathrm{H}}Pv)+M_{1}^{\top}\otimes(\lambda v^{\mathrm{H}}Pv)$
$=v^{\mathrm{H}}Pv(M_{0}+M_{1}\lambda+M_{1}^{\top}\overline{\lambda})$
$=v^{\mathrm{H}}Pvf_{\mathcal{D}}(\lambda).$
The assumption $P\succ 0$ implies $v^{\mathrm{H}}Pv>0$, and $M_{\mathcal{D}}(A,P)\succeq 0$ further implies $f_{\mathcal{D}}(\lambda)\succeq 0$. Therefore $\lambda\in\textnormal{cl}(\mathcal{D})$.
Next suppose $\lambda\in\lambda(A)$ is non-simple and $\lambda\in\partial\mathcal{D}$. Then there exists nonzero $v_{1},v_{2}\in\mathbb{C}^{n}$ (linearly independent) such that $v^{\mathrm{H}}f_{\mathcal{D}}(\lambda)v=0$, $v_{1}^{\mathrm{H}}A=\lambda v_{1}^{\mathrm{H}}$, and $v_{2}^{\mathrm{H}}A=\lambda v_{2}^{\mathrm{H}}+v_{1}$. Because $\mathcal{D}$ is open, $\lambda\in\partial\mathcal{D}=\textnormal{cl}(\mathcal{D})\setminus\mathcal{D}$ must satisfy both $f_{\mathcal{D}}(\lambda)\succeq 0$ and $f_{\mathcal{D}}(\lambda)\not\succ 0$. Therefore $f_{\mathcal{D}}(\lambda)$ is singular, and there exists a nonzero vector $v\in\mathbb{C}^{m}$ such that $v^{\mathrm{H}}f_{\mathcal{D}}(\lambda)v=0$. With the $2\times 2$ matrices
$\tilde{P}$ $=\begin{bmatrix}p_{11}&p_{12}\\ p_{12}&p_{22}\end{bmatrix}\coloneqq\begin{bmatrix}v_{1}^{\mathrm{H}}\\ v_{2}^{\mathrm{H}}\end{bmatrix}P\begin{bmatrix}v_{1}&v_{2}\end{bmatrix}\succ 0$
$\tilde{J}$ $\coloneqq\lambda I_{2}+\begin{bmatrix}0&1\\ 0&0\end{bmatrix}$
we have $\begin{bmatrix}v_{1}&v_{2}\end{bmatrix}^{\mathrm{H}}A=\tilde{J}\begin{bmatrix}v_{1}&v_{2}\end{bmatrix}^{\mathrm{H}}$ and therefore
$(I_{m}$ $\otimes\begin{bmatrix}v_{1}&v_{2}\end{bmatrix})^{\mathrm{H}}M_{\mathcal{D}}(A,P)(I_{m}\otimes\begin{bmatrix}v_{1}&v_{2}\end{bmatrix})$
$=M_{0}\otimes\tilde{P}+M_{1}\otimes\tilde{J}\tilde{P}+M_{1}^{\top}\otimes(\overline{\tilde{J}}\tilde{P})^{\top}$
$=M_{\mathcal{D}}(\tilde{J},\tilde{P})\succeq 0.$
Next, we have
$\tilde{M}$ $\coloneqq K_{2,m}M_{\mathcal{D}}(\tilde{J},\tilde{P})K_{2,m}^{\top}$
$=\tilde{P}\otimes M_{0}+\tilde{J}\tilde{P}\otimes M_{1}+(\overline{\tilde{J}}\tilde{P})^{\top}\otimes M_{1}^{\top}$
$=\tilde{P}\otimes f_{\mathcal{D}}(\lambda)+\begin{bmatrix}p_{12}(M_{1}+M_{1}^{\top})&p_{22}M_{1}\\ p_{22}M_{1}^{\top}&0\end{bmatrix}\succeq 0.$
Finally,
$(I_{2}\otimes v)^{\mathrm{H}}\tilde{M}(I_{2}\otimes v)\\ =\begin{bmatrix}p_{12}v^{\mathrm{H}}(M_{1}+M_{1}^{\top})v&p_{22}v^{\mathrm{H}}M_{1}v\\ p_{22}v^{\mathrm{H}}M_{1}^{\top}v&0\end{bmatrix}\succeq 0.$
But $\tilde{P}\succ 0$ implies $p_{22}>0$, so the above matrix inequality implies $v^{\mathrm{H}}M_{1}v=0$. Moreover, with $v^{\mathrm{H}}f_{\mathcal{D}}(\lambda)v=0$, we also have $v^{\mathrm{H}}M_{0}v=0$ and therefore $f(z)\equiv 0$ and $\mathcal{D}$ is empty, a contradiction. Therefore each $\lambda\in\lambda(A)$ non-simple implies $\lambda\in\mathcal{D}$.
($\Rightarrow$) Suppose $\lambda(A)\subset\textnormal{cl}(\mathcal{D})$ and $\lambda\in\lambda(A)$ non-simple implies $\lambda\in\mathcal{D}$.
If $A=\lambda$ is a (possibly complex) scalar, then it lies in $\textnormal{cl}(\mathcal{D})$ by assumption, with $M_{\mathcal{D}}(\lambda,p)=pf_{\mathcal{D}}(\lambda)\succeq 0$ for all $p>0$.
If $A=\lambda I_{n}+N$ is a (possibly complex) Jordan block, where $N\in{}^{n\times n}$ is a shift matrix and $n>1$, then $\lambda\in\mathcal{D}$ and $f_{\mathcal{D}}(\lambda)\succ 0$. Let $T_{k}\coloneqq\textnormal{diag}(k^{n-1},\ldots,k,1)$ for each $k\in\mathbb{N}$. Then $T_{k}^{-1}AT_{k}=\lambda I_{n}+k^{-1}N\rightarrow\lambda I_{n}$ as $k\rightarrow\infty$. Moreover, because $M_{\mathcal{D}}$ is continuous, we have
$M_{\mathcal{D}}(T_{k}^{-1}AT_{k},I_{n})\rightarrow M_{\mathcal{D}}(\lambda I_{n},I_{n})=f_{\mathcal{D}}(\lambda)\otimes I_{n}\succ 0.$
Therefore there exists some $k_{0}\in\mathbb{N}$ such that $M_{\mathcal{D}}(T_{k}^{-1}AT_{k},I_{n})\succ 0$ for all $k\geq k_{0}$. With $P\coloneqq T_{k}T_{k}^{\top}$, we have
$M_{\mathcal{D}}(A,P)$ $=M_{0}\otimes T_{k}T_{k}^{\top}+M_{1}\otimes(AT_{k}T_{k}^{\top})$
$\qquad+M_{1}^{\top}\otimes(AT_{k}T_{k}^{\top})^{\top}$
$=(I_{m}\otimes T_{k})(M_{0}\otimes I_{n}+M_{1}\otimes T_{k}^{-1}AT_{k}$
$\qquad+M_{1}^{\top}\otimes(T_{k}^{-1}AT_{k})^{\top})(I_{m}\otimes T_{k})^{\top}$
$=(I_{m}\otimes T_{k})M_{\mathcal{D}}(T_{k}^{-1}AT_{k},I_{n})(I_{m}\otimes T_{k})^{\top}\succ 0.$
Finally, for any $A\in{}^{n\times n}$, let $A=V(\bigoplus_{i=1}^{p}J_{i})V^{-1}$ denote the Jordan decomposition of $A$, where $J_{i}=\lambda_{i}I_{n_{i}}+N_{i}$, $\lambda_{i}\in\lambda(A)$, $N_{i}$ are shift matrices, and $n=\sum_{i=1}^{p}n_{i}$. We have already shown that for each $i\in\mathbb{I}_{1:p}$, there exists $P_{i}\succ 0$ such that $M_{\mathcal{D}}(J_{i},P_{i})\succeq 0$. Then with $\tilde{P}\coloneqq V(\bigoplus_{i=1}^{p}P_{i})V^{-1}$, we have
$(I_{m}$ $\otimes V^{-1})M_{\mathcal{D}}(A,\tilde{P})(I_{m}\otimes V^{-1})^{\mathrm{H}}$
$=M_{0}\otimes\left(\bigoplus_{i=1}^{p}P_{i}\right)+M_{1}\otimes\left(\bigoplus_{i=1}^{p}J_{i}P_{i}\right)$
$\qquad+M_{1}\otimes\left(\bigoplus_{i=1}^{p}J_{i}P_{i}\right)^{\top}$
$=K_{n,m}\left(\bigoplus_{i=1}^{p}K_{m,n_{i}}M_{\mathcal{D}}(J_{i},P_{i})K_{m,n_{i}}^{\top}\right)K_{n,m}^{\top}\succeq 0$
and therefore $M_{\mathcal{D}}(A,\tilde{P})\succeq 0$. Last, Lemma 40 gives $M_{\mathcal{D}}(A,P)\succeq 0$ with $P\coloneqq\textrm{Re}(\tilde{P})$ since
$M_{\mathcal{D}}(A,P)=M_{\mathcal{D}}(A,\textrm{Re}(\tilde{P}))=\textrm{Re}(M_{\mathcal{D}}(A,\tilde{P})).\qed$
Appendix C Proof of Proposition 21
To show Proposition 21(a), we first require the following eigenvalue sensitivity result due to [63, Thm. 7.2.3].
Theorem 41 ([63, Thm. 7.2.3]).
For any $A\in\mathbb{C}^{n\times n}$, denote its Schur decomposition by $A=Q(D+N)Q^{\mathrm{H}}$, where $Q\in\mathbb{C}^{n\times n}$ is unitary, $D\in\mathbb{C}^{n\times n}$ is diagonal, and $N\in\mathbb{C}^{n\times n}$ is strictly upper triangular. [^4] A matrix $U$ is strictly upper triangular if $U_{ij}=0$ for all $i\geq j$. Let $p$ be the smallest integer for which $M^{p}=0$ where $M_{ij}\coloneqq|N_{ij}|$. Then
$\min_{\lambda\in\lambda(A)}|\mu-\lambda|\leq\max\set{c}{E\|,(c\|E\|)^{1/p}}$
where $c\coloneqq\sum_{k=0}^{p-1}\|N\|^{k}$.
Proof of Proposition 21.
Throughout this proof, we show a set $S$ is not open (or not closed) by demonstrating that $S^{c}$ (or $S$) does not contain all its limit points.
(a)—For any $A\in\mathbb{A}_{\mathcal{D}}^{n}$, continuity of $f_{\mathcal{D}}$ gives the existence of a function $\delta(\lambda)>0$ such that $f_{\mathcal{D}}(z)\succ 0$ for all $|z-\lambda|<\delta(\lambda)$ and $\lambda\in\lambda(A)$. Let $\delta\coloneqq\min_{\lambda\in\lambda(A)}\delta(\lambda)$. By Theorem 41 and norm equivalence, there exist $c>0$ and $p\in\mathbb{I}_{1:n}$ such that
$\max_{\mu\in\lambda(A+E)}\min_{\lambda\in\lambda(A)}|\lambda-\mu|\leq\max\set{c}{E\|_{\textnormal{F}},(c\|E\|_{\textnormal{F}})^{1/p}}$
for all $E\in{}^{n\times n}$. Therefore there exists a $\varepsilon>0$ such that
$\max_{\mu\in\lambda(A+E)}\min_{\lambda\in\lambda(A)}|\lambda-\mu|<\delta$
for all $E\in\mathcal{B}\coloneqq\set{E^{\prime}\in{}^{n\times n}}{\|E^{\prime}\|_{\textnormal{F}}<\varepsilon}$. Finally, $A+\mathcal{B}$ is a neighborhood of $A$ contained in $\mathbb{A}_{\mathcal{D}}^{n}$, and, since $A\in\mathbb{A}_{\mathcal{D}}^{n}$ was chosen arbitrarily, $\mathbb{A}_{\mathcal{D}}^{n}$ is open.
(b)(i)—Because $\mathcal{D}$ is open, nonempty, and not equal to $\mathcal{D}$, $\partial\mathcal{D}$ is nonempty. Let $\lambda\in\partial\mathcal{D}$ and $\lambda_{k}\in\mathcal{D}^{c}$ be a sequence for which $\lambda_{k}\rightarrow\lambda$. By symmetry, we also have $\overline{\lambda}\in\mathcal{D}$ and $\overline{\lambda}_{k}\in\mathcal{D}^{c}$.
For $n=2$, we have $A\coloneqq\begin{bmatrix}\textnormal{Re}(\lambda)&-\textnormal{Im}(\lambda)\\ \textnormal{Im}(\lambda)&\textnormal{Re}(\lambda)\end{bmatrix}\in{}^{2\times 2}$ has eigenvalues $\lambda,\overline{\lambda}\in\mathcal{D}$, and $A_{k}\coloneqq\begin{bmatrix}\textnormal{Re}(\lambda_{k})&-\textnormal{Im}(\lambda_{k})\\ \textnormal{Im}(\lambda_{k})&\textnormal{Re}(\lambda_{k})\end{bmatrix}\in{}^{2\times 2}$ has eigenvalues $\lambda_{k},\overline{\lambda_{k}}\in\mathcal{D}^{c}$ for each $k\in\mathbb{N}$. The corresponding eigenvectors are [±$\begin{bmatrix}\pm\iota\\ 1\end{bmatrix}\in\mathbb{C}^{2}$. Therefore $A\in\tilde{\mathbb{A}}_{\mathcal{D}}^{2}$ but $A_{k}\in(\tilde{\mathbb{A}}_{\mathcal{D}}^{2})^{c}$ for each $k\in\mathbb{N}$, and the limit $A_{k}\rightarrow A$ gives us that $(\tilde{\mathbb{A}}_{\mathcal{D}}^{2})^{c}$ does not contain all its limit points.
For $n>2$, let $A_{0}\in\tilde{\mathbb{A}}_{\mathcal{D}}^{n-2}$, and we can extend the prior argument with the sequence $B_{k}\coloneqq A_{k}\oplus A_{0}\in(\tilde{\mathbb{A}}_{\mathcal{D}}^{n})^{c},k\in\mathbb{N}$ that converges to $B\coloneqq A\oplus A_{0}\in\tilde{\mathbb{A}}_{\mathcal{D}}^{n}$.
(b)(ii)—By part (b)(i), it suffices to consider the case $n=1$. By closure and convexity of $\mathcal{D}$, $\mathcal{D}\cap\real$ is either a closed line segment, a closed ray, or itself. In other words, $\mathcal{D}\cap\real$ is open if and only if it has no endpoints. Moreover, since $\partial\mathcal{D}\cap\real$ is the set of the endpoints of $\mathcal{D}\cap\real$, $\mathcal{D}\cap\real$ is open if and only if $\partial\mathcal{D}\cap\real$ is empty. Finally, since $\tilde{\mathbb{A}}_{\mathcal{D}}^{1}=\mathcal{D}\cap\real$, $\tilde{\mathbb{A}}_{\mathcal{D}}^{1}$ is open if and only if $\partial\mathcal{D}\cap\real$ is empty.
(c)(i)—Let $\lambda\in\partial\mathcal{D}$. Suppose $n=4$. Then $\overline{\lambda}\in\partial\mathcal{D}$ by symmetry. Because $\mathcal{D}$ is open, there exists a sequence $\lambda_{k}\in\mathcal{D}$ such that $\lambda_{k}\rightarrow\lambda$, and by symmetry, we also have $\overline{\lambda}_{k}\in\mathcal{D}$ and $\overline{\lambda}_{k}\rightarrow\overline{\lambda}$. Consider again the $2\times 2$ matrices $A$ and $A_{k}$ from part (b)(i), which have eigenvalues $\lambda,\overline{\lambda}\in\mathcal{D}$ and $\lambda_{k},\overline{\lambda_{k}}\in\mathcal{D}^{c}$, respectively. Then the block matrices $B\coloneqq\begin{bmatrix}A&I_{2}\\ 0&A\end{bmatrix}\in{}^{4\times 4}$ and $B_{k}\coloneqq\begin{bmatrix}A_{k}&I_{2}\\ 0&A_{k}\end{bmatrix}\in{}^{4\times 4}$ have the same eigenvalues, but this time the eigenvectors are [±$\begin{bmatrix}\pm\iota\\ 1\\ 0\\ 0\end{bmatrix},\begin{bmatrix}0\\ 0\\ \pm\iota\\ 1\end{bmatrix}\in\mathbb{C}^{4}$ and the eigenvalues are non-simple. Since $\lambda$ is a non-simple eigenvalue on the boundary of $\mathcal{D}$, we have $B\not\in\tilde{\mathbb{A}}_{\mathcal{D}}^{4}$. However, $\lambda_{k}$ are all in the interior of $\mathcal{D}$, so $B_{k}\in\tilde{\mathbb{A}}_{\mathcal{D}}^{4}$. Since $B_{k}\rightarrow B$, the set $\tilde{\mathbb{A}}_{\mathcal{D}}^{4}$ does not contain all its limit points.
On the other hand, let $\lambda\in\partial\mathcal{D}$ and suppose $n>4$. Similarly to part (b)(i), with any $\tilde{A}_{0}\in\tilde{\mathbb{A}}_{\mathcal{D}}^{n-4}$, we can extend the argument for the $n=4$ case with the sequence $\tilde{A}_{k}\coloneqq B_{k}\oplus\tilde{A}_{0}\in\tilde{\mathbb{A}}_{\mathcal{D}}^{n},k\in\mathbb{N}$ that converges to $\tilde{A}\coloneqq B\oplus\tilde{A}_{0}\in(\tilde{\mathbb{A}}_{\mathcal{D}}^{n})^{c}$.
(c)(ii)—Let $\lambda\in\partial\mathcal{D}\cap\real$ and $n\geq 2$. Because $\mathcal{D}$ is convex, open, and nonempty, there exists $\varepsilon>0$ such that exactly one of the real intervals $(\lambda,\lambda+\varepsilon)$ or $(\lambda-\varepsilon,\lambda)$ is contained in $\mathcal{D}$, whereas the other is contained in $\textnormal{int}(\mathcal{D}^{c})$. Without loss of generality, assume $(\lambda-\varepsilon,\lambda)\subseteq\mathcal{D}$ [^5] Otherwise, take the reflection about the imaginary axis $-\mathcal{D}$ and $-\tilde{\mathbb{A}}_{\mathcal{D}}^{n}$. Then $A_{k}\coloneqq(\lambda-\varepsilon/k)I_{n}+N_{n}\in\tilde{\mathbb{A}}_{\mathcal{D}}^{n}$ for each $k\in\mathbb{N}$, but $A_{k}\rightarrow\lambda I_{n}+N_{n}\in(\tilde{\mathbb{A}}_{\mathcal{D}}^{n})^{c}$ and therefore $\tilde{\mathbb{A}}_{\mathcal{D}}^{n}$ does not contain all its limit points. (d)—Since $\overline{\mathbb{A}}_{\mathcal{D}}^{n}\coloneqq\set{A\in{}^{n\times n}}{\lambda(A)\subset\textnormal{cl}(\mathcal{D})}$ contains $\mathbb{A}_{\mathcal{D}}^{n}$, it suffices to show any $A\in\overline{\mathbb{A}}_{\mathcal{D}}^{n}$ is a limit point of $\mathbb{A}_{\mathcal{D}}^{n}$. Denote the Jordan form by $A=V\left(\bigoplus_{i=1}^{p}\mu_{i}I_{n_{i}}+N_{n_{i}}\right)V^{-1}$, where $V\in{}^{n\times n}$ is invertible, $\mu_{i}\in\lambda(A)$, $n=\sum_{i=1}^{p}n_{i}$, and $N_{i}\in{}^{n_{i}\times n_{i}}$ is a shift matrix. Because $\mu_{i}\in\textnormal{cl}(\mathcal{D})$, there exists a sequence $\mu_{i,k}\in\mathcal{D}$ such that $\mu_{i,k}\rightarrow\mu_{i}$. Then $A_{k}\coloneqq V\left(\bigoplus_{i=1}^{p}\mu_{i,k}I_{n_{i}}+N_{i}\right)V^{-1}\in\mathbb{A}_{\mathcal{D}}^{n}$ and $A_{k}\rightarrow A$. ∎
Appendix D Proof of Proposition 22
To prove Proposition 22(a,b), we use sensitivity results on the value functions of parameterized nonlinear SDPs,
$V(y)\coloneqq\inf_{x\in\mathbb{X}(y)}F(x,y)$ (44)
where the set-valued function $\mathbb{X}:{}^{m}\rightarrow\mathcal{P}({}^{n})$ is defined by
$\mathbb{X}(y)\coloneqq\set{x\in{}^{n}}{G(x,y)\succeq 0}.$
Consider also the graph of the set-valued function $\mathbb{X}$,
$\mathbb{Z}\coloneqq\set{(x,y)\in{}^{n+m}}{G(x,y)\succeq 0}.$
Notice that $\mathbb{Z}$ is closed if $G$ is continuous. We say Slater’s condition holds at $y\in{}^{m}$ if there exists $x\in{}^{n}$ such that $x\in\textnormal{int}(\mathbb{X}(y))$, or equivalently, $G(x,y)\succ 0$.
In [71, Prop. 4.4], continuity of a general class of optimization problems is considered. In the following proposition, we state the specialization to nonlinear SDPs.
Proposition 42 ([71, Prop. 4.4]).
Let $y_{0}\in{}^{m}$ and suppose
  1. (i)
    $F$ and $G$ are continuous on n+m;
  2. (ii)
    there exist $\alpha\in\real$ and compact $C\subset{}^{n}$ such that, for each $y$ in a neighborhood of $y_{0}$, the level set
    $\textnormal{lev}_{\leq\alpha}F(\cdot,y)\coloneqq\set{x\in\mathbb{X}(y)}{F(x,y)\leq\alpha}$
    is nonempty and contained in $C$; and
  3. (iii)
    Slater’s condition holds at $y_{0}$.
Then $F(\cdot,y)$ attains a minimum on $\mathbb{X}(y)$ for all $y\in N_{y}$, and $V(y)$ is continuous at $y=y_{0}$.
Proof.
See [71, Prop. 4.4] and the discussions in [71, pp. 264, 483–484, 491–492]. ∎
Finally, we prove Proposition 22.
Proof of Proposition 22.
Let $\textnormal{vec}:{}^{n\times n}\rightarrow{}^{n^{2}}$ and $\textnormal{vecs}:{}^{n\times n}\rightarrow{}^{(1/2)(n+1)n}$ denote the vectorization and symmetric vectorization operators, respectively.
(a)—With $x\coloneqq\textnormal{vecs}(P)$, $y\coloneqq\textnormal{vec}(A)$, $F(x,y)\coloneqq\textnormal{tr}(VP)$, and $G(x,y)\coloneqq P\oplus(M_{\mathcal{D}}(A,P)-M)$, we can use Proposition 42 to show the continuity of $\phi_{\mathcal{D}}$ on $\mathbb{A}_{\mathcal{D}}^{n}$. Let $A_{0}\in\mathbb{A}_{\mathcal{D}}^{n}$. Condition (i) of Proposition 42 holds by assumption. Slater’s condition (iii) holds because for any $P\succ 0$ such that $M_{\mathcal{D}}(A_{0},P)\succ 0$, we can define $P_{0}\coloneqq\gamma P\succ 0$ for some $\gamma>\gamma_{0}\coloneqq\|M\|\times\|[M_{\mathcal{D}}(A_{0},P)]^{-1}\|$ to give
$M_{\mathcal{D}}(A_{0},P_{0})=\gamma M_{\mathcal{D}}(A_{0},P)\succ\gamma_{0}M_{\mathcal{D}}(A_{0},P)\succeq M.$
Moreover, by continuity of $M_{\mathcal{D}}$, there exists a neighborhood $N_{A}$ of $A_{0}$ such that $M_{\mathcal{D}}(A,P_{0})\succ M$ for all $A\in N_{A}$. Letting $\alpha\coloneqq\textnormal{tr}(VP_{0})>0$, we have that the set
$\set{P\in\mathbb{S}^{n}_{+}}{\textnormal{tr}(VP)\leq\alpha}$
is compact and contains the nonempty level set
$\set{P\in\mathbb{P}(A)}{\textnormal{tr}(VP)\leq\alpha}$
for all $A\in N_{A}$. Taking the image of each of the above sets under the vecs operation gives condition (ii) of Proposition 42. All the conditions of Proposition 42 are thus satisfied for each $A_{0}\in\mathbb{A}_{\mathcal{D}}^{n}$, and we have $\phi_{\mathcal{D}}$ is continuous on $\mathbb{A}_{\mathcal{D}}^{n}$.
(b)—Continuity of $\phi_{\mathcal{D}}$ on $\mathbb{A}_{\mathcal{D}}^{n}$ implies closure of the sublevel sets of $\phi_{\mathcal{D}}$, and eq. 24 follows by definition of $\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon)$.
(c)—First, $M_{\mathcal{D}}(A,P)\succ 0$ implies $P\succ 0$ since, if $M_{\mathcal{D}}(A,P)\succ 0$ and $P\succeq 0$ but $P\not\succ 0$, there exists a nonzero $v\in{}^{n}$ such that $Pv=0$ and
$(I_{m}\otimes v)^{\top}(M_{\mathcal{D}}(A,P))(I_{m}\otimes v)=M\otimes(v^{\top}Pv)\\ +M_{1}\otimes(v^{\top}APv)+M_{1}^{\top}\otimes(v^{\top}PA^{\top}v)=0$
a contradiction of the assumption $M_{\mathcal{D}}(A,P)\succ 0$. Moreover, for any $P\succ 0$ such that $M_{\mathcal{D}}(A,P)\succ 0$, we have $M_{\mathcal{D}}(A,P)\succeq\gamma M_{\mathcal{D}}(A,P)\succeq M$ with $P\coloneqq\gamma P$ and $\gamma\coloneqq\|M\|\times\|[M_{\mathcal{D}}(A,P)]^{-1}\|$, so feasibility of eq. 17 is equivalent to feasibility of
$M_{\mathcal{D}}(A,P)$ $\succ M,$ $P$ $\succeq 0$
and therefore $\bigcup_{\varepsilon>0}\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon)=\mathbb{A}_{\mathcal{D}}^{n}$. But $\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon)$ is monotonically decreasing, [^6] By “monotonically decreasing” we mean $\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon)\supseteq\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon^{\prime})$ for all $\varepsilon\leq\varepsilon^{\prime}$. so $\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon)\nearrow\bigcup_{\varepsilon>0}\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon)=\mathbb{A}_{\mathcal{D}}^{n}$ as $\varepsilon\searrow 0$. ∎
Appendix E Proof of Propositions 33 and 34
Starting with Proposition 33:
Proof of Proposition 33.
Since $\mu_{\varepsilon}$ is nondecreasing and bounded from below by $\mu$, it suffices to show that for each $\delta>0$, there exists a $\overline{\varepsilon}>0$ such that $\mu_{\overline{\varepsilon}}-\mu<\delta$.
Let $(\beta^{*},\Sigma^{*})\in\Theta$ denote a point for which $\mu=f(\beta^{*},\Sigma^{*})$. If $(\beta^{*},\Sigma^{*})\in\Theta_{++}$, we could simply choose $\overline{\varepsilon}>0$ small enough to put $(\beta^{*},\Sigma^{*})$ in $\Theta_{\overline{\varepsilon}}$ and achieve $\mu_{\overline{\varepsilon}}-\mu=0<\delta$.
Instead, we assume $(\beta^{*},\Sigma^{*})\not\in\Theta_{++}$. By 2, there exists a sequence $(\beta_{k},\Sigma_{k})\in\Theta_{++},k\in\mathbb{N}$ such that $\beta_{k}\rightarrow\beta$ and $\Sigma_{k}\rightarrow\Sigma$ as $k\rightarrow\infty$. Defining $\nu_{k}\coloneqq f(\beta_{k},\Sigma_{k})$, we have $\nu_{k}\rightarrow\mu$ by continuity of $f$. Therefore, there exists some $k_{0}\in\mathbb{N}$ such that $\nu_{k}-\mu<\delta$ for all $k\geq k_{0}$.
For each $(\beta_{k},\Sigma_{k})\in\Theta_{++}$, there exist unique $L^{\mathcal{I}_{\Sigma}}_{k}\in\mathbb{L}^{n_{\Sigma}}_{++}[\mathcal{I}_{\Sigma}]$ and $L^{\mathcal{I}_{\mathcal{A}}}_{k}\in\mathbb{L}^{n_{\mathcal{A}}}_{++}[\mathcal{I}_{\mathcal{A}}]$ such that the constraints
$\tilde{g}(\beta_{k},L^{\mathcal{I}_{\Sigma}}_{k},L^{\mathcal{I}_{\mathcal{A}}}_{k})$ $=0$ $h(\beta_{k},L^{\mathcal{I}_{\Sigma}}_{k})$ $\leq 0$
are satisfied (by Lemmas 29 and 31). Let $\overline{\varepsilon}$ be the minimum value over all the diagonal elements of $L^{\mathcal{I}_{\Sigma}}_{k_{0}}$ and $L^{\mathcal{I}_{\mathcal{A}}}_{k_{0}}$. Then $(\beta_{k_{0}},L^{\mathcal{I}_{\Sigma}}_{k_{0}})\in\Theta_{\overline{\varepsilon}}$ by construction, so $\nu_{k_{0}}\geq\mu_{\overline{\varepsilon}}$ by optimality, and therefore $\mu_{\overline{\varepsilon}}-\mu\leq\nu_{k_{0}}-\mu<\delta$. ∎
As in Appendix D, we use sensitivity results of [71] on optimization problems to to prove Proposition 34. This time, however, we consider the continuity of the value function for parameterized NLPs on Banach spaces. Let $\mathcal{X}$, $\mathcal{Y}$, and $\mathcal{K}$ be Banach spaces and consider the parameterized NLP,
$V(y)\coloneqq\inf_{x\in\mathbb{X}(y)}F(x,y)$ (45)
where the set-valued function $\mathbb{X}:\mathcal{Y}\rightarrow\mathcal{P}(\mathcal{X})$ is defined by
$\mathbb{X}(y)\coloneqq\set{x\in\mathcal{X}}{G(x,y)\in K}$
for some $G:\mathcal{X}\times\mathcal{Y}\rightarrow\mathcal{K}$ and $K\subseteq\mathcal{K}$ is closed. Let $X^{0}(y)$ denote the (possibly empty) set of solutions to eq. 45. Define the graph of the set-valued function $\mathbb{X}(\cdot)$ by
$\mathbb{Z}\coloneqq\set{(x,y)\in\mathcal{X}\times\mathcal{Y}}{G(x,y)\in K}.$
Notice that $\mathbb{Z}$ is closed if $G$ is continuous and $K$ is closed.
Proposition 43 ([71, Prop. 4.4]).
Let $y_{0}\in\mathcal{Y}$ and assume:
  1. (i)
    $F$ and $G$ are continuous on $\mathcal{X}\times\mathcal{Y}$ and $K$ is closed;
  2. (ii)
    there exist $\alpha\in\real$ and a compact set $C\subseteq\mathcal{X}$ such that, for every $y$ in a neighborhood of $y_{0}$, the level set
    $\set{x\in\mathbb{X}(y)}{f(x,y)\leq\alpha}$
    is nonempty and contained in $C$; and
  3. (iii)
    for any neighborhood $N_{x}$ of the solution set $X^{0}(y_{0})$, there exists a neighborhood $N_{y}$ of $y_{0}$ such that $N_{x}\cap\mathbb{X}(y)$ is nonempty for all $y\in N_{y}$;
then $V(y)$ is continuous and $X^{0}(y)$ is outer semicontinuous at $y=y_{0}$.
Proof of Proposition 34.
First, we must specify $\overline{\varepsilon}$. For each $\theta\in\Theta_{++}$, let
$\varepsilon(\theta)\coloneqq\max\set{\varepsilon>0}{\theta\in\Psi(\Theta_{\varepsilon})}$
where the maximum is achieved since there is a finite number of diagonal elements of the Cholesky factors that must be lower bounded. Now we specify $\overline{\varepsilon}$ as the supremum of $\varepsilon(\theta)$ over all $\theta\in\Theta_{f\leq\alpha}\cap\Theta_{++}$,
$\overline{\varepsilon}\coloneqq\sup\set{\varepsilon(\theta)}{\theta\in\Theta_{f\leq\alpha}\cap\Theta_{++}}$
so that, for any $\varepsilon\in(0,\overline{\varepsilon})$, $\Theta_{f\leq\alpha}\cap\Psi(\Theta_{\varepsilon})$ is nonempty and is contained in the compact set $C$.
(a)—Following the proof of [71, Prop. 4.4], we have (i) $F$ is continuous and (ii) the level set $\Theta_{f\leq\alpha}$ is nonempty and contained in the compact set $C$, which implies $\Theta_{f\leq\alpha}$ is a compact level set and therefore the minimum of $f$ over $\Theta_{f\leq\alpha}$ is achieved and equals the minimum over $\Theta$. Moreover, $\hat{\theta}_{0}$ must be nonempty.
(b)—Similarly to part (a), we have, for each $\varepsilon\in(0,\overline{\varepsilon})$, that the level set $\Theta_{f\leq\alpha}\cap\Psi(\Theta_{\varepsilon})$ is nonempty and contained in the compact set $C$, so $f$ achieves its minimum over $\Psi(\Theta_{\varepsilon})$ and $\hat{\theta}_{\varepsilon}$ is nonempty.
(c)—Consider the graph of the constraint function,
$\mathbb{Z}\coloneqq\set{(\theta,\varepsilon)\in\Theta\times\mathbb{R}_{\geq 0}}{\theta\in\Psi(\Theta_{\varepsilon})\textnormal{ if }\varepsilon>0}.$
Consider a sequence $(\theta_{k},\varepsilon_{k})\in\mathbb{Z},k\in\mathbb{N}$ that is convergent $(\theta_{k},\varepsilon_{k})\rightarrow(\theta,\varepsilon)$. Then $\varepsilon\geq 0$, otherwise the sequence would not converge. Moreover, $\theta\in\Theta$ since $\theta_{k}\in\Psi(\Theta_{\varepsilon_{k}})\subseteq\Theta$ for all $k\in\mathbb{N}$ and $\Theta$ contains all its limit points. If $\varepsilon=0$, then $(\theta,\varepsilon)\in\mathbb{Z}$ trivially. On the other hand, if $\varepsilon>0$, then $\varepsilon(\theta_{k})$ converges to $\varepsilon(\theta)$ because $\Psi$ is continuous and the max can be taken over a finite number of elements of $\Psi^{-1}(\theta_{k})$. Moreover, $\varepsilon(\theta_{k})$ and upper bounds $\varepsilon_{k}$ because $\theta_{k}\in\Psi(\Theta_{\varepsilon_{k}})$, so $\varepsilon(\theta)\geq\varepsilon$. Finally, we have $\theta\in\Psi(\Theta_{\varepsilon})$, $(\theta,\varepsilon)\in\mathbb{Z}$, and $\mathbb{Z}$ is closed.
Let $\varepsilon_{0}\geq 0$ and $N_{\theta}$ be a neighborhood of $\hat{\theta}_{\varepsilon_{0}}$. With
$\delta\coloneqq\sup\set{\varepsilon(\theta)}{\theta\in N_{\theta}}>0$
we have $N_{\theta}\cap\Theta$ and $N_{\theta}\cap\Psi(\Theta_{\varepsilon})$ are nonempty for all $\varepsilon\in(0,\varepsilon_{0}+\delta)$.
Finally, the requirements of [71, Prop. 4.4] are satisfied for all $\varepsilon_{0}\in[0,\overline{\varepsilon})$, so $\mu_{\varepsilon}$ is continuous and $\hat{\theta}_{\varepsilon}$ is outer semicontinuous at $\varepsilon=\varepsilon_{0}$.
[…]