### v3

# Steven J. Kuntz and James B. Rawlings

Abstract
This report addresses the maximum likelihood identification of models for offset-free model predictive control, where linear time-invariant models are augmented with (fictitious) uncontrollable integrating modes, called integrating disturbances. The states and disturbances are typically estimated with a Kalman filter. The disturbance estimates effectively provide integral control, so the quality of the disturbance model (and resulting filter) directly influences the control performance. We implement eigenvalue constraints to protect against undesirable filter behavior (unstable or marginally stable modes, high-frequency oscillations). Specifically, we consider the class of linear matrix inequality (LMI) regions for eigenvalue constraints. These LMI regions are open sets by default, so we introduce a barrier function method to create tightened, but closed, eigenvalue constraints. To solve the resulting nonlinear semidefinite program, we approximate it as a nonlinear program using a Cholesky factorization method that exploits known sparsity structures of semidefinite optimization variables and matrix inequalities. The algorithm is applied to real-world data taken from two physical systems: a low-cost benchmark temperature microcontroller suitable for classroom laboratories, and an industrial-scale chemical reactor at Eastman Chemical’s plant in Kingsport, TN.
{twmcc@abskw}
Please cite this work as:-1ex S. J. Kuntz and J. B. Rawlings, “Maximum Likelihood Identification of Linear Models with Integrating Disturbances for Offset-Free Control,” IEEE Trans. Auto. Cont., vol. 70, no. 9, pp. 5675–5689, 2025, doi: 10.1109/TAC.2025.3547607.
1 Introduction
Offset-free model predictive control (MPC) is a widely used advanced control method that combines regulation, estimation, and steady-state optimization problems to track prescribed setpoints [1, 2]. In linear offset-free MPC, a stochastic linear time-invariant (LTI) model is augmented with uncontrollable integrating modes, called integrating disturbances, providing integral action through the estimator (typically a Kalman filter) and allowing offset-free tracking even in the presence of plant-model mismatch and persistent disturbances [3, 4]. We call such a model a linear augmented disturbance model (LADM).
The LADM or its corresponding Kalman filter can either “tuned” by hand or identified automatically from data. Common tuning methods include pole placement [5, 6, 7, 8], covariance matrix selection [9, 10, 11], and filter gain selection [12, 13, 14]. Disturbance models can be identified with autocovariance least squares estimation [15] or maximum likelihood (ML) identification [16, 17, 18, 19].
Tuning of integrating disturbance models can be a time-consuming and ad hoc procedure, requiring simplified parameterizations (e.g., diagonal covariance matrices). In prior work, we have suggested identification as the preferred strategy for acquiring LADMs [16, 17]. In this work, we further develop ML identification because of its desirable statistical properties (consistency, asymptotic efficiency) and ability to handle general model structures and constraints [20, 21].
Design constraints can be included in tuning procedures to avoid undesirable filter behaviors (slow response time, fictitious high frequencies) that are passed to the control performance through the integrating disturbance estimates. Control-relevant design constraints and prior knowledge have sometimes been incorporated into identification problems [22, 23, 24]. However, there are no general approaches to shaping the closed-loop filter behavior in ML identification. To address this gap, we consider ML identification with eigenvalue constraints implemented via the linear matrix inequality (LMI) regions commonly used in robust control [25, 26].
LMI region constraints have been used in subspace identification [27]. However, subspace identification cannot be used for LADM identification as it is not possible to impose the required disturbance model structure. Open-loop stability constraints have been included in the expectation-maximization (EM) algorithm [28], but this formulation is not obviously generalized to filter stability or general LMI region constraints.
While EM is an algorithm for ML, it does not have strong convergence guarantees. While it can be shown that the EM iterates produce, almost surely, an increasing sequence of likelihood values [29, 30], slow convergence at low noise levels has been reported on a range of problems [28, 31, 32, 33, 34, 35]. Interior point, and even gradient methods [35], are therefore preferable to the EM approach.
As originally posed by Chilali and colleagues [25, 26], LMI regions are strict semidefinite matrix inequalities. While Miller and de Callafon [27] used relaxed LMI regions with nonstrict inequalities, as we show in Section 4, the constraint sets are not closed, and thus problematic as optimization constraints. To address this issue, we formulate tightened LMI region constraints that define a closed constraint set. This formulation introduces nonlinear matrix inequalities and semidefinite matrix arguments, making the ML problem a nonlinear semidefinite program (NSDP).
To efficiently convert the NSDP to a nonlinear program (NLP), we generalize the Burer-Monteiro-Zhang (BMZ) method [36, 37], which was originally used to convert sparse semidefinite matrix arguments into vector arguments with minimal dimension. An additional advantage of the BMZ method over standard Cholesky factor substitution is that structural knowledge (e.g., flowsheet or network structure) can be efficiently imposed in the model parameterization. Finally, while this work is primarily motivated by identification of LADMs and offset-free MPC implementations, we remark that any linear Gaussian state-space model can be identified, with eigenvalue constraints, using this approach.
The rest of this report is organized as follows. In Section 2, the ML identification problem is stated. In Section 3 the algorithm is outlined (Algorithm 1). In Section 4, we introduce tightened LMI region constraints show they define closed sets of system matrices (Theorem 2). In Section 5, we present our substitution and elimination scheme for approximating NSDPs as NLPs (Theorems 4 and 5). In Section 6, we apply the algorithm to two real-world applications of offset-free MPC: first, a benchmark temperature microcontroller used for classroom laboratories and prototyping [38], and second, an industrial-scale chemical reactor at Eastman Chemical’s plant in Kingsport, TN [17]. Finally, in Section 7 we discuss broader implications and potential future research.
This report is an extended version of the published work [39], and contains additional review, discussion, and proofs of minor results that were omitted from the journal version due to page limitations. Compared to the journal version, this report contains the following additions:
  • a longer discussion of problem formulations in Section 2;
  • a characterization of models that can be converted to innovation form (Proposition 1, see Appendix A for proof);
  • additional basic LMI regions in Section 3 (see Lemma 1);
  • an explicit counterexample of [27, Thm. 1], which (incorrectly) characterized the eigenvalues of relaxed LMI regions, in 1;
  • a proof of our (correct) characterization of the eigenvalues of relaxed LMI regions (Proposition 2) in Appendix B;
  • a proof of the fact that relaxed LMI regions define neither open nor closed sets of system matrices (Proposition 4(b,c)) in Appendix C;
  • an additional results on solution uniqueness and minimum/infimum equivalences for the BMZ and generalized BMZ method properties (Lemma 2, Proposition 5, and Theorems 3 and 4) in Section 5;
  • and additional remarks throughout.
Notation
For any $z\in\mathbb{C}$, let $\overline{z}$ denote its complex conjugate. Denote the set of $n\times n$ symmetric, positive definite, and positive semidefinite matrices by $\mathbb{S}^{n}$, $\mathbb{S}^{n}_{++}$, and $\mathbb{S}^{n}_{+}$. Denote the set of lower triangular matrices and lower triangular matrices with positive diagonal entries by $\mathbb{L}^{n}$ and $\mathbb{L}^{n}_{++}$. Recall $M\in\real^{n\times n}$ is positive definite if and only if there exists a unique $L\in\mathbb{L}^{n}_{++}$, called the Cholesky factor, such that $M=LL^{\top}$. A $2\times 2$ Hermitian matrix $M=\begin{bmatrix}a&b\\ \overline{b}&c\end{bmatrix}\in\mathbb{H}^{2}$ is positive (semi)definite if and only if $a,c>0$ ($a,c\geq 0$) and $ac>|b|^{2}$ ($ac\geq|b|^{2}$). Denote the matrix direct sum and the Kronecker product by $\oplus$ and $\otimes$, respectively, defined as in [40]. Define the set of eigenvalues of a matrix $A\in\real^{n\times n}$ by $\lambda(A)\subset\mathbb{C}$. The spectral radius and spectral abscissa are defined as $\rho(A)\coloneqq\max_{\lambda\in\lambda(A)}|\lambda|$ and $\alpha(A)\coloneqq\max_{\lambda\in\lambda(A)}\mathrm{Re}(\lambda)$, respectively. We say a matrix $A$ is Schur (Hurwitz) stable if $\rho(A)<1$ ($\alpha(A)<0$). We use $\sim$ as a shorthand for “distributed as” and $\overset{\textnormal{iid}}{\sim}$ as a shorthand for “independent and identically distributed as.” The complement, interior, closure, and boundary of a set $S$ are denoted $S^{c}$, $\textnormal{int}(S)$, $\textnormal{cl}(S)$, and $\partial S$, respectively.
2 Problem statement
We consider stochastic LTI models in innovation form:
$\hat{x}_{k+1}$ $=A(\theta)\hat{x}_{k}+B(\theta)u_{k}+K(\theta)e_{k}$ (1a)
$y_{k}$ $=C(\theta)\hat{x}_{k}+D(\theta)u_{k}+e_{k}$ (1b)
$e_{k}$ $\overset{\textnormal{iid}}{\sim}\mathcal{N}(0,R_{e}(\theta))$ (1c)
where $\hat{x}\in\real^{n}$ are the model states, $u\in\real^{m}$ are the inputs, $y\in\real^{p}$ are the outputs, $e\in\real^{p}$ are the innovation errors, and $\theta\in\Theta$ are the model parameters. The model functions $\mathcal{M}(\cdot)\coloneqq(A(\cdot),B(\cdot),C(\cdot),D(\cdot),\hat{x}_{0}(\cdot),K(\cdot),R_{e}(\cdot))$ are assumed to be known. While the model $\mathcal{M}$ is kept fairly general throughout, it is advantageous to assume the model is identifiable in $\Theta$. Last, for brevity, we often drop the dependence on the parameters $\theta\in\Theta$ and write the model functions as $\mathcal{M}=(A,B,C,D,\hat{x}_{0},K,R_{e})$.
While the subsequent developments apply to any model of the form eq. 1, our main motivation is to identify the LADM,
$\hat{s}_{k+1}$ $=A_{s}(\theta)\hat{s}_{k}+B_{d}(\theta)\hat{d}_{k}+B_{s}(\theta)u_{k}+K_{s}(\theta)e_{k}$ (2a)
$\hat{d}_{k+1}$ $=\hat{d}_{k}+K_{d}(\theta)e_{k}$ (2b)
$y_{k}$ $=C_{s}(\theta)\hat{s}_{k}+C_{d}(\theta)\hat{d}_{k}+D(\theta)u_{k}+e_{k}$ (2c)
$e_{k}$ $\overset{\textnormal{iid}}{\sim}\mathcal{N}(0,R_{e}(\theta))$ (2d)
where $\hat{s}\in\real^{n_{s}}$ denote plant states and $\hat{d}\in\real^{n_{d}}$ denote integrating disturbances. The LADM eq. 2 is clearly a special case of eq. 1 and can be put back into the standard form eq. 1 by consolidating the plant and disturbance states $\hat{x}_{k}\coloneqq\begin{bmatrix}\hat{s}_{k}^{\top}&\hat{d}_{k}^{\top}\end{bmatrix}^{\top}$ and defining
$A$ $\coloneqq\begin{bmatrix}A_{s}&B_{d}\\ 0&I\end{bmatrix}$, $B$ $\coloneqq\begin{bmatrix}B_{s}\\ 0\end{bmatrix}$,
$C$ $\coloneqq\begin{bmatrix}C_{s}&C_{d}\end{bmatrix}$, $K$ $\coloneqq\begin{bmatrix}K_{s}\\ K_{d}\end{bmatrix}$.
Typically the LADM eq. 2 is parameterized with $(A_{s},C_{s})$ in observability canonical form [41], $(B_{d},C_{d})$ fixed, [^1] With $(A_{s},B_{s},C_{s},D)$ fixed, all $(B_{d},C_{d})$ such that eq. 2 is observable are equivalent up to a similarity transform [42]. Thus, $(B_{d},C_{d})$ are chosen by the practitioner to maximize interpretability of the disturbance estimates. $(B_{s},K_{s},K_{d},R_{e})$ fully parameterized, and $(D,\hat{s}_{0},\hat{d}_{0})=(0,0,0)$. Alternatively, a physics-based model may be used for $(A_{s},B_{s},C_{s},D)$.
2.1 Constrained maximum likelihood identification
The ML identification problem is defined as follows:
$\min_{\theta\in\Theta}L_{N}(\theta)\coloneqq\frac{N}{2}\ln\det R_{e}(\theta)+\frac{1}{2}\sum_{k=0}^{N-1}|e_{k}(\theta)|_{[R_{e}(\theta)]^{-1}}^{2}$ (3)
where $N\in\mathbb{I}_{>0}$ is the sample size, $L_{N}$ is the log-likelihood, $e_{k}(\theta)$ are given by eq. 1 [20, p. 557][21, p. 219]. To regularize the estimates with respect to a previous parameter estimate $\overline{\theta}$ or incorporate an available prior distribution $p_{0}(\theta)$, we consider the maximum a posteriori (MAP) identification problem,
$\min_{\theta\in\Theta}L_{N}(\theta)+R_{0}(\theta)$ (4)
where $R_{0}(\theta)\propto-\ln p_{0}(\theta)$ is the regularization term, typically chosen as a distance from $\overline{\theta}$ [43, 44].
For a Gaussian prior or generalized $\ell_{2}$ penalty, we use
$R_{0}(\theta)\coloneqq\frac{1}{2}|\textnormal{vec}(\theta)-\textnormal{vec}(\overline{\theta})|_{V^{-1}}^{2}$ (5)
where $\overline{\theta}\in\Theta$ is the prior estimate, vec is a vectorization operator,  [^2] The vectorization operator may depend on the parameterization, as $\theta$ may contain both a vector portion and a sparse (semidefinite) matrix portion. The vectorization should only preserve the uniquely definednonzero elements of the sparse matrix. and $V\succ 0$ is the prior estimate variance. Such penalties are useful for model updating and re-identification. We typically use the penalty eq. 5 with $V=\rho^{-1}I$. Later on, we transform the parameters $\theta$ into a more convenient space for optimization and find it more convenient to define the prior directly in the transformed space.
In a standard identification setting, the plant takes the form eq. 1 with AKCA-KC stable, the estimates found by eq. 3 are consistent [20], and, with sufficient data, the identified filter is stable. However, the plant is intentionally misspecified by the LADM eq. 2, so this property breaks down in our problem.
Under certain regularity assumptions, we can show consistency with respect to the estimates nearest in relative entropy rate, denoted $\theta^{*}$, taken between the plant and model measurement distributions,
$\theta^{*}\coloneqq\min_{\theta\in\Theta}N^{-1}\mathbb{E}[L_{N}(\theta)]$
where the expectation is taken over the true distribution of measurements $(y_{k})_{k=0}^{N-1}$ [45, 46]. The optimal estimates $\theta^{*}$ may not represent a stable LADM filter, and the identified LADM filter may be unstable even in the large-sample limit. Moreover, there are fundamental limitations to the performance of the Kalman filter for LADMs, as shown in [47]. Specifically, the largest real filter eigenvalue is often bounded from below by the largest real open-loop eigenvalue (as we also find in Section 6). Thus, it is necessary to design $\Theta$ to guarantee high-performance offset-free control.
2.2 Constraints
The constraint set $\Theta$ should capture both estimator design specifications and system knowledge. At a bare minimum, we require nondegeneracy of the innovation errors,
$R_{e}(\theta)\succ 0$ (6)
and stability of the estimator,
$\rho(A(\theta)-K(\theta)C(\theta))<1.$ (7)
Other useful constraints include spectral abscissa bounds,
$\alpha(-\tilde{A}(\theta))<0,$ (8)
and bounds on the argument of the eigenvalues,
$0<|\textrm{Im}(\mu)|/\textrm{Re}(\mu)$ $<s,$ $\forall\mu$ $\in\lambda(\tilde{A}(\theta))$ (9)
where $s>0$, for either the open-loop or estimator stability matrices ($\tilde{A}=A$ or $\tilde{A}=A-KC$), to eliminate artificial high-frequency dynamics that may affect the control performance.
Chemical processes exhibit sparse interactions between units (mass and energy flows), especially for large-scale plants [48, 49]. Sparse parameterization of $(A,B,C,D,K)$ is easily encoded, but that of $R_{e}$ is less obviously accomplished. While, in general, $R_{e}$ is dense even for sparse plants, correlations between distant units are small [50]. Thus, it suffices to consider only nearest-neighbor correlations, e.g.,
$R_{e}=\begin{bmatrix}R_{1,1}&R_{1,2}\\ R_{1,2}^{\top}&R_{2,2}&\ddots\\ &\ddots&\ddots&R_{N_{u}-1,N_{u}}\\ &&R_{N_{u}-1,N_{u}}^{\top}&R_{N_{u},N_{u}}\end{bmatrix}$ (10)
where $R_{i,j}\in\real^{p_{u}\times p_{u}}$ is the covariance between the innovations of the $i$-th and $j$-th process unit innovations. In eq. 10, the sparse formulation introduces just $O(N_{u}p_{u}^{2})$ variables compared to $O(N_{u}^{2}p_{u}^{2})$ variables for the dense formulation. Such constraints can be applied to the ML identification of any networked system with a time-invariant topology, as in [51].
2.3 Other parameterizations
The remainder of this section presents some other formulations of the ML identification problem. While we do not consider these formulations explicitly in our algorithm formulation (Section 3) or case studies (Section 6), the methods are readily generalized to these formulations.
2.3.1 Time-varying Kalman filter formulations
More generally, we could consider models of the following form:
$x_{k+1}$ $=A(\theta)x_{k}+B(\theta)u_{k}+w_{k}$ (11a)
$y_{k}$ $=C(\theta)x_{k}+D(\theta)u_{k}+v_{k}$ (11b)
$x_{0}$ $\sim\mathcal{N}(\hat{x}_{0}(\theta),\hat{P}_{0}(\theta))$ (11c)
$\begin{bmatrix}w_{k}\\ v_{k}\end{bmatrix}$ $\overset{\textnormal{iid}}{\sim}\mathcal{N}(0,S(\theta))$ (11d)
where $\mathcal{M}\coloneqq(A,B,C,D,\hat{x}_{0},\hat{P}_{0},S)$ are the model functions and $w\in\real^{n}$ and $v\in\real^{p}$ are the process and measurement noises. The noise covariance matrix $S(\theta)$ may be partitioned as
$S(\theta)=\begin{bmatrix}Q_{w}(\theta)&S_{wv}(\theta)\\ [S_{wv}(\theta)]^{\top}&R_{v}(\theta)\end{bmatrix}$ (12)
where $Q_{w}(\theta)\in\mathbb{S}^{n}_{+}$ is the process noise covariance, $S_{wv}(\theta)$ is the cross-covariance, and $R_{v}(\theta)\in\mathbb{S}^{p}_{+}$ is the measurement noise covariance. Throughout, we impose the stronger requirement $R_{v}(\theta)\succ 0$ on the measurement noise covariance.
For the model eq. 11, the ML problem is defined as
$\min_{\theta\in\Theta}L_{N}(\theta)\coloneqq\frac{1}{2}\sum_{k=0}^{N-1}\ln\det\mathcal{R}_{k}(\theta)+|e_{k}(\theta)|_{[\mathcal{R}_{k}(\theta)]^{-1}}^{2}$ (13)
where the $e_{k}$ and $\mathcal{R}_{k}$ are defined by the Kalman filtering equations
$\hat{x}_{k+1}$ $=A\hat{x}_{k}+Bu_{k}+\mathcal{K}_{k}e_{k}$ (14a)
$y_{k}$ $=C\hat{x}_{k}+Du_{k}+e_{k}$ (14b)
$e_{k}$ $\sim\mathcal{N}(0,\mathcal{R}_{k})\quad\textnormal{(indep.)}$ (14c)
where
$\hat{P}_{k+1}$ $\coloneqq A\hat{P}_{k}A^{\top}+Q_{w}-\mathcal{K}_{k}\mathcal{R}_{k}\mathcal{K}_{k}^{\top}$ (14d)
$\mathcal{K}_{k}$ $\coloneqq(A\hat{P}_{k}C^{\top}+S_{wv})\mathcal{R}_{k}^{-1}$ (14e)
$\mathcal{R}_{k}$ $\coloneqq C\hat{P}_{k}C^{\top}+R_{v}.$ (14f)
We remark that $R_{v}\succ 0$ suffices to guarantee the innovations are uniformly nondegenerate, i.e., $\mathcal{R}_{k}\succ 0$. However, stability of the filter is more difficult to guarantee as the early iterates $A-\mathcal{K}_{k}C$ may not be stable, even though the overall filter is stable, or vice versa. Instead, it is necessary to check that a stabilizing solution to the Riccati equation exists, which we elaborate on in the next formulation.
2.3.2 Time-invariant Kalman filter formulations
In most situations, the state error covariance matrix converges exponentially fast to a steady-state solution $\hat{P}_{k}\rightarrow\hat{P}$, so it suffices to consider the original steady-state filter model eq. 1. In terms of the model eq. 11, the steady-state filter takes the form $K\coloneqq(A\hat{P}C^{\top}+S_{wv})R_{e}^{-1}$ and $R_{e}\coloneqq C\hat{P}C^{\top}+R_{v}$, where $\hat{P}$ is the unique, stabilizing solution to the discrete algebraic Riccati equation (DARE),
$\hat{P}=A\hat{P}A^{\top}+Q_{w}-(A\hat{P}C^{\top}+S_{wv})(C\hat{P}C^{\top}+R_{v})^{-1}(A\hat{P}C^{\top}+S_{wv})^{\top}.$ (15)
Recall a solution to the DARE eq. 15 is stabilizing if the resulting $A_{K}\coloneqq A-KC$ is stable.
Convergence of $\hat{P}_{k}$ to $\hat{P}$ is equivalent to the solution to the DARE eq. 15 being unique and stabilizing. We generally assume such a solution exists, but for completeness, we state the following proposition, adapted from [52, Thm. 18(iii)] (see Appendix A for proof).
Proposition 1.
Assume $R_{v}\succ 0$ and consider the full rank factorization
$\begin{bmatrix}Q_{w}&S_{wv}\\ S_{wv}^{\top}&R_{v}\end{bmatrix}=\begin{bmatrix}\tilde{B}\\ \tilde{D}\end{bmatrix}\begin{bmatrix}\tilde{B}^{\top}&\tilde{D}^{\top}\end{bmatrix}$
Then the following statements are equivalent:
  1. 1.
    The DARE eq. 15 has a unique, stabilizing solution $\hat{P}\succeq 0$.
  2. 2.
    The error covariance converges exponentially fast $\hat{P}_{k}\rightarrow\hat{P}$ for any $\hat{P}_{0}\succeq 0$.
  3. 3.
[…]
Remark 1.
The hypothesis of Proposition 1 holds if we constrain $A$ to be stable or $(A,C)$ to be observable.
Remark 2.
The cross-covariance $S_{wv}$ complicates the filter stability analysis. With $S_{wv}=0$, it would suffice to assume $(A,C)$ detectable and $(A,Q_{w})$ stabilizable. With nonzero $S_{wv}$, however, a more elaborate stabilizability condition is needed. [52, Thm. 18] considers the regulation problem with a cross-weighting term and semidefinite input weights. Proposition 1 specializes this result to the filter problem with positive definite $R_{v}$.
Remark 3.
While $R_{e}(\theta)$ and $K(\theta)$ could be defined via $\hat{P}(\theta)$, taken as the function that returns solutions to the DARE eq. 15 and therefore enforcing filters stability, it is more convenient to directly parameterize these matrices as in eq. 1.
2.3.3 Minimum determinant formulation
Suppose in the model eq. 1, that $R_{e}$ is parameterized fully, and separately from the other terms, i.e.,
$\mathcal{M}(\tilde{\theta},R_{e})=\left(A(\tilde{\theta}),B(\tilde{\theta}),C(\tilde{\theta}),D(\tilde{\theta}),\hat{x}_{0}(\tilde{\theta}),K(\tilde{\theta}),R_{e}\right).$
Moreover, assume $R_{e}$ is constrained separately as well, i.e.,
$\Theta=\tilde{\Theta}\times\mathbb{S}^{p}_{++}.$
Then we can always solve eq. 3 stagewise, first in $R_{e}$, and then in the remaining variables $\tilde{\theta}$, i.e.,
$\min_{\tilde{\theta}\in\tilde{\Theta}}\min_{R_{e}\in\mathbb{S}^{p}_{++}}L_{N}(\tilde{\theta},R_{e})=\frac{N}{2}\ln\det R_{e}+\frac{1}{2}\sum_{k=0}^{N-1}|e_{k}(\tilde{\theta})|_{R_{e}^{-1}}^{2}.$
Solving the inner problem gives the solution
[…]
where we use the fact that $e_{k}$ is only dependent on $\tilde{\theta}$, and we assume $\hat{R}_{e}(\tilde{\theta})\succ 0$ for all $\tilde{\theta}\in\tilde{\Theta}$. The outer problem can be written
$\min_{\tilde{\theta}\in\tilde{\Theta}}L_{N}(\tilde{\theta},\hat{R}_{e}(\tilde{\theta}))=\frac{N}{2}\det\hat{R}_{e}(\tilde{\theta})+\frac{p}{2}$ (16)
which is the minimum determinant problem.
The problem eq. 16 is of relevance because it avoids posing eq. 3 as a NSDP. It has been used both in the early ML identification literature [53, 54, 55] and in recent works [56, 57, 58]. None of these works consider filter stability constraints. To the best of our knowledge, only [28] consider the ML problem eq. 13 with stability constraints, but they consider open-loop stability (i.e., $\rho(A)<1$) and use the EM algorithm.
Remark 4.
For real-world data, $\det\hat{R}_{e}(\tilde{\theta},\hat{R}_{e}(\tilde{\theta}))=0$ is not attainable because that would imply some direction of $y_{k}$ were perfectly modeled. Therefore, $\hat{R}_{e}(\tilde{\theta})\succ 0$ for all $\tilde{\theta}\in\tilde{\Theta}$ is a reasonable assumption.
3 Algorithm outline
3.1 Constraint set formulation
More generally, we seek to (i) impose eigenvalue constraints on any model function $\tilde{A}(\theta)$ and (ii) impose a sparsity structure on any semidefinite model function $\tilde{Q}(\theta)$.
3.1.1 Eigenvalue constraints
First, we define a LMI region.
Definition 1.
We call $\mathcal{D}\subseteq\mathbb{C}$ an LMI region if
[…]
for some generating matrices $(M_{0},M_{1})\in\mathbb{S}^{m}\times\real^{m\times m}$. We call $f_{\mathcal{D}}:\mathbb{C}\rightarrow\mathbb{S}^{m}$ the characteristic function of $\mathcal{D}$.
The following lemma defines the four basic LMI regions: shifted half-planes, circles centered on the real axis, conic sections, and horizontal bands. For a general discussion of LMI regions properties, see [25, 59].
Lemma 1.
For each $s,x_{0}\in\real$, the subsets
[…]
are LMI regions with characteristic functions
$f_{\mathcal{D}_{1}(x_{0})}(z)$ $\coloneqq-2x_{0}+z+\overline{z}$
$f_{\mathcal{D}_{2}(s,x_{0})}(z)$ $\coloneqq\begin{bmatrix}s&-x_{0}\\ -x_{0}&s\end{bmatrix}+\begin{bmatrix}0&1\\ 0&0\end{bmatrix}z+\begin{bmatrix}0&0\\ 1&0\end{bmatrix}\overline{z}$
$f_{\mathcal{D}_{3}(s,x_{0})}(z)$ $\coloneqq-2sx_{0}I_{2}+\begin{bmatrix}s&1\\ -1&s\end{bmatrix}z+\begin{bmatrix}s&-1\\ 1&s\end{bmatrix}\overline{z}$
$f_{\mathcal{D}_{4}(s)}(z)$ $\coloneqq-2sI_{2}+\begin{bmatrix}0&1\\ -1&0\end{bmatrix}z+\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}\overline{z}.$
Proof.
The first identity follows from the formula $2\textnormal{Re}(z)=z+\overline{z}$. For the second identity, we have $f_{\mathcal{D}_{2}(s,x_{s})}(z)=\begin{bmatrix}s&z-x_{0}\\ \overline{z}-x_{0}&s\end{bmatrix}\succ 0$ if and only if $s>0$ and $s^{2}>|z-x_{0}|^{2}$, or equivalently, $|z-x_{0}|<s$. For the third identity, we have $f_{\mathcal{D}_{3}(s,x_{0})}(z)=\begin{bmatrix}2s(\textnormal{Re}(z)-x_{0})&2\iota\textnormal{Im}(z)\\ -2\iota\textnormal{Im}(z)&2s(\textnormal{Re}(z)-x_{0})\end{bmatrix}\succ 0$ if and only if $2s(\textnormal{Re}(z)-x_{0})>0$ and $4s^{2}(\textnormal{Re}(z)-x_{0})^{2}>4|\textnormal{Im}(z)|^{2}$, or equivalently, $|\textnormal{Im}(z)|<s(\textnormal{Re}(z)-x_{0})$. For the fourth identity, we have $f_{\mathcal{D}_{4}(s)}(z)=\begin{bmatrix}2s&2\iota\textnormal{Im}(z)\\ -2\iota\textnormal{Im}(z)&2s\end{bmatrix}\succ 0$ if and only if $2s>0$ and $4s^{2}>4|\textnormal{Im}(z)|^{2}$, or equivalently, $|\textnormal{Im}(z)|<s$. ∎
Remark 5.
For continuous-time systems, $\mathcal{D}_{1}(\alpha)$ corresponds to a minimum decay rate of $\alpha>0$, $\mathcal{D}_{3}(-\tan(\omega),0)$ corresponds to a minimum damping ratio $\cos(\omega)$, and $\mathcal{D}_{2}(r,0)\cap\mathcal{D}_{3}(-\tan(\omega),0)$ implies to a maximum undamped natural frequency $r\sin(\omega)$, where $\alpha,r>0$ and $\omega\in[0,\pi/2]$ [25]. For discrete-time systems, $\mathcal{D}_{2}(r,0)$ corresponds to a minimum decay rate of $\ln r/\Delta$, and $\mathcal{D}_{2}(r,0)\cap\mathcal{D}_{3}(\tan(\omega),0)$ implies a minimum damping ratio $\cos(\tan^{-1}(\omega/\ln r))$ and maximum natural frequency $\sqrt{\ln(r)^{2}+\omega^{2}}/\Delta$, where $r\in(0,1)$, $\omega\in[0,\pi/2)$, and $\Delta$ is the sample time.
Remark 6.
For any LMI region $\mathcal{D}$ (including those in Lemma 1), the set $\mathcal{D}$ is convex, open, and symmetric about the real axis. The intersection of two LMI regions $\mathcal{D}\coloneqq\mathcal{D}_{1}\cap\mathcal{D}_{2}$ is an LMI region with the characteristic function $f_{\mathcal{D}}(z)=f_{\mathcal{D}_{1}}(z)\oplus f_{\mathcal{D}_{2}}(z)$. By this property, we can construct any convex polyhedron that is symmetric about the real axis by intersecting left and right half-planes, horizontal strips, and conic sections. Moreover, since any convex region can be approximated, to any desired accuracy, by a convex polyhedron, the set of LMI regions is dense in the space of convex subsets of $\mathbb{C}$ that are symmetric about the real axis. An LMI region $\mathcal{D}$ with characteristic function $f_{\mathcal{D}}$ also has characteristic function $Mf_{\mathcal{D}}(\cdot)M^{\top}$ for any nonsingular $M\in\real^{m\times m}$. For an in-depth discussion of LMI region geometry and other properties, see [59].
In [25], it is shown that a matrix $\tilde{A}\in\real^{\tilde{n}\times\tilde{n}}$ has eigenvalues in a LMI region $\mathcal{D}$ if and only if the following system of matrix inequalities is feasible:
$M_{\mathcal{D}}(\tilde{A},P)$ $\succ 0,$ $P$ $\succ 0$ (17)
where the matrix characteristic function $M_{\mathcal{D}}:\real^{\tilde{n}\times\tilde{n}}\times\mathbb{S}^{\tilde{n}}\rightarrow\mathbb{S}^{\tilde{n}\tilde{m}}$ of $\mathcal{D}$ is defined by
$M_{\mathcal{D}}(\tilde{A},P)\coloneqq M_{0}\otimes P+M_{1}\otimes(\tilde{A}P)+M_{1}^{\top}\otimes(\tilde{A}P)^{\top}.$ (18)
[…]
$\begin{bmatrix}P&(A-KC)P\\ P(A-KC)^{\top}&P\end{bmatrix}$ $\succ 0,$ $P$ $\succ 0$
[…]
The system of matrix inequalities eq. 17 contains only strict inequalities, but we can “tighten” them as follows:
$M_{\mathcal{D}}(\tilde{A},P)$ $\succeq M,$ $P$ $\succeq 0,$ $\textnormal{tr}(VP)$ $\leq\varepsilon^{-1}$ (19)
where $\varepsilon>0$, $V\in\mathbb{S}^{\tilde{n}}_{++}$, and $M\in\mathbb{S}^{\tilde{n}\tilde{m}}_{++}$. The set of $\tilde{A}\in\real^{\tilde{n}\times\tilde{n}}$ for which eq. 19 is feasible defines a closed set for which $\lambda(\tilde{A})\subseteq\mathcal{D}$. In Section 4, we show this fact and other properties of the constraint eq. 19.
3.1.2 Sparsity structure
[…]
For each $\mathcal{I}\subseteq\mathcal{L}^{n}$, define the sets
[…]
Finally, let $\textnormal{vecs}_{\mathcal{I}}:\mathbb{S}^{n}\rightarrow\real^{|\mathcal{I}|}$ denote the operator that vectorizes the $|\mathcal{I}|$ entries of the argument corresponding to the index set $\mathcal{I}$.
3.1.3 Constraint definition
To combine the LMI region and sparsity constraints, we partition the parameter into vector and sparse symmetric matrix parts, i.e., $\theta=(\beta,\Sigma)$, and define the constraint set $\Theta$ by
[…]
where $\mathcal{D}^{n_{\Sigma}}\subseteq\mathcal{I}_{\Sigma}\subseteq\mathcal{L}^{n_{\Sigma}}$, $\mathcal{D}^{n_{\mathcal{A}}}\subseteq\mathcal{I}_{\mathcal{A}}\subseteq\mathcal{L}^{n_{\mathcal{A}}}$, $g:\real^{n_{\beta}}\times\mathbb{S}^{n_{\Sigma}}\rightarrow\real^{n_{g}}$, $h:\real^{n_{\beta}}\times\mathbb{S}^{n_{\Sigma}}\rightarrow\real^{n_{h}}$, $H:\real^{n_{\beta}}\rightarrow\mathbb{S}^{n_{\Sigma}}$, and $\mathcal{A}:\real^{n_{\beta}}\times\mathbb{S}^{n_{\Sigma}}\rightarrow\mathbb{S}^{n_{\mathcal{A}}}[\mathcal{I}_{\mathcal{A}}]$. The purpose of the partition $\theta=(\beta,\Sigma)$ is to clearly delineate the sparse semidefinite matrix argument $\Sigma$ from the remaining parameters $\beta$. The index set $\mathcal{I}_{\Sigma}$ defines the sparsity pattern of $\Sigma$ and $H$, and the index set $\mathcal{I}_{\mathcal{A}}$ defines the sparsity pattern of $\mathcal{A}$.
Remark 7.
1 rules out direct use strict inequalities, e.g., $R_{e}(\theta)\succ 0$ or $R_{v}(\theta)\succ 0$. To satisfy nondegeneracy requirements, we use the closed constraint $R_{e}(\theta)\succeq\delta I_{p}$ with a small back-off $\delta>0$.
Remark 8.
Typically, the index set $\mathcal{I}_{\Sigma}$ encodes block diagonal structures, e.g., for the model eq. 11, $\Sigma=\hat{P}_{0}\oplus Q_{w}\oplus R_{v}\in\mathbb{S}^{2n+p}[\mathcal{I}_{\Sigma}]$ where $\mathcal{I}_{\Sigma}\coloneqq\mathcal{L}^{n}\oplus\mathcal{L}^{n}\oplus\mathcal{L}^{p}$. However, more general structures (e.g., eq. 10) can be stated. For the time-varying formulation eq. 13, we may further restrict $Q_{w}$ and $R_{v}$ to take block tridiagonal and diagonal structures, e.g.,
$Q_{w}$ $=\begin{bmatrix}Q_{1,1}&Q_{1,2}\\ Q_{1,2}^{\top}&Q_{2,2}&\ddots\\ &\ddots&\ddots&Q_{\tilde{n}-1,\tilde{n}}\\ &&Q_{\tilde{n}-1,\tilde{n}}^{\top}&Q_{\tilde{n},\tilde{n}}\end{bmatrix}$, $R_{v}$ $=R_{1}\oplus\ldots\oplus R_{\tilde{n}}$
that arise in sequentially interconnected processes such as chemical plants. Adding a $Q_{1,\tilde{n}}$ block can account for an overall recycle loop. Note that if we parameterize the block tridiagonal $Q_{w}$ via a sparse shaping matrix (i.e., $Q_{w}=G_{w}G_{w}^{\top}$), then there are more parameters than if the sparsity of $Q_{w}$ is known.
Remark 9.
As alluded to in Section 2, the Riccati equation solution has a dense solution, but the entries far from the core sparsity pattern decay rapidly. Thus, we can approximate an eigenvalue constraint, e.g., $P-APA^{\top}\succ 0$, as a function that maps to the same sparsity pattern as AA [50, 61, 62, 63].
3.2 Cholesky factorization and elimination
At this juncture, the ML and MAP problems eqs. 3 and 4 with the constraints eq. 20 are in standard NSDP form and can be solved with any dedicated NSDP solver, e.g., [64, 65]. However, such solvers are neither as widely available nor as well-understood as NLP solvers such as IPOPT [66].
The Burer-Monteiro-Zhang (BMZ) method is a Cholesky factorization-based substitution and elimination algorithm that can convert certain NSDPs to NLPs [36, 37]. In Section 5, we consider a generalization of this algorithm to (approximately) transform a given NSDP into a NLP while only introducing $|\mathcal{I}_{\mathcal{A}}|$ new variables. This generalization requires the following assumption.
Assumption 1.
The model functions $\mathcal{M}$ are twice differentiable and the constraint functions $\mathcal{C}\coloneqq(g,h,H,\mathcal{A})$ are differentiable. Moreover, $\textnormal{cl}(\Theta_{++})=\Theta$ where
[…]
In Section 5, we construct functions
$\mathcal{T}$ $:\real^{n_{\beta}}\times\mathbb{L}^{n_{\Sigma}}_{++}[\mathcal{I}_{\Sigma}]\times\mathbb{L}^{n_{\mathcal{A}}}_{++}[\mathcal{I}_{\mathcal{A}}]\rightarrow\real^{n_{\beta}}\times\mathbb{S}^{n_{\Sigma}}[\mathcal{I}_{\Sigma}]$
$\mathcal{A}_{\mathcal{T}}$ $:\real^{n_{\beta}}\times\mathbb{L}^{n_{\Sigma}}_{++}[\mathcal{I}_{\Sigma}]\times\mathbb{L}^{n_{\mathcal{A}}}_{++}[\mathcal{I}_{\mathcal{A}}]\rightarrow\mathbb{S}^{n_{\mathcal{A}}}_{++}[\mathcal{I}_{\mathcal{A}}]$
and define transformed constraint functions
$g_{\mathcal{T}}(\phi)$ $\coloneqq\begin{bmatrix}g(\mathcal{T}(\phi))\\ \textnormal{vecs}_{\mathcal{I}_{\mathcal{A}}}(\mathcal{A}(\mathcal{T}(\phi))-\mathcal{A}_{\mathcal{T}}(\phi))\end{bmatrix}$ (22a)
$h_{\mathcal{T}}(\phi)$ $\coloneqq h(\mathcal{T}(\phi))$ (22b)
and a transformed constraint set
[…]
such that, under 1, $\mathcal{T}$ is an invertible map from $\Phi$ to $\Theta_{++}$. Finally, to eliminate the strict inequalities on the diagonal entries of $(L_{\Sigma},L_{\mathcal{A}})\in\mathbb{L}^{n_{\Sigma}}_{++}[\mathcal{I}_{\Sigma}]\times\mathbb{L}^{n_{\mathcal{A}}}_{++}[\mathcal{I}_{\mathcal{A}}]$, we introduce a fixed lower bound $\varepsilon>0$ on the diagonal entries,
[…]
where we have defined, for any $\varepsilon>0$ and $\mathcal{I}\subseteq\mathcal{L}^{n}$,
[…]
We define the approximate transformed problem as
$\min_{\phi\in\Phi_{\varepsilon}}L_{N}(\mathcal{T}(\phi))+R_{0}(\mathcal{T}(\phi)).$ (25)
[…]
3.3 Algorithm summary
Algorithm 1 provides an example of our approach towards solving the identification problem eq. 4 with eigenvalue constraints and the Cholesky factor-based substitution and elimination scheme.
Algorithm 1 Identification of an innovation form model eq. 1 with eigenvalue constraints and the Cholesky factor-based substitution and elimination scheme.
0: Model functions $\mathcal{M}=(A,B,C,D,\hat{x}_{0},K,R_{e})$, regularization term $R_{0}$, initial parameters $\theta_{0}=(\beta,\Sigma_{0})$ constraint functions $\mathcal{C}_{0}=(g,h_{0},H_{0},\mathcal{A}_{0})$ and sparsity patterns $(\mathcal{I}_{\Sigma_{0}},\mathcal{I}_{\mathcal{A}_{0}})$, a series of LMI region constraints $(\mathcal{D}_{i},\tilde{A}_{i}(\cdot))_{i=1}^{n_{c}}$ with small tightening constants $\varepsilon_{i}>0,i\in\mathbb{I}_{1:n_{c}}$, and a small constant $\varepsilon>0$ (for eq. 24).
1: For each $i\in\mathbb{I}_{1:n_{c}}$, let $M_{\mathcal{D}_{i}}:\real^{n_{i}\times n_{i}}\times\mathbb{S}^{n_{i}}\rightarrow\mathbb{S}^{n_{i}m_{i}}$ denote the matrix characteristic function for $\mathcal{D}_{i}$.
2: Extend the parameters $\Sigma\coloneqq\Sigma_{0}\oplus\left(\bigoplus_{i=1}^{n_{c}}P_{i}\right)$ and $\theta\coloneqq(\beta,\Sigma)$ with $P_{i}\in\mathbb{S}^{n_{i}}$.
3: Extend the constraint functions
$h(\theta)$ $\coloneqq\begin{bmatrix}h_{0}(\theta_{0})\\ \textnormal{tr}(V_{1}P_{1})-\varepsilon_{1}^{-1}\\ \ldots\\ \textnormal{tr}(V_{n_{c}}P_{n_{c}})-\varepsilon_{n_{c}}^{-1}\end{bmatrix}$,
$H(\beta)$ $\coloneqq H_{0}(\beta)\oplus\textstyle\left(\bigoplus_{i=1}^{n_{c}}0_{n_{i}\times n_{i}}\right),$
$\mathcal{A}(\theta)$ $\coloneqq\mathcal{A}_{0}(\theta_{0})\oplus\textstyle\left(\bigoplus_{i=1}^{n_{c}}M_{\mathcal{D}_{i}}(\tilde{A}_{i}(\theta_{0}),P_{i})-\varepsilon_{i}I\right).$
4: Extend the index sets $\mathcal{I}_{\Sigma}\coloneqq\mathcal{I}_{\Sigma_{0}}\oplus\textstyle\left(\bigoplus_{i=1}^{n_{c}}\mathcal{L}^{n_{i}}\right)$ and $\mathcal{I}_{\mathcal{A}}\coloneqq\mathcal{I}_{\mathcal{A}_{0}}\oplus\textstyle\left(\bigoplus_{i=1}^{n_{c}}\mathcal{L}^{n_{i}m_{i}}\right)$.
5: Form the functions $\mathcal{T}$, $\mathcal{T}^{-1}$, and $\tilde{\mathcal{A}}$ as in Section 5.
6: Form the transformed constraint functions eq. 22.
7: Solve eqs. 25 and 24, and let $\hat{\phi}$ denote the solution.
[…]
4 Eigenvalue constraints
In this section, we elaborate on the LMI region constraints previewed in Section 3. Throughout, assume the LMI region $\mathcal{D}$ is nonempty, not equal to $\mathbb{C}$, and its characteristic function $f_{\mathcal{D}}$ and generating matrices $(M_{0},M_{1})$ are fixed. Our goal in this section is to define, using only matrix inequalities, a closed set of matrices $A\in\real^{n\times n}$ such that $\lambda(A)\subseteq\mathcal{D}$. For this section, the matrix $A\in\real^{n\times n}$ need not have any relation to the model function in eq. 1, and can in fact be any square matrix of any dimension (e.g., the filter stability matrix AKCA-KC, the plant stability matrix $A_{s}$ from eq. 2, or any submatrix thereof). Throughout this section, we assume the matrix characteristic function $M_{\mathcal{D}}$ is fixed.
4.1 LMI region constraints
Originally, Chilali and Gahinet [25] proved the following theorem relating the eigenvalues of $A\in\real^{n\times n}$ to feasibility of a system of matrix inequalities.
Theorem 1 ([25, Thm. 2.2]).
For any $A\in\real^{n\times n}$, we have $\lambda(A)\subseteq\mathcal{D}$ if and only if
$M_{\mathcal{D}}(A,P)$ $\succ 0,$ $P$ $\succ 0.$ (26)
holds for some $P\in\mathbb{S}^{n}$.
Ultimately, we seek matrix inequalities that define a closed set of constraints. Due to the strictness of the inequalities eq. 26, it is unlikely that Theorem 1 achieves this goal.
4.2 Relaxed constraints
In [27], the following relaxation of eq. 26 was considered,
$M_{\mathcal{D}}(A,P)$ $\succeq 0,$ $P$ $\succ 0.$ (27)
Since $M_{\mathcal{D}}(A,P)$ is linear in $P$, feasibility of eq. 28 is equivalent to feasibility of
$M_{\mathcal{D}}(A,P)$ $\succeq 0,$ $P$ $\succeq P_{0}$ (28)
for some fixed $P_{0}\in\mathbb{S}^{n}_{++}$. [^3] For any $P_{0}\succ 0$ and $P$ satisfying eq. 27, define the scaling factor $\gamma\coloneqq\|P_{0}\|_{2}\|P^{-1}\|_{2}$ and a rescaled solution $P^{*}\coloneqq\gamma P$. Then $P^{*}\succeq P_{0}$ and $M_{\mathcal{D}}(A,P^{*})=\gamma M_{\mathcal{D}}(A,P)\succeq 0$.
An attempt was made in [27, Thm. 1] to characterize the eigenvalues of matrices $A\in\real^{n\times n}$ for which eq. 27 is feasible, but this theorem does not correctly treat eigenvalues on the LMI region’s boundary $\partial\mathcal{D}$. We restate [27, Thm. 1] below as a conjecture and disprove it with a simple counterexample.
Conjecture 1 ([27, Thm. 1]).
The matrix $A\in\real^{n\times n}$ satisfies $\lambda(A)\subset\textnormal{cl}(\mathcal{D})$ if and only if eq. 27 holds for some $P\in\mathbb{S}^{n}$.
Counterexample.
Let $\mathcal{D}$ be the left half-plane, consider the Jordan block $A=\begin{bmatrix}0&1\\ 0&0\end{bmatrix}$, and suppose $P=\begin{bmatrix}p_{11}&p_{12}\\ p_{12}&p_{22}\end{bmatrix}\in\mathbb{S}^{2}$ such that eq. 27 holds. Then $\lambda(A)\subset\textnormal{cl}(\mathcal{D})$ and
$0\preceq M_{\mathcal{D}}(A,P)=-\begin{bmatrix}2p_{12}&p_{22}\\ p_{22}&0\end{bmatrix}$
which implies $p_{12}=p_{22}=0$, a contradiction of eq. 27. ∎
The correction to 1 requires a more careful treatment of eigenvalues lying on the the LMI region’s boundary $\partial\mathcal{D}$. Specifically, we show in the following proposition that feasibility of eq. 27 for a given $A\in\real^{n\times n}$ is equivalent to the eigenvalues of $A$ being in $\textnormal{cl}(\mathcal{D})$, with all non-simple eigenvalues lying in $\mathcal{D}$ (see Appendix B for proof).
Proposition 2.
The matrix $A\in\real^{n\times n}$ satisfies $\lambda(A)\subseteq\textnormal{cl}(\mathcal{D})$ and $\lambda\in\mathcal{D}$ for all non-simple eigenvalues $\lambda\in\lambda(A)$ if and only if eq. 27 holds for some $P\in\mathbb{S}^{n}$.
4.3 Tightened constraints
Instead of the “relaxed” constraints eq. 27, we consider “tightened” constraints of the form
$M_{\mathcal{D}}(A,P)$ $\succeq M,$ $P$ $\succeq 0,$ $\textnormal{tr}(VP)$ $\leq\varepsilon^{-1}$ (29)
where $M\in\mathbb{S}^{nm}_{+}$ and $V\in\mathbb{S}^{n}_{++}$ are fixed and chosen in a way that eq. 29 implies eq. 26. While we allow $M$ to be semidefinite, [^4] For some LMI regions, $M\succeq 0$ is advantageous. For example, we can always take $M\coloneqq\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\otimes Q$ with $Q\succ 0$ for circular LMI regions. Then we can reduce the constraint dimension by taking the Schur complement. in the following proposition, we show $M\succ 0$ always suffices.
Proposition 3.
Suppose $M\in\mathbb{S}^{nm}_{++}$ and $V\in\mathbb{S}^{n}_{++}$. Then eq. 29 implies eq. 26 for all $A\in\real^{n\times n}$ and $\varepsilon>0$.
Proof.
With $M\succ 0$ and eq. 29, we automatically have $M_{\mathcal{D}}(A,P)\succ 0$. It remains to show eq. 29 implies $P\succ 0$. For contradiction suppose eq. 29 and $M\succ 0$, but $P\not\succ 0$. Then there exists a nonzero $v\in\real^{n}$ such that $Pv=0$, and
$(I_{m}\otimes v)^{\top}M_{\mathcal{D}}(A,P)(I_{m}\otimes v)=M_{0}\otimes(v^{\top}Pv)+M_{1}\otimes(v^{\top}APv)+M_{1}^{\top}\otimes(v^{\top}PA^{\top}v)=0$
a contradiction of the assumption $M_{\mathcal{D}}(A,P)\succeq M\succ 0$. ∎
Remark 10.
The tightened constraint eq. 29 was inspired by a similar set of constraints was introduced by Diehl and colleagues [67] to “smooth” the spectral radius. Specifically, feasibility of the nonlinear system
$s^{2}P-APA^{\top}$ $=W,$ $P$ $\succeq 0,$ $\textnormal{tr}(VP)$ $\leq\varepsilon^{-1}$ (30)
implies $\rho(A)<s$ where $W,V\in\mathbb{S}^{n}_{++}$ and $s,\varepsilon>0$ are fixed [67, Thms. 5.4, 5.6]. Similarly, the spectral abscissa was “smoothed” in [68, Thms. 2.5, 2.6], and it is straightforward to generalize [67, Thms. 5.4, 5.6] to show feasibility of
$(A-sI)P+P(A-sI)^{\top}$ $=-W,$ $P$ $\succeq 0,$ $\textnormal{tr}(VP)$ $\leq\varepsilon^{-1}$ (31)
implies $\alpha(A)<s$ where $W,V\in\mathbb{S}^{n}_{++}$, $s\in\real$, and $\varepsilon>0$ are fixed. The authors do not discuss LMI regions and the results are not obviously generalizable to them.
4.4 Constraint topology
Consider the constraint sets,
[…]
The following proposition characterizes the topology of $\mathbb{A}_{\mathcal{D}}^{n}$ and $\tilde{\mathbb{A}}_{\mathcal{D}}^{n}$ (see Appendix C for proof).
Proposition 4.
The following holds.
  1. (a)
    $\mathbb{A}_{\mathcal{D}}^{n}$ is open.
  2. (b)
    $\tilde{\mathbb{A}}_{\mathcal{D}}^{n}$ is not open if (i) $n\geq 2$ or (ii) $\partial\mathcal{D}\cap\real$ is nonempty.
  3. (c)
    $\tilde{\mathbb{A}}_{\mathcal{D}}^{n}$ is not closed if (i) $n\geq 4$ or (ii) $\partial\mathcal{D}\cap\real$ is nonempty and $n\geq 2$.
  4. (d)
[…]
Proposition 4 reveals a weakness of the relaxed constraints eqs. 27 and 28. Since $\tilde{\mathbb{A}}_{\mathcal{D}}^{n}$ is not closed, any feasible path towards a matrix $A\in\textnormal{cl}(\mathbb{A}_{\mathcal{D}}^{n})\setminus\tilde{\mathbb{A}}_{\mathcal{D}}^{n}$ has no feasible limiting $P$. In fact, $P$ will grow unbounded along the path of iterates.
To analyze the topology of $\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon)$, we take a barrier function approach. Consider the parameterized linear SDP,
$\phi_{\mathcal{D}}(A)\coloneqq\inf_{P\in\mathbb{S}^{n}_{+}}\textnormal{tr}(VP)\;\textnormal{subject to}\;M_{\mathcal{D}}(A,P)\succeq M.$ (32)
The optimal value function $\phi_{\mathcal{D}}:\real^{n\times n}\rightarrow\mathbb{R}_{\geq 0}\cup\set{\infty}$ is a barrier function for the constraint $A\in\mathbb{A}_{\mathcal{D}}^{n}$. Theorem 2 establishes properties of $\phi_{\mathcal{D}}$ and its $\varepsilon^{-1}$-sublevel sets (see Appendix D for proof).
Theorem 2.
Let $V\in\mathbb{S}^{n}_{++}$ and $M\in\mathbb{S}^{n}_{+}$ such that $M_{\mathcal{D}}(A,P)\succeq M$ implies $M_{\mathcal{D}}(A,P)\succ 0$. The following statements hold.
  1. (a)
    $\phi_{\mathcal{D}}$ is continuous on $\mathbb{A}_{\mathcal{D}}$.
  2. (b)
    For each $\varepsilon>0$, $\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon)$ is equivalent to the $\varepsilon^{-1}$-sublevel set of $\phi_{\mathcal{D}}$, i.e.,
[…]
    and both are closed.
  3. (c)
    $\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon)\nearrow\mathbb{A}_{\mathcal{D}}^{n}$ as $\varepsilon\searrow 0$.
Remark 11.
To reconstruct eq. 30 via Theorem 2, we set $M=sW\oplus 0_{n\times n}$ for any $W,V\succ 0$ and $s>0$ and apply the Schur complement lemma to $M_{\mathcal{D}_{2}}(A,P)/s-M/s$, where $\mathcal{D}_{2}$ is the circle defined in Lemma 1 with $x_{0}=0$, and $M_{\mathcal{D}_{2}}$ is defined by the generating matrices used in Lemma 1. Then the $\varepsilon^{-1}$-sublevel set of $\phi_{\mathcal{D}_{2}}$ equals the set of $A\in\real^{n\times n}$ for which eq. 30 is feasible.
Similarly, we can reconstruct the set of $A\in\real^{n\times n}$ for which eq. 31 is feasible as $\varepsilon^{-1}$-sublevel sets of $\phi_{\mathcal{D}_{1}}$, where $\mathcal{D}_{1}$ is the shifted half-plane defined in Lemma 1, and $M=W$ for any $W,V\succ 0$.
5 Cholesky substitution and elimination
In this section, we seek to approximate certain NSDPs by NLPs. Specifically, we consider the NSDP
$\min_{(\beta,\Sigma)\in\Theta}f(\beta,\Sigma)$ (34)
where $\Theta$ is defined as in eq. 20. This covers both ML eq. 3 and MAP eq. 4 problems with constraints eq. 20. We combine Cholesky factor-based substitution with an elimination scheme to convert the NSDP to a NLP while adding just $|\mathcal{I}_{\mathcal{A}}|$ variables to the optimization problem.
For this section, we define the following notation. For each $\mathcal{I}\subseteq\mathcal{L}^{n}$, let $\pi_{\mathcal{I}}^{L}:\real^{n\times n}\rightarrow\mathbb{L}^{n}[\mathcal{I}]$ and $\pi_{\mathcal{I}}:\real^{n\times n}\rightarrow\mathbb{S}^{n}[\mathcal{I}]$ denote the orthogonal projections (in the Frobenius norm) from n×n onto the subspaces $\mathbb{L}^{n}[\mathcal{I}]$ and $\mathbb{S}^{n}[\mathcal{I}]$, respectively. Let $\textnormal{chol}:\mathbb{S}^{n}_{++}\rightarrow\mathbb{L}^{n}_{++}$ denote the invertible function that maps a positive definite matrix to its Cholesky factor.
5.1 Burer-Monteiro-Zhang method
We first consider the simplified constraint set
[…]
where $\mathcal{D}^{n}\subseteq\mathcal{I}\subseteq\mathcal{L}^{n}$ and $H:\real^{m}\rightarrow\mathbb{S}^{n}$. As in [36], we parameterize the matrix argument $Q$ in a way that automatically enforces the constraint $Q\succ H(x)$ while introducing just $n$ scalar inequality constraints.
Recall $Q\succ H$ if and only if $Q=H+LL^{\top}$ for the unique matrix $L=\textnormal{chol}(Q-H)\in\mathbb{L}^{n}_{++}$. With $\mathcal{J}\coloneqq\mathcal{L}^{n}\setminus\mathcal{I}$, we can split $L$ into a sum of $L^{\mathcal{I}}\in\mathbb{L}^{n}_{++}[\mathcal{I}]$ and $L^{\mathcal{J}}\in\mathbb{L}^{n}[\mathcal{J}]$, giving
$Q=H+(L^{\mathcal{I}}+L^{\mathcal{J}})(L^{\mathcal{I}}+L^{\mathcal{J}})^{\top}.$ (36)
But $Q\in\mathbb{S}^{n}[\mathcal{I}]$, so we can apply the vectorization operator $\textnormal{vecs}_{\mathcal{J}}$ on both sides to give
$\textnormal{vecs}_{\mathcal{J}}(H+(L^{\mathcal{I}}+L^{\mathcal{J}})(L^{\mathcal{I}}+L^{\mathcal{J}})^{\top})=0.$ (37)
Equation 37 defines $|\mathcal{J}|$ equations to solve for the $|\mathcal{J}|$ variables of $L^{\mathcal{J}}$, where each $L_{ij}^{\mathcal{J}}$ is fully specified by $H_{ij}$ and the $L_{i^{\prime}j^{\prime}}$ with $(i^{\prime},j^{\prime})<(i,j)$. [^5] The lexicographic order << on $\mathbb{I}^{2}$ is defined by $(i,j)<(i^{\prime},j^{\prime})$ if $i<i^{\prime}$ or $(i=i^{\prime})\land(j<j^{\prime})$. In Algorithm 2, we compute the $L_{ij}^{\mathcal{J}}$ in ascending lexicographic order via Cholesky factorization.
Algorithm 2 Cholesky factorization algorithm for solving systems of the form eq. 37 based on [36, Lem. 1].
[…]
2:for each $(i,j)\in\mathcal{J}$ in ascending lexicographic order do
3:  $L^{\mathcal{J}}_{ij}\leftarrow-\frac{1}{L_{jj}^{\mathcal{I}}}(H_{ij}+\sum_{k=1}^{j-1}(L^{\mathcal{I}}_{ik}+L^{\mathcal{J}}_{ik})(L^{\mathcal{I}}_{jk}+L^{\mathcal{J}}_{jk}))$
4:end for
5:return  $L^{\mathcal{J}}$
Notice that each $L^{\mathcal{J}}$ is fully defined by $H$ and $L^{\mathcal{I}}$ via algorithm 2, so we have proven the following lemma.
Lemma 2 ([36, Lem. 1]).
For each $(H,L^{\mathcal{I}})\in\mathbb{S}^{n}\times\mathbb{L}^{n}[\mathcal{I}]$ such that $L^{\mathcal{I}}_{ii}\neq 0$ for each $i\in\mathbb{I}_{1:n}$, there is a unique $L^{\mathcal{J}}\in\mathbb{L}^{n}[\mathcal{J}]$ satisfying eq. 37.
With a slight abuse of notation, we let $L^{\mathcal{J}}=L^{\mathcal{J}}(H,L^{\mathcal{I}})$ denote the function defined by Algorithm 2, which maps each $(H,L^{\mathcal{I}})\in\mathbb{S}^{n}\times\mathbb{L}^{n}_{++}[\mathcal{I}]$ to the matrix $L^{\mathcal{J}}\in\mathbb{L}^{n}[\mathcal{J}]$ uniquely satisfying eq. 37. Moreover, we let
$Q(H,L^{\mathcal{I}})\coloneqq H+(L^{\mathcal{I}}+L^{\mathcal{J}}(H,L^{\mathcal{I}}))(L^{\mathcal{I}}+L^{\mathcal{J}}(H,L^{\mathcal{I}}))^{\top}$
as in eq. 36. Clearly $Q(H,L^{\mathcal{I}})\succ H$ is satisfied by definition. Finally, we define the transformation
$T(x,L^{\mathcal{I}})\coloneqq\left(x,Q(H(x),L^{\mathcal{I}})\right)$ (38)
which has the inverse
$T^{-1}(x,Q)\coloneqq\left(x,\pi_{\mathcal{I}}^{L}[\textnormal{chol}(Q-H(x))]\right)$ (39)
and we have the following lemma.
Lemma 3 ([36, Lem. 2]).
The function $T$ defined by eq. 38 is a bijection between $\real{}^{m}\times\mathbb{L}^{n}_{++}[\mathcal{I}]$ and $\textnormal{int}(\mathcal{P})$.
Differentiability of $T$ and $T^{-1}$ follow from differentiability of $H$ and algorithm 2. In fact, these functions are as smooth as $H$. More importantly, the bijection $T$ allows us to transform the minimum of a continuous function over $\mathcal{P}$ to an infimum over $\real{}^{m}\times\mathbb{L}^{n}_{++}[\mathcal{I}]$, given by the following theorem.
Theorem 3 ([36, Thm. 1]).
If $f:\mathcal{P}\rightarrow\real$ is continuous and attains a minimum in $\mathcal{P}$, then
$\min_{(x,Q)\in\mathcal{P}}f(x,Q)=\inf_{(x,L^{\mathcal{I}})\in\real^{m}\times\mathbb{L}^{n}_{++}[\mathcal{I}]}f(T(x,L^{\mathcal{I}})).$ (40)
We reiterate the proof of Theorem 3 for illustrative purposes.
Proof.
Continuity of $f$ implies its minimum over $\mathcal{P}$ equals its infimum over $\textnormal{int}(\mathcal{P})$, i.e.,
$\min_{(x,Q)\in\mathcal{P}}f(x,Q)=\inf_{(x,Q)\in\textnormal{int}(\mathcal{P})}f(x,Q)$
Since $T$ is a bijection, we can transform the optimization variables as follows:
$\inf_{(x,Q)\in\textnormal{int}(\mathcal{P})}f(x,Q)=\inf_{(x,L^{\mathcal{I}})\in T^{-1}(\textnormal{int}(\mathcal{P}))}f(T(x,L^{\mathcal{I}})).$
Finally, since $\real{}^{m}\times\mathbb{L}^{n}_{++}[\mathcal{I}]=T^{-1}(\textnormal{int}(\mathcal{P}))$, we have eq. 40. ∎
5.2 Generalized Burer-Monteiro-Zhang method
We return to constraints of the form eq. 20. Recall 1 requires the matrix inequalities are strictly feasible in the constraint set. We use a similar procedure to Section 5.1, but Algorithm 2 must be applied to each strict inequality $\Sigma\succ H$ and $\mathcal{A}(\beta,\Sigma)\succ 0$.
For the sparse symmetric matrix $\Sigma$ and matrix inequality $\Sigma\succ H(\beta)$, the procedure is the same as in Section 5.1. Let $L^{\mathcal{J}_{\Sigma}}=L^{\mathcal{J}_{\Sigma}}(H,L^{\mathcal{I}_{\Sigma}})$ denote the function defined by Algorithm 2 with $L^{\mathcal{I}}=L^{\mathcal{I}_{\Sigma}}$, $\mathcal{I}=\mathcal{I}_{\Sigma}$, and $n=n_{\Sigma}$. Then
$\Sigma(\beta,L^{\mathcal{I}_{\Sigma}})\coloneqq H+(L^{\mathcal{I}_{\Sigma}}+L^{\mathcal{J}_{\Sigma}}(H,L^{\mathcal{I}_{\Sigma}}))(L^{\mathcal{I}_{\Sigma}}+L^{\mathcal{J}_{\Sigma}}(H,L^{\mathcal{I}_{\Sigma}}))^{\top}$
guarantees $\Sigma(H,L^{\mathcal{I}_{\Sigma}})\succ H$ and $\Sigma(H,L^{\mathcal{I}_{\Sigma}})\in\mathbb{S}^{n_{\Sigma}}[\mathcal{I}_{\Sigma}]$ for all $(H,L^{\mathcal{I}_{\Sigma}})\in\mathbb{S}^{n_{\Sigma}}\times\mathbb{L}^{n_{\Sigma}}_{++}[\mathcal{I}_{\Sigma}]$. In other words, $\Sigma$ is fully defined and the constraint $\Sigma\succ H$ automatically satisfied by $(H,L^{\mathcal{I}_{\Sigma}})\in\mathbb{S}^{n_{\Sigma}}\times\mathbb{L}^{n_{\Sigma}}_{++}[\mathcal{I}_{\Sigma}]$.
For the general matrix inequality $\mathcal{A}(\beta,\Sigma)\succeq 0$, the procedure is slightly different. Let $L^{\mathcal{J}_{\mathcal{A}}}=L^{\mathcal{J}_{\mathcal{A}}}(L^{\mathcal{I}_{\mathcal{A}}})$ denote function defined by Algorithm 2 with $L^{\mathcal{I}}=L^{\mathcal{I}_{\mathcal{A}}}$, $\mathcal{I}=\mathcal{I}_{\mathcal{A}}$, $n=n_{\mathcal{A}}$, and $H=0$. Define the functions
$\mathcal{A}(L^{\mathcal{I}_{\mathcal{A}}})\coloneqq(L^{\mathcal{I}_{\mathcal{A}}}+L^{\mathcal{J}_{\mathcal{A}}}(L^{\mathcal{I}_{\mathcal{A}}}))(L^{\mathcal{I}_{\mathcal{A}}}+L^{\mathcal{J}_{\mathcal{A}}}(L^{\mathcal{I}_{\mathcal{A}}}))^{\top}$
which guarantees $\mathcal{A}(L^{\mathcal{I}_{\mathcal{A}}})\in\mathbb{S}^{n_{\mathcal{A}}}_{++}[\mathcal{I}_{\mathcal{A}}]$ for all $L^{\mathcal{I}_{\mathcal{A}}}\in\mathbb{L}^{n_{\mathcal{A}}}_{++}[\mathcal{I}_{\mathcal{A}}]$. However, the constraint is not fully eliminated; we are left with $|\mathcal{I}_{\mathcal{A}}|$ equality constraints in the transform space,
$\textnormal{vecs}_{\mathcal{I}_{\mathcal{A}}}(\mathcal{A}(\beta,\Sigma(H(\beta),L^{\mathcal{I}_{\Sigma}}))-\mathcal{A}(L^{\mathcal{I}_{\mathcal{A}}}))=0$
with the other $|\mathcal{L}^{n_{\mathcal{A}}}\setminus\mathcal{I}_{\mathcal{A}}|$ constraints automatically guaranteed by Algorithm 2.
To define the new constraints, we require the variable transformations
$\mathcal{T}(\beta,L^{\mathcal{I}_{\Sigma}},L^{\mathcal{I}_{\mathcal{A}}})$ $\coloneqq\left(\beta,\Sigma(H(\beta),L^{\mathcal{I}_{\Sigma}})\right)$ (41a)
$\mathcal{A}_{\mathcal{T}}(\beta,L^{\mathcal{I}_{\Sigma}},L^{\mathcal{I}_{\mathcal{A}}})$ $\coloneqq\mathcal{A}(L^{\mathcal{I}_{\mathcal{A}}})$ (41b)
which are well-defined for all $(\beta,L^{\mathcal{I}_{\Sigma}},L^{\mathcal{I}_{\mathcal{A}}})\in\real^{n_{\beta}}\times\mathbb{L}^{n_{\Sigma}}[\mathcal{I}_{\Sigma}]\times\mathbb{L}^{n_{\mathcal{A}}}[\mathcal{I}_{\mathcal{A}}]$. With the functions eq. 41, we define the transformed constraint functions $(g_{\mathcal{T}},h_{\mathcal{T}})$ and the transformed constraint set $\Phi\subseteq\real^{n_{\beta}}\times\mathbb{L}^{n_{\Sigma}}[\mathcal{I}_{\Sigma}]\times\mathbb{L}^{n_{\mathcal{A}}}[\mathcal{I}_{\mathcal{A}}]$ according to eqs. 22 and 23. The inverse transform is
$\mathcal{T}^{-1}(\beta,\Sigma)\coloneqq\left(\beta,\pi_{\mathcal{I}_{\Sigma}}^{L}[\textnormal{chol}(\Sigma-H(\beta))],\pi_{\mathcal{I}_{\mathcal{A}}}^{L}[\textnormal{chol}(\mathcal{A}(\beta,\Sigma))]\right)$ (42)
for all $(\beta,\Sigma)\in\Theta_{++}$, and we have the following lemma.
Lemma 4.
The function $\mathcal{T}$ defined by eq. 41 is a bijection between $\Phi$ and $\Theta_{++}$.
Proof.
First, we have $\mathcal{T}(\Phi)\subseteq\Theta_{++}$ since the transformed constraints guarantee the constraints $g(\beta,\Sigma)=0$, $h(\beta,\Sigma)\leq 0$, $\Sigma\succ H(\beta)$, and $\mathcal{A}(\beta,\Sigma)\succ 0$ for any $(\beta,\Sigma)\coloneqq\mathcal{T}(\phi)$ and $\phi\in\Phi$. Next, it is clear by construction that $\mathcal{T}^{-1}\circ\mathcal{T}$ is the identity map on $\Phi$. Therefore $\mathcal{T}$ is injective. Similarly, we have $\mathcal{T}^{-1}(\Theta_{++})\subseteq\Phi$ by construction, and $\mathcal{T}\circ\mathcal{T}^{-1}$ is the identity map on $\Theta_{++}$, so $\mathcal{T}:\Phi$ is surjective. ∎
Under 1, the functions $\mathcal{T}$, $\mathcal{T}^{-1}$, and $\mathcal{A}_{\mathcal{T}}$ are as smooth as $H$, and moreover, the bijection $\mathcal{T}$ transforms a minimum over $\Theta$ into an infimum over $\Phi$.
Proposition 5.
If 1 holds and $f:\Theta\rightarrow\real$ is continuous and attains a minimum in $\Theta$, then
$\min_{\theta\in\Theta}f(\theta)=\inf_{\phi\in\Phi}f(\mathcal{T}(\phi)).$
Proof.
The proof follows that of Theorem 3, noting that 1 gives $\textnormal{cl}(\Theta_{++})=\Theta$ and therefore the minimum of $f$ over $\Theta$ equals the infimum of $f$ over $\Theta_{++}$. ∎
5.3 Approximate solutions
As mentioned in Section 3, we consider a lower bound $\varepsilon>0$ on the diagonal elements of $(L^{\mathcal{I}_{\Sigma}},L^{\mathcal{I}_{\mathcal{A}}})$. We define the tightened constraint set $\Phi_{\varepsilon}$ by eq. 24. In the following theorem we show, under 1 and continuity of $f$, the infimum of $f\circ\mathcal{T}$ over $\Phi_{\varepsilon}$ converges to the minimum of $f$ over $\Theta$ (see Appendix E for proof).
Theorem 4.
Suppose $f$ is continuous and attains a minimum in $\Theta$. Define
$\mu_{0}\coloneqq\min_{\theta\in\Theta}f(\theta)$
and
$\mu_{\varepsilon}\coloneqq\inf_{\phi\in\Phi_{\varepsilon}}f(\mathcal{T}(\phi))$ (43)
for each $\varepsilon>0$. If 1 holds, then $\mu_{\varepsilon}\searrow\mu$ as $\varepsilon\searrow 0$.
In fact, with a few additional requirements on the objective $f$, convergence of approximate problem solutions to the solution of the original problem is guaranteed by the following theorem (see Appendix E for proof).
Theorem 5.
Suppose $f$ is continuous and 1 holds. Consider the set-valued function $\hat{\theta}:\mathbb{R}_{\geq 0}\rightarrow\mathcal{P}(\Theta)$, defined as $\hat{\theta}_{\varepsilon}\coloneqq\operatorname*{argmin}_{\theta\in\mathcal{T}(\Phi_{\varepsilon})}f(\theta)$ for all $\varepsilon>0$, and $\hat{\theta}_{0}\coloneqq\operatorname*{argmin}_{\theta\in\Theta}f(\theta)$. If there exists $\alpha\in\real$ and compact $C\subseteq\Theta$ such that
[…]
is contained in $C$ and $\Theta_{f\leq\alpha}\cap\Theta_{++}$ is nonempty, then there exists $\overline{\varepsilon}>0$ such that, for all $\varepsilon_{0}\in[0,\overline{\varepsilon})$,
  1. (a)
    $f$ achieves a minimum in $\Theta$ and $\hat{\theta}_{0}$ is nonempty;
  2. (b)
    if $\varepsilon_{0}>0$, then $f$ achieves a minimum in $\mathcal{T}(\Phi_{\varepsilon_{0}})$ and $\hat{\theta}_{\varepsilon_{0}}$ is nonempty;
  3. (c)
    $\mu_{\varepsilon}$ is continuous and $\hat{\theta}_{\varepsilon}$ is outer semicontinuous at $\varepsilon=\varepsilon_{0}$; and
  4. (d)
    if $\hat{\theta}_{0}$ is a singleton, then $\limsup_{\varepsilon\searrow 0}\hat{\theta}_{\varepsilon}=\hat{\theta}_{0}$.
Remark 12.
A log-barrier approach was used by Burer et al. [36] to handle strict inequalities in $L\in\mathbb{L}^{n}_{++}[\mathcal{I}]$ and achieve global convergence for linear SDPs. For the problem eq. 40, the log-barrier term eliminates all remaining constraints. However, for the problem eq. 34, many constraints remain in addition to the strict inequalities on the diagonal elements of $(L^{\mathcal{I}_{\Sigma}},L^{\mathcal{I}_{\mathcal{A}}})\in\mathbb{L}^{n_{\Sigma}}_{++}[\mathcal{I}_{\Sigma}]\times\mathbb{L}^{n_{\mathcal{A}}}_{++}[\mathcal{I}_{\mathcal{A}}]$.
6 Case Studies
In this section, we present two real-world case studies in which Algorithm 1 is used to identify the LADM eq. 2 and implement offset-free MPC. In the first case study, we consider the temperature control laboratory (TCLab) (see Figure 1), an Arduino-based temperature control laboratory that serves as a low-cost [^6] The TCLab is available for under \$40 from https://apmonitor.com/heat.htm and https://www.amazon.com/gp/product/B07GMFWMRY. benchmark for linear MIMO control [38]. We identify the TCLab from open-loop data and use the resulting model to design an offset-free MPC. We compare closed-loop control and estimation performance of these models to that of offset-free MPCs designed with the identification methods from [16, 17]. In the second case study, data from an industrial-scale chemical reactor is used to design Kalman filters for the linear augmented disturbance model, and the closed-loop estimation performance is compared to that of the designs proposed in [17].
Figure 1: Benchmark temperature Control Laboratory (TCLab) [38].
Throughout these experiments, we use an $\ell_{2}$ regularization term in the transformed space, [^7] With $L_{\Sigma}(\overline{\beta},\overline{L}^{\mathcal{I}_{\Sigma}})=0$, the last term of eq. 44 becomes proportional to $\textnormal{tr}(L_{\Sigma}L_{\Sigma}^{\top})=\textnormal{tr}(\Sigma)$ where $L_{\Sigma}=L_{\Sigma}(\beta,L^{\mathcal{I}_{\Sigma}})$ and $\Sigma=\Sigma(\beta,L^{\mathcal{I}_{\Sigma}})$., [^8] With $L^{\mathcal{J}_{\Sigma}}(\beta,L^{\mathcal{I}_{\Sigma}})\equiv 0$ (e.g., $\Sigma$ is block diagonal and $H(\beta)\equiv 0$) the last term of eq. 44 is proportional to $\|L^{\mathcal{I}_{\Sigma}}-\overline{L}^{\mathcal{I}_{\Sigma}}\|_{\mathrm{F}}^{2}=|\textnormal{vec}_{\mathcal{I}_{\Sigma}}(L^{\mathcal{I}_{\Sigma}}-\overline{L}^{\mathcal{I}_{\Sigma}})|^{2}$.
$\ln p_{0}(\beta,L^{\mathcal{I}_{\Sigma}})\propto R_{0}(\beta,L^{\mathcal{I}_{\Sigma}})\coloneqq\frac{\rho}{2}\left(|\beta-\overline{\beta}|^{2}+\|L_{\Sigma}(\beta,L^{\mathcal{I}_{\Sigma}})-L_{\Sigma}(\overline{\beta},\overline{L}^{\mathcal{I}_{\Sigma}})\|_{\mathrm{F}}^{2}\right).$ (44)
where $\rho\geq 0$ is the regularization weight and $(\overline{\beta},\overline{L}^{\mathcal{I}_{\Sigma}},\overline{L}^{\mathcal{I}_{\mathcal{A}}})$ denote the initial guess for the optimizer. The variable $L_{\mathcal{A}}$ is not regularized. With $\rho=0$, the MAP problem eq. 4 with the regularizer eq. 44 simplifies to the standard ML identification problem eq. 3.
The initial guess for the ML and MAP problems is based on a nested ML estimation approach described in [16, 17]. The initial guess methods effectively augment standard identification methods (e.g., principal component analysis (PCA), Ho-Kalman (HK), canonical correlation analysis (CCA) algorithms), so we refer to the initial guess models as “augmented” versions of the standard method being used (e.g., augmented PCA, augmented HK, augmented CCA). Each optimization problem is formulated in CasADi via Algorithm 1 and solved with IPOPT. Information about each model fit and configuration is presented in Table 1. Wall times for a single-thread of an Intel Core i9-10850K processor are reported.
6.1 Benchmark temperature controller
Table 1: TCLab model fitting results.  The augmented PCA/ARX identification methods are not iterative. ∗∗ The maximum number of iterations was set at 500.
Model Results Configuration
[…]
Augmented PCA 0.02 N/A 3823.4 [16] N/A N/A N/A N/A
Augmented ARX 0.03 N/A 3807.3 see text N/A N/A N/A N/A
Unregularized ML 125.48 500∗∗ -9430.9 Algo. 1 0 $\mathbb{C}$ 10-6 N/A
Regularized ML 1 126.05 500∗∗ -9431.7 Algo. 1 0.002 $\mathbb{C}$ 10-6 N/A
Regularized ML 2 9.23 21 -9416.6 Algo. 1 0.005 $\mathbb{C}$ 10-6 N/A
Constrained ML 1 74.88 97 -9347.2 Algo. 1 0 $\mathcal{D}_{1}(0.3)\cap\mathcal{D}_{2}(0.998,0)$ 10-6 0.03
Constrained ML 2 51.17 62 -9358.2 Algo. 1 0 $\mathcal{D}_{1}(0.3)\cap\mathcal{D}_{2}(0.999,0)$ 10-6 0.03
Reg. & Cons. ML 37.07 40 -9338.4 Algo. 1 0.001 $\mathcal{D}_{1}(0.3)\cap\mathcal{D}_{2}(0.998,0)$ 10-6 0.03
Figure 2: TCLab identification data and noise-free responses $\hat{y}_{k}=\sum_{j=1}^{k}\hat{C}\hat{A}^{j-1}\hat{B}u_{k-j}$ of a few selected models.
Unless otherwise specified, the TCLab is modeled as a two-state, two-disturbance system of the form eq. 2, with internal temperatures as plant states $s=\begin{bmatrix}T_{1}&T_{2}\end{bmatrix}^{\top}$, heater voltages as inputs $u=\begin{bmatrix}V_{1}&V_{2}\end{bmatrix}^{\top}$, and measured temperatures $y=\begin{bmatrix}T_{m,1}&T_{m_{2}}\end{bmatrix}^{\top}$ as outputs. Throughout, we choose $n_{d}=p$ to satisfy the offset-free necessary conditions in [3, 4], and we consider output disturbance models $(B_{d},C_{d})=(0_{2\times 2},I_{2})$. We use $(A_{s},B_{s})$ fully parameterized and $C=I_{2}$ to guarantee model identifiability and make the states interpretable as internal temperatures. For the remaining model terms, we have $(K_{x},K_{d},R_{e})$ fully parameterized and $(D,\hat{s}_{0},\hat{d}_{0})=(0,0,0)$.
Eight TCLab models are presented.
  1. 1.
    Augmented PCA: the 6-state TCLab model used in [16], where principle component analysis on a $400\times 5100$ data Hankel matrix is used to determine the states in the disturbance-free model.
  2. 2.
    Augmented ARX: a VARX(1,1)(1,1) model, equivalent to a stochastic LTI model with process noise but zero measurement noise.
  3. 3–5)
    Unregularized ML, Regularized ML 1 and 2: classic ML and MAP models.
  4. 6–8)
    Constrained ML 1 and 2, Reg. & Cons. ML: eigenvalue-constrained ML and MAP models. LMI region constraints enforce filter stability and impose a lower bound on the real part of the filter eigenvalues.
Each ML model uses Augmented ARX as the initial guess. See Table 1 for specifics of the formulation of models 3–8.
[…]
Except for the augmented PCA model, all of the open-loop eigenvalues cluster around the same region of the complex plane (figure 3). The closed-loop filter eigenvalues are also placed similarly, although the classic ML models (Unregularized ML, Regularized ML 1 and 2) suffer from slow or even unstable filter eigenvalues. The models with unstable eigenvalues fail to converge (see Table 1) as the unstable filter modes make the problem extremely sensitive to changes in the parameter values. While a high $\rho$ is sufficient to achieve filter stability, there is no clear minimum value of $\rho$ to achieve this. On the other hand, the constrained ML models have stable filter eigenvalues without regularization, and have well-defined estimator performance guarantees based on the applied constraints. We remark the relative performance is unchanged by removal of the bad sensor value (around $t=1800$ s in Figure 2).
Figure 3: TCLab models open-loop and closed-loop (filter) eigenvalues.
Figure 4: TCLab setpoint tracking tests.
Figure 5: TCLab disturbance rejection tests.
Figure 6: TCLab setpoint tracking test performance.
Figure 7: TCLab disturbance rejection test performance.
To test offset-free control performance, we performed two sets of closed-loop experiments on offset-free MPCs designed with the models. In Figure 4, identical setpoint changes were applied to a TCLab running at a steady-state power output of 50%. The setpoint changes were tracked with the offset-free MPC design described in [16]. In Figure 5, step disturbances in the output $p_{i}$ and the input $m_{i}$ are injected into a plant trying to maintain a given steady-state temperature. The setpoints are tracked with the offset-free MPC design described in [16].
Control performance is quantified by the squared distance from the setpoint $\ell_{k}\coloneqq|y_{k}-y_{\textnormal{sp},k}|^{2}$. Estimation performance is quantified by the squared filter errors $e_{k}^{\top}e_{k}$. For any signal $a_{k}$, we define a $T$-sample moving average by $\langle a_{k}\rangle_{T}\coloneqq T^{-1}\sum_{j=0}^{T-1}a_{k-j}$. Setpoint tracking performance is reported in Figure 6, and disturbance rejection performance is reported in Figure 7. The worst performing models are those with unstable filters (Unregularized ML and Regularized ML 1). These models shut off over the course of the experiment as the integrating disturbance estimates grow unbounded. The remaining classic ML model (Regularized ML 2) has slow filter eigenvalues that contribute to poor control performance on the disturbance rejection test (see Figure 7, left). The augmented models (Augmented PCA/ARX) perform poorly in either control or estimation aspect on both tests. The best performance is achieved by the remaining ML models, which all perform approximately the same across the tests.
Figure 8: TCLab identification index data for (left) setpoint tracking and (right) disturbance rejection tests.
To investigate the distributional accuracy of the models, we consider the identification index $q\coloneqq e^{\top}R_{e}^{-1}e$. Recall the signal $e_{k}\in\real^{2}$ is an i.i.d., zero-mean Gaussian process, i.e., $e_{k}\overset{\textnormal{iid}}{\sim}\mathcal{N}(0,R_{e})$, and therefore the index $q_{k}$ is i.i.d. with a $\chi^{2}_{2}$ distribution. Moreover, the moving average $\langle q_{k}\rangle_{T}$ is distributed as $\chi^{2}_{2T}/T$, although it is no longer independent in time. In Figure 8, histograms of $\langle q\rangle_{T},T\in\set{1,10,100}$ are plotted against their expected distribution for a few selected models (Augmented PCA/ARX, Unregularized ML, and Reg. & Cons. ML). The extreme discrepancies between the augmented models’ performance index $\langle q\rangle_{T}$ and the reference distribution $\chi^{2}_{2T}/T$ are primarily due to the augmented models significantly overestimating $R_{e}$ compared to the ML models,
$\hat{R}_{e}^{\textnormal{Aug.~PCA}}$ $=\begin{bmatrix}0.5871&0.3365\\ 0.3365&0.2878\end{bmatrix}$,
$\hat{R}_{e}^{\textnormal{Aug.~ARX}}$ $=\begin{bmatrix}0.5084&0.2198\\ 0.2198&0.2980\end{bmatrix}$,
$\hat{R}_{e}^{\textnormal{Unreg.~ML}}$ $=\begin{bmatrix}0.0106&0.0007\\ 0.0007&0.008\end{bmatrix}$,
$\hat{R}_{e}^{\textnormal{Reg.~Cons.~ML}}$ $=\begin{bmatrix}0.0107&0.0007\\ 0.0007&0.008\end{bmatrix}$.
The reference distribution and the ML models’ $\langle q\rangle_{T}$ distribution diverge at large $T$ since, due to plant-model mismatch, the filter’s innovation errors are autocorrelated. Filter instability of the unregularized ML model causes frequent right-tail errors.
6.2 Eastman reactor
$\overline{F}_{2}$
$\overline{T}_{H}$
$\overline{F}_{1}$
$F_{1,\textrm{sp}}$
$T_{\textrm{sp}}$
$F_{2,\textrm{sp}}$
MeOH flow rate, $F_{1}$
Reactor temp, $T$
Slurry flow rate, $F_{2}$
Hot oil TC.op, $T_{H}$
Quality param 1, $r_{1}$
Quality param 2, $r_{2}$
MPC
Hot oil return
Hot oil supply
Recycle
Sludge
Xylene
Crude DMT,
H2O,
MeOH
Slurry
MeOH
Figure 9: Schematic of the DMT reactor and MPC control strategy.
Table 2: Eastman reactor model fitting results. The data is a reduced version of the closed-loop data from [17].  The augmented identification methods are not iterative. ∗∗ The maximum number of iterations was set at 500.
Model Results Configuration
[…]
Augmented CCA 0.06 N/A -11399.3 [17] N/A N/A N/A N/A
Unregularized ML 5.47 19 -14383.1 Algo. 1 0 $\mathbb{C}$ 10-6 N/A
Regularized ML 1 5.28 17 -14362.5 Algo. 1 0.01 $\mathbb{C}$ 10-6 N/A
Regularized ML 2 5.61 20 -14346.7 Algo. 1 0.10 $\mathbb{C}$ 10-6 N/A
Regularized ML 3 4.80 13 -14108.0 Algo. 1 1.00 $\mathbb{C}$ 10-6 N/A
Constrained ML 1 20.03 92 -13944.9 Algo. 1 0 $\mathcal{D}_{1}(0.3)$ 10-6 0.01
Constrained ML 2 16.60 73 -13941.1 Algo. 1 0 $\mathcal{D}_{1}(0.3)$ 10-6 0.02
Constrained ML 3 13.99 58 -13928.5 Algo. 1 0 $\mathcal{D}_{1}(0.3)$ 10-6 0.04
Figure 10: Training data and noise-free responses for the Eastman reactor models (Augmented HK and ML models using Augmented HK as the initial guess).
Figure 11: Eastman reactor models open-loop and closed-loop (filter) eigenvalues.
Figure 12: Test performance for the Eastman reactor models on the test data sets from [17].
Figure 13: Eastman reactor models’ closed-loop (filter) response to the eigenvector corresponding to the fastest eigenvalue.
A schematic of the chemical reactor considered in the next case study is presented in Figure 9.The control objective of the chemical reactor is to track three setpoints (the output, a specified reactor temperature $y=T$, and the flowrates $\begin{bmatrix}u_{1}&u_{3}\end{bmatrix}^{\top}=\begin{bmatrix}F_{1}&F_{2}\end{bmatrix}^{\top}$), without offset, by controlling the three inputs (the reactant flow rates and utility temperatures $u=\begin{bmatrix}F_{1}&T_{H}&F_{2}\end{bmatrix}^{\top}$). [^9] The flowrates are both manipulated variables and controlled variables. At steady-state, we should reach the setpoints in $y=T$ and $\begin{bmatrix}u_{1}&u_{3}\end{bmatrix}^{\top}=\begin{bmatrix}F_{1}&F_{2}\end{bmatrix}^{\top}$, but $u_{2}=T_{H}$ will not reach a predefined setpoint. See [17] for more details about the reactor operation. As in Section 6.1, we choose $n_{d}=p$ and consider output disturbance models $(B_{d},C_{d})=(0_{2\times 1},1)$. This time, we use an observability canonical form [41] with $A_{s}=\begin{bmatrix}0&1\\ a_{1}&a_{2}\end{bmatrix}$ and $C_{s}=\begin{bmatrix}1&0\end{bmatrix}$. For the remaining model terms, we have $(B_{s},K_{x},K_{d},R_{e})$ fully parameterized and $(D,\hat{s}_{0},\hat{d}_{0})=(0,0,0)$.
Eight reactor models were fit to closed-loop data:
  1. 1.
    Augmented CCA: a CCA model [69] augmented with a disturbance model, as detailed in [17]. [^10] This is not the same model used in [17], as a different input-output model is considered, although the same data is used. Specifically, Kuntz et al. [17] considered a model with both regulatory-layer setpoints and measured values.
  2. 2–5)
    Unregularized ML, Regularized ML 1 to 3: classic ML and MAP models.
  3. 6–8)
    Constrained ML 1 to 3: eigenvalue-constrained ML and MAP models. LMI region constraints impose a lower bound on the real part of the filter eigenvalues.
Each ML model uses the augmented CCA model as the initial guess. See Table 2 for specifics of the formulation of models 2–8. In Figure 10, the closed-loop identification data and noise-free responses are presented. Again, these responses do not reflect testing performance. Computational details, the unregularized log-likelihood value, and model configurations are reported in Table 2. The open-loop $A_{s}$ and closed-loop $A_{K}$ eigenvalues are plotted in Figure 11.
The main difference between eigenvalues of the unconstrained ML models (Unregularized ML and Regularized ML 1–3) and the constrained ML models (Constrained ML 1–3) are faster open-loop eigenvalues and closed-loop eigenvalues with possibly negative real part (see Figure 11). For the constrained ML models, the real part of this fast filter eigenvalue is bounded from below using the LMI region constraint $\mathcal{D}_{1}(0.3)$. As in the TCLab case study, high $\rho$ is sufficient to avoid negative eigenvalues, but there is no clear cutoff.
In Figure 12, the estimation performance for these filters is compared on two test data sets (from [17]). While the unconstrained models appear to have the best test performance, it is at a cost of undesirable estimate dynamics. In Figure 13, we plot the filter response to an initial guess equal to the eigenvector corresponding to the smallest eigenvalue of $A_{K}$. Those filters with eigenvalues having negative real parts exhibit overshoot in the estimate. The best performing filters without this behavior are the constrained ML models.
Figure 14: Eastman reactor models simulated closed-loop test performance.
Control performance could not be compared on the real plant due to cost and safety considerations. However, the closed-loop responses can be compared in simulation. In Figure 14, we plot simulated responses to a setpoint change. Each simulation considers the nominal closed-loop response (i.e., plant as the model, no noise) using the offset-free MPC design in [17] with $Q_{s}=1$ and $R_{s}=\textnormal{diag}(0.01,1,0.01)$. The regularized ML models exhibit significant overshoot in the response, whereas the unregularized ML model and constrained ML models do not.
6.3 Discussion
The main limitation of eigenvalued-constrained ML is computational cost. While constrained ML retains linear scaling in sample size NN, [^11] It is well-known that computation of the log-likelihood $L_{N}(\theta)$ and its derivatives scale linearly in $N$ [20]. Moreover, the constraint function computation scales independently of $N$. each LMI region constraint on an arbitrary system matrix $\tilde{A}\in\real^{\tilde{n}\times\tilde{n}}$ requires an additional $O(\tilde{n}^{2}(m^{2}+1))$ variables and $O(\tilde{n}^{2}m^{2})$ equality constraints. These requirements can be significantly reduced for spectral abscissa bounds $\mathcal{D}_{1}(s)$ and stability constraints $\mathcal{D}_{2}(s,0)$. As mentioned in remark 10, these constraints are quite similar to the “smooth” spectral radii and abscissa constraints of [67, 68], which only add $O(\tilde{n}^{2})$ variables and $O(\tilde{n}^{2})$ equality constraints. For eigenvalues constrained to the LMI regions $\mathcal{D}_{1}(s)$ or $\mathcal{D}_{2}(s,x_{0})$, implementing these constraints as a special case can reduce the computational cost significantly.
For a standard, black-box LADM eq. 2 with $n_{d}=p$, a canonical form for $(A_{s},B_{s},C_{s})$, and $(D,\hat{s}_{0},\hat{d}_{0})=(0,0,0)$, there are $O(n_{s}(p+m)+p^{2})$ variables before constraints are added, and $O(n_{s}^{2})$ variables after. Thus, fitting black-box models of large-scale systems is computationally prohibitive. However, as discussed in Section 2, networked systems may be represented by significantly fewer variables: $O(N_{u}n_{u}(p_{u}+m_{u})+N_{u}p_{u}^{2})$ without constraints, or $O(N_{u}n_{u}^{2})$ with constraints, where $N_{u}$ is the number of units or nodes, and $n_{u},m_{u},p_{u}$ are the number of states, inputs, and outputs per unit or node.
7 Conclusion
We propose an algorithm for identifying offset-free MPC-relevant models with ML identification. The algorithm is validated on real-world data in two case studies: a low-cost benchmark temperature controller, and an industrial-scale reactor. Our method maintains LADM filter stability, avoids fitting artificial high-frequency dynamics, and outperforms standard ML identification in both control and estimation performance.
The code, including sample TCLab datasets and scripts for model fitting, is made available at github.com/rawlings-group/mlid_2024. The Eastman reactor data is proprietary and is not made available. We will develop this code further for reliable applications on large-scale systems.
We conclude with some suggestions of future research. Since ML identified models are more distributionally accurate, they are more suitable to the performance monitoring technique of [70]. Integrated identification and offset-free controller validation may be possible by combining this method with ours. Recall that there are limitations to the performance of the Kalman filter for LADMs, as shown by Bageshwar and Borrelli [47] and in the case studies. Filter designs with eigenvalue constraints may overcome these limitations and deliver superior offset-free MPC performance. Finally, we have found ML identification sometimes produces poor open-loop gain estimates. Using prediction error minimization with a least squares objective (as a regularizer or in a multiobjective approach) may alleviate these issues.
Appendix A Proof of Proposition 1
Silverman [52] contains a more complete characterization of the DARE solutions for regulation problems with cross terms. However, this admits additional nullspace terms into the gain matrix which the Kalman filtering problem does not allow. We avoid nullspace terms through the assumption $R_{v}\succ 0$ and therefore streamline the proof of Proposition 1.
For the following definitions and lemmas, we consider the matrices $\mathcal{W}\coloneqq(A,B,C,D)$ corresponding to a noise-free system.
Definition 2.
The system $\mathcal{W}$ is left invertible on $\mathbb{I}_{0:k-1}$ if
$0=\begin{bmatrix}D\\ CB&D\\ \vdots&\ddots&\ddots\\ CA^{k-2}B&\ldots&CB&D\end{bmatrix}\begin{bmatrix}u_{0}\\ \vdots\\ u_{k-1}\end{bmatrix}$
implies $u_{0}=0$. The system $\mathcal{W}$ is left invertible if there is some $j\in\mathbb{I}_{>0}$ such that $\mathcal{W}$ is left invertible on $\mathbb{I}_{0:k-1}$ for all $k\geq j$.
Definition 3.
The system $\mathcal{W}$ is strongly detectable if $y_{k}\rightarrow 0$ implies $x_{k}\rightarrow 0$.
The following lemmas are taken directly from [52, Thms. 8, 18(iii)], but the proofs are omitted for the sake of brevity.
Lemma 5 ([52, Thm. 8]).
If $\mathcal{W}$ is left invertible, then $\mathcal{W}$ is strongly detectable if and only if (ABF,CDF)(A-BF,C-DF) is detectable for all $F$ of appropriate dimension.
Lemma 6 ([52, Thm. 18(iii)]).
If $\mathcal{W}$ is left invertible, then the DARE
$P=A^{\top}PA-(A^{\top}PB+C^{\top}D)(B^{\top}PB+D^{\top}D)^{-1}(B^{\top}PA+D^{\top}C)$
has a unique, stabilizing solution [^12] Contrary to in Section 2, here we mean the solution $P$ is stabilizing when ABK(P)A-BK(P) is stable, where $K(P)\coloneqq(B^{\top}PB+D^{\top}D)^{-1}B^{\top}P$. if and only if $\mathcal{W}$ is stabilizable and semistrongly detectable.
For the remainder of this section, we consider the full rank factorization
$\begin{bmatrix}Q_{w}&S_{wv}\\ S_{wv}^{\top}&R_{v}\end{bmatrix}=\begin{bmatrix}\tilde{B}\\ \tilde{D}\end{bmatrix}\begin{bmatrix}\tilde{B}^{\top}&\tilde{D}^{\top}\end{bmatrix}$
[…]
Lemma 7.
If $R_{v}\succ 0$ then $\tilde{\mathcal{W}}$ is left invertible.
Proof.
Left invertability on $\mathbb{I}_{0:k-1}$ is equivalent to
$0=\begin{bmatrix}\tilde{D}^{\top}\\ \tilde{B}^{\top}C^{\top}&\tilde{D}^{\top}\\ \vdots&\ddots&\ddots\\ \tilde{B}^{\top}(A^{\top})^{k-2}C^{\top}&\ldots&\tilde{B}^{\top}C^{\top}&\tilde{D}^{\top}\end{bmatrix}\begin{bmatrix}u_{0}\\ \vdots\\ u_{k-1}\end{bmatrix}$ (45)
implying $u_{0}=0$. But $R_{v}=\tilde{D}\tilde{D}^{\top}\succ 0$, so $\tilde{D}^{\top}$ has a zero nullspace. For each $k\in\mathbb{I}_{>0}$, the coefficient matrix of eq. 45 has a zero nullspace. Thus, $u_{0}=0$ and $\tilde{\mathcal{W}}$ is left invertible. ∎
Finally, we can prove Proposition 1.
Proof of Proposition 1.
[…]
Appendix B Proof of Proposition 2
Throughout this appendix, we define the set of $n\times n$ Hermitian, Hermitian positive definite, and Hermitian positive semidefinite matrices as $\mathbb{H}^{n}$, $\mathbb{H}^{n}_{++}$, and $\mathbb{H}^{n}_{+}$. Notice that $f_{\mathcal{D}}$ maps to Hermitian matrices so we can write it as $f:\mathbb{C}\rightarrow\mathbb{H}^{m}$. We define the extension of $M_{\mathcal{D}}$ to complex arguments $M_{\mathcal{D}}:\mathbb{C}^{n\times n}\times\mathbb{H}^{n}_{+}\rightarrow\mathbb{H}^{nm}$ as
$M_{\mathcal{D}}(A,P)\coloneqq M_{0}\otimes P+M_{1}\otimes(AP)+M_{1}^{\top}\otimes(AP)^{\mathrm{H}}.$
To show Proposition 2, we need a preliminary result about Hermitian positive semidefinite matrices, generalized from Lemma A.1 in [25].
Lemma 8.
For any $M\in\mathbb{H}^{n}$, if $M\succeq 0$ ($M\succ 0$) then $\textnormal{Re}(M)\succeq 0$ ($\textnormal{Re}(M)\succ 0$).
Proof.
With $M=\textnormal{Re}(M)+\iota\textnormal{Im}(M)$, it is clear $M$ Hermitian implies $\textnormal{Re}(M)$ is symmetric and $\textnormal{Im}(M)$ is skew-symmetric. Thus $v^{\top}Mv=v^{\top}\textnormal{Re}(M)v$ for all $v\in\real^{n}$, and positive (semi)definiteness of $M$ implies positive (semi)definiteness of $\textnormal{Re}(M)$. ∎
In proving Proposition 2, we take the approach of [25] but are careful to distinguish eigenvalues on the interior $\mathcal{D}$ from those on the boundary $\partial\mathcal{D}$.
Proof of Proposition 2.
($\Leftarrow$) Suppose that $M_{\mathcal{D}}(A,P)\succeq 0$ for some $P\succ 0$ and let $\lambda\in\lambda(A)$. Then there exists a nonzero $v\in\mathbb{C}^{n}$ for which $v^{\mathrm{H}}A=\lambda v^{\mathrm{H}}$. Consider the identity
$(I_{m}\otimes v)^{\mathrm{H}}M_{\mathcal{D}}(A,P)(I_{m}\otimes v)$ $=M_{0}\otimes v^{\mathrm{H}}Pv+M_{1}\otimes(v^{\mathrm{H}}APv)+M_{1}^{\top}\otimes(v^{\mathrm{H}}PA^{\top}v)$
$=M_{0}\otimes v^{\mathrm{H}}Pv+M_{1}\otimes(\overline{\lambda}v^{\mathrm{H}}Pv)+M_{1}^{\top}\otimes(\lambda v^{\mathrm{H}}Pv)$
$=v^{\mathrm{H}}Pv(M_{0}+M_{1}\lambda+M_{1}^{\top}\overline{\lambda})$
$=v^{\mathrm{H}}Pvf_{\mathcal{D}}(\lambda).$
The assumption $P\succ 0$ implies $v^{\mathrm{H}}Pv>0$, and $M_{\mathcal{D}}(A,P)\succeq 0$ further implies $f_{\mathcal{D}}(\lambda)\succeq 0$. Therefore $\lambda\in\textnormal{cl}(\mathcal{D})$.
Next suppose $\lambda\in\lambda(A)$ is non-simple and $\lambda\in\partial\mathcal{D}$. Then there exists nonzero $v_{1},v_{2}\in\mathbb{C}^{n}$ (linearly independent) such that $v^{\mathrm{H}}f_{\mathcal{D}}(\lambda)v=0$, $v_{1}^{\mathrm{H}}A=\lambda v_{1}^{\mathrm{H}}$, and $v_{2}^{\mathrm{H}}A=\lambda v_{2}^{\mathrm{H}}+v_{1}$. Because $\mathcal{D}$ is open, $\lambda\in\partial\mathcal{D}=\textnormal{cl}(\mathcal{D})\setminus\mathcal{D}$ must satisfy both $f_{\mathcal{D}}(\lambda)\succeq 0$ and $f_{\mathcal{D}}(\lambda)\not\succ 0$. Therefore $f_{\mathcal{D}}(\lambda)$ is singular, and there exists a nonzero vector $v\in\mathbb{C}^{m}$ such that $v^{\mathrm{H}}f_{\mathcal{D}}(\lambda)v=0$. With the $2\times 2$ matrices
$\tilde{P}$ $=\begin{bmatrix}p_{11}&p_{12}\\ p_{12}&p_{22}\end{bmatrix}\coloneqq\begin{bmatrix}v_{1}^{\mathrm{H}}\\ v_{2}^{\mathrm{H}}\end{bmatrix}P\begin{bmatrix}v_{1}&v_{2}\end{bmatrix}\succ 0$
$\tilde{J}$ $\coloneqq\lambda I_{2}+\begin{bmatrix}0&1\\ 0&0\end{bmatrix}$
we have $\begin{bmatrix}v_{1}&v_{2}\end{bmatrix}^{\mathrm{H}}A=\tilde{J}\begin{bmatrix}v_{1}&v_{2}\end{bmatrix}^{\mathrm{H}}$ and therefore
[…]
$=M_{\mathcal{D}}(\tilde{J},\tilde{P})\succeq 0.$
Next, we have
$\tilde{M}$ $\coloneqq K_{2,m}M_{\mathcal{D}}(\tilde{J},\tilde{P})K_{2,m}^{\top}$
$=\tilde{P}\otimes M_{0}+\tilde{J}\tilde{P}\otimes M_{1}+(\overline{\tilde{J}}\tilde{P})^{\top}\otimes M_{1}^{\top}$
$=\tilde{P}\otimes f_{\mathcal{D}}(\lambda)+\begin{bmatrix}p_{12}(M_{1}+M_{1}^{\top})&p_{22}M_{1}\\ p_{22}M_{1}^{\top}&0\end{bmatrix}\succeq 0.$
Finally,
$(I_{2}\otimes v)^{\mathrm{H}}\tilde{M}(I_{2}\otimes v)=\begin{bmatrix}p_{12}v^{\mathrm{H}}(M_{1}+M_{1}^{\top})v&p_{22}v^{\mathrm{H}}M_{1}v\\ p_{22}v^{\mathrm{H}}M_{1}^{\top}v&0\end{bmatrix}\succeq 0.$
But $\tilde{P}\succ 0$ implies $p_{22}>0$, so the above matrix inequality implies $v^{\mathrm{H}}M_{1}v=0$. Moreover, with $v^{\mathrm{H}}f_{\mathcal{D}}(\lambda)v=0$, we also have $v^{\mathrm{H}}M_{0}v=0$ and therefore $f(z)\equiv 0$ and $\mathcal{D}$ is empty, a contradiction. Therefore each $\lambda\in\lambda(A)$ non-simple implies $\lambda\in\mathcal{D}$.
($\Rightarrow$) Suppose $\lambda(A)\subset\textnormal{cl}(\mathcal{D})$ and $\lambda\in\lambda(A)$ non-simple implies $\lambda\in\mathcal{D}$.
If $A=\lambda$ is a (possibly complex) scalar, then it lies in $\textnormal{cl}(\mathcal{D})$ by assumption, with $M_{\mathcal{D}}(\lambda,p)=pf_{\mathcal{D}}(\lambda)\succeq 0$ for all $p>0$.
If $A=\lambda I_{n}+N$ is a (possibly complex) Jordan block, where $N\in\real^{n\times n}$ is a shift matrix and $n>1$, then $\lambda\in\mathcal{D}$ and $f_{\mathcal{D}}(\lambda)\succ 0$. Let $T_{k}\coloneqq\textnormal{diag}(k^{n-1},\ldots,k,1)$ for each $k\in\mathbb{I}_{>0}$. Then $T_{k}^{-1}AT_{k}=\lambda I_{n}+k^{-1}N\rightarrow\lambda I_{n}$ as $k\rightarrow\infty$. Moreover, because $M_{\mathcal{D}}$ is continuous, we have
$M_{\mathcal{D}}(T_{k}^{-1}AT_{k},I_{n})\rightarrow M_{\mathcal{D}}(\lambda I_{n},I_{n})=f_{\mathcal{D}}(\lambda)\otimes I_{n}\succ 0.$
Therefore there exists some $k_{0}\in\mathbb{I}_{>0}$ such that $M_{\mathcal{D}}(T_{k}^{-1}AT_{k},I_{n})\succ 0$ for all $k\geq k_{0}$. With $P\coloneqq T_{k}T_{k}^{\top}$, we have
$M_{\mathcal{D}}(A,P)$ $=M_{0}\otimes T_{k}T_{k}^{\top}+M_{1}\otimes(AT_{k}T_{k}^{\top})+M_{1}^{\top}\otimes(AT_{k}T_{k}^{\top})^{\top}$
$=(I_{m}\otimes T_{k})(M_{0}\otimes I_{n}+M_{1}\otimes T_{k}^{-1}AT_{k}+M_{1}^{\top}\otimes(T_{k}^{-1}AT_{k})^{\top})(I_{m}\otimes T_{k})^{\top}$
$=(I_{m}\otimes T_{k})M_{\mathcal{D}}(T_{k}^{-1}AT_{k},I_{n})(I_{m}\otimes T_{k})^{\top}\succ 0.$
Finally, for any $A\in\real^{n\times n}$, let $A=V(\bigoplus_{i=1}^{p}J_{i})V^{-1}$ denote the Jordan decomposition of $A$, where $J_{i}=\lambda_{i}I_{n_{i}}+N_{i}$, $\lambda_{i}\in\lambda(A)$, $N_{i}$ are shift matrices, and $n=\sum_{i=1}^{p}n_{i}$. We have already shown that for each $i\in\mathbb{I}_{1:p}$, there exists $P_{i}\succ 0$ such that $M_{\mathcal{D}}(J_{i},P_{i})\succeq 0$. Then with $\tilde{P}\coloneqq V(\bigoplus_{i=1}^{p}P_{i})V^{-1}$, we have
$(I_{m}$ $\otimes V^{-1})M_{\mathcal{D}}(A,\tilde{P})(I_{m}\otimes V^{-1})^{\mathrm{H}}$
$=M_{0}\otimes\left(\bigoplus_{i=1}^{p}P_{i}\right)+M_{1}\otimes\left(\bigoplus_{i=1}^{p}J_{i}P_{i}\right)+M_{1}\otimes\left(\bigoplus_{i=1}^{p}J_{i}P_{i}\right)^{\top}$
$=K_{n,m}\left(\bigoplus_{i=1}^{p}K_{m,n_{i}}M_{\mathcal{D}}(J_{i},P_{i})K_{m,n_{i}}^{\top}\right)K_{n,m}^{\top}\succeq 0$
[…]
Appendix C Proof of Proposition 4
To show Proposition 4(a), we first require the following eigenvalue sensitivity result due to [40, Thm. 7.2.3].
Theorem 6 ([40, Thm. 7.2.3]).
For any $A\in\mathbb{C}^{n\times n}$, denote its Schur decomposition by $A=Q(D+N)Q^{\mathrm{H}}$, where $Q\in\mathbb{C}^{n\times n}$ is unitary, $D\in\mathbb{C}^{n\times n}$ is diagonal, and $N\in\mathbb{C}^{n\times n}$ is strictly upper triangular. [^13] A matrix $U$ is strictly upper triangular if $U_{ij}=0$ for all $i\geq j$. Let $p$ be the smallest positive integer for which $M^{p}=0$ where $M_{ij}\coloneqq|N_{ij}|$. Then, for any $E\in\real^{n\times n}$ and $\mu\in\lambda(A+E)$,
$\min_{\lambda\in\lambda(A)}|\mu-\lambda|\leq\max\set{c}{E\|,(c\|E\|)^{1/p}}$
where $c\coloneqq\sum_{k=0}^{p-1}\|N\|^{k}$.
Proof of Proposition 4.
Throughout this proof, we show a set $S$ is not open (or not closed) by demonstrating that $S^{c}$ (or $S$) does not contain all its limit points.
(a)—For any $A\in\mathbb{A}_{\mathcal{D}}^{n}$, continuity of $f_{\mathcal{D}}$ gives the existence of a function $\delta(\lambda)>0$ such that $f_{\mathcal{D}}(z)\succ 0$ for all $|z-\lambda|<\delta(\lambda)$ and $\lambda\in\lambda(A)$. Let $\delta\coloneqq\min_{\lambda\in\lambda(A)}\delta(\lambda)$. By Theorem 6 and norm equivalence, there exist $c>0$ and $p\in\mathbb{I}_{1:n}$ such that
$\max_{\mu\in\lambda(A+E)}\min_{\lambda\in\lambda(A)}|\lambda-\mu|\leq\max\set{c}{E\|_{\textnormal{F}},(c\|E\|_{\textnormal{F}})^{1/p}}$
for all $E\in\real^{n\times n}$. Therefore there exists a $\varepsilon>0$ such that
$\max_{\mu\in\lambda(A+E)}\min_{\lambda\in\lambda(A)}|\lambda-\mu|<\delta$
[…]
(b)(i)—Because $\mathcal{D}$ is open, nonempty, and not equal to $\mathcal{D}$, $\partial\mathcal{D}$ is nonempty. Let $\lambda\in\partial\mathcal{D}$ and $\lambda_{k}\in\mathcal{D}^{c}$ be a sequence for which $\lambda_{k}\rightarrow\lambda$. By symmetry, we also have $\overline{\lambda}\in\mathcal{D}$ and $\overline{\lambda}_{k}\in\mathcal{D}^{c}$.
For $n=2$, we have $A\coloneqq\begin{bmatrix}\textnormal{Re}(\lambda)&-\textnormal{Im}(\lambda)\\ \textnormal{Im}(\lambda)&\textnormal{Re}(\lambda)\end{bmatrix}\in\real^{2\times 2}$ has eigenvalues $\lambda,\overline{\lambda}\in\mathcal{D}$, and $A_{k}\coloneqq\begin{bmatrix}\textnormal{Re}(\lambda_{k})&-\textnormal{Im}(\lambda_{k})\\ \textnormal{Im}(\lambda_{k})&\textnormal{Re}(\lambda_{k})\end{bmatrix}\in\real^{2\times 2}$ has eigenvalues $\lambda_{k},\overline{\lambda_{k}}\in\mathcal{D}^{c}$ for each $k\in\mathbb{I}_{>0}$. The corresponding eigenvectors are $\begin{bmatrix}\pm\iota\\ 1\end{bmatrix}\in\mathbb{C}^{2}$. Therefore $A\in\tilde{\mathbb{A}}_{\mathcal{D}}^{2}$ but $A_{k}\in(\tilde{\mathbb{A}}_{\mathcal{D}}^{2})^{c}$ for each $k\in\mathbb{I}_{>0}$, and the limit $A_{k}\rightarrow A$ gives us that $(\tilde{\mathbb{A}}_{\mathcal{D}}^{2})^{c}$ does not contain all its limit points.
For $n>2$, let $A_{0}\in\tilde{\mathbb{A}}_{\mathcal{D}}^{n-2}$, and we can extend the prior argument with the sequence $B_{k}\coloneqq A_{k}\oplus A_{0}\in(\tilde{\mathbb{A}}_{\mathcal{D}}^{n})^{c},k\in\mathbb{I}_{>0}$ that converges to $B\coloneqq A\oplus A_{0}\in\tilde{\mathbb{A}}_{\mathcal{D}}^{n}$.
(b)(ii)—By part (b)(i), it suffices to consider the case $n=1$. By closure and convexity of $\mathcal{D}$, $\mathcal{D}\cap\real$ is either a closed line segment, a closed ray, or itself. In other words, $\mathcal{D}\cap\real$ is open if and only if it has no endpoints. Moreover, since $\partial\mathcal{D}\cap\real$ is the set of the endpoints of $\mathcal{D}\cap\real$, $\mathcal{D}\cap\real$ is open if and only if $\partial\mathcal{D}\cap\real$ is empty. Finally, since $\tilde{\mathbb{A}}_{\mathcal{D}}^{1}=\mathcal{D}\cap\real$, $\tilde{\mathbb{A}}_{\mathcal{D}}^{1}$ is open if and only if $\partial\mathcal{D}\cap\real$ is empty.
(c)(i)—Let $\lambda\in\partial\mathcal{D}$. Suppose $n=4$. Then $\overline{\lambda}\in\partial\mathcal{D}$ by symmetry. Because $\mathcal{D}$ is open, there exists a sequence $\lambda_{k}\in\mathcal{D}$ such that $\lambda_{k}\rightarrow\lambda$, and by symmetry, we also have $\overline{\lambda}_{k}\in\mathcal{D}$ and $\overline{\lambda}_{k}\rightarrow\overline{\lambda}$. Consider again the $2\times 2$ matrices $A$ and $A_{k}$ from part (b)(i), which have eigenvalues $\lambda,\overline{\lambda}\in\mathcal{D}$ and $\lambda_{k},\overline{\lambda_{k}}\in\mathcal{D}^{c}$, respectively. Then the block matrices $B\coloneqq\begin{bmatrix}A&I_{2}\\ 0&A\end{bmatrix}\in\real^{4\times 4}$ and $B_{k}\coloneqq\begin{bmatrix}A_{k}&I_{2}\\ 0&A_{k}\end{bmatrix}\in\real^{4\times 4}$ have the same eigenvalues, but this time the eigenvectors are $\begin{bmatrix}\pm\iota\\ 1\\ 0\\ 0\end{bmatrix},\begin{bmatrix}0\\ 0\\ \pm\iota\\ 1\end{bmatrix}\in\mathbb{C}^{4}$ and the eigenvalues are non-simple. Since $\lambda$ is a non-simple eigenvalue on the boundary of $\mathcal{D}$, we have $B\not\in\tilde{\mathbb{A}}_{\mathcal{D}}^{4}$. However, $\lambda_{k}$ are all in the interior of $\mathcal{D}$, so $B_{k}\in\tilde{\mathbb{A}}_{\mathcal{D}}^{4}$. Since $B_{k}\rightarrow B$, the set $\tilde{\mathbb{A}}_{\mathcal{D}}^{4}$ does not contain all its limit points.
On the other hand, let $\lambda\in\partial\mathcal{D}$ and suppose $n>4$. Similarly to part (b)(i), with any $\tilde{A}_{0}\in\tilde{\mathbb{A}}_{\mathcal{D}}^{n-4}$, we can extend the argument for the $n=4$ case with the sequence $\tilde{A}_{k}\coloneqq B_{k}\oplus\tilde{A}_{0}\in\tilde{\mathbb{A}}_{\mathcal{D}}^{n},k\in\mathbb{I}_{>0}$ that converges to $\tilde{A}\coloneqq B\oplus\tilde{A}_{0}\in(\tilde{\mathbb{A}}_{\mathcal{D}}^{n})^{c}$.
(c)(ii)—Let $\lambda\in\partial\mathcal{D}\cap\real$ and $n\geq 2$. Because $\mathcal{D}$ is convex, open, and nonempty, there exists $\varepsilon>0$ such that exactly one of the real intervals $(\lambda,\lambda+\varepsilon)$ or $(\lambda-\varepsilon,\lambda)$ is contained in $\mathcal{D}$, whereas the other is contained in $\textnormal{int}(\mathcal{D}^{c})$. Without loss of generality, assume $(\lambda-\varepsilon,\lambda)\subseteq\mathcal{D}$. [^14] Otherwise, take the reflection about the imaginary axis $\mathcal{D}$ and $\tilde{\mathbb{A}}_{\mathcal{D}}^{n}$. Then $A_{k}\coloneqq(\lambda-\varepsilon/k)I_{n}+N_{n}\in\tilde{\mathbb{A}}_{\mathcal{D}}^{n}$ for each $k\in\mathbb{I}_{>0}$, but $A_{k}\rightarrow\lambda I_{n}+N_{n}\in(\tilde{\mathbb{A}}_{\mathcal{D}}^{n})^{c}$ and therefore $\tilde{\mathbb{A}}_{\mathcal{D}}^{n}$ does not contain all its limit points.
[…]
Appendix D Proof of Theorem 2
To prove Theorem 2(a,b), we use sensitivity results on the value functions of parameterized nonlinear SDPs,
$V(y)\coloneqq\inf_{x\in\mathbb{X}(y)}F(x,y)$ (46)
where the set-valued function $\mathbb{X}:\real^{m}\rightarrow\mathcal{P}(\real^{n})$ is defined by
[…]
Consider also the graph of the set-valued function $\mathbb{X}$,
[…]
Notice that $\mathbb{Z}$ is closed if $G$ is continuous. We say Slater’s condition holds at $y\in\real^{m}$ if there exists $x\in\real^{n}$ such that $x\in\textnormal{int}(\mathbb{X}(y))$, or equivalently, $G(x,y)\succ 0$. In the following proposition, we specialize [71, Prop. 4.4] to nonlinear SDPs.
Proposition 6 ([71, Prop. 4.4]).
Let $y_{0}\in\real^{m}$ and suppose
  1. (i)
    $F$ and $G$ are continuous on n+m;
  2. (ii)
    there exist $\alpha\in\real$ and compact $C\subset\real^{n}$ such that, for each $y$ in a neighborhood of $y_{0}$, the level set
[…]
    is nonempty and contained in $C$; and
  3. (iii)
    Slater’s condition holds at $y_{0}$.
Then $F(\cdot,y)$ attains a minimum on $\mathbb{X}(y)$ for all $y\in N_{y}$, and $V(y)$ is continuous at $y=y_{0}$.
Proof.
See [71, Prop. 4.4] and the discussions in [71, pp. 264, 483–484, 491–492]. ∎
Proof of Theorem 2.
Let $\textnormal{vec}:\real^{n\times n}\rightarrow\real^{n^{2}}$ and $\textnormal{vecs}:\real^{n\times n}\rightarrow\real^{(1/2)(n+1)n}$ denote the vectorization and symmetric vectorization operators, respectively.
(a)—With $x\coloneqq\textnormal{vecs}(P)$, $y\coloneqq\textnormal{vec}(A)$, $F(x,y)\coloneqq\textnormal{tr}(VP)$, and $G(x,y)\coloneqq P\oplus(M_{\mathcal{D}}(A,P)-M)$, we can use Proposition 6 to show the continuity of $\phi_{\mathcal{D}}$ on $\mathbb{A}_{\mathcal{D}}^{n}$. Let $A_{0}\in\mathbb{A}_{\mathcal{D}}^{n}$. Condition (i) of Proposition 6 holds by assumption. Slater’s condition (iii) holds because for any $P\succ 0$ such that $M_{\mathcal{D}}(A_{0},P)\succ 0$, we can define $P_{0}\coloneqq\gamma P\succ 0$ for some $\gamma>\gamma_{0}\coloneqq\|M\|\times\|[M_{\mathcal{D}}(A_{0},P)]^{-1}\|$ to give
$M_{\mathcal{D}}(A_{0},P_{0})=\gamma M_{\mathcal{D}}(A_{0},P)\succ\gamma_{0}M_{\mathcal{D}}(A_{0},P)\succeq M.$
Moreover, by continuity of $M_{\mathcal{D}}$, there exists a neighborhood $N_{A}$ of $A_{0}$ such that $M_{\mathcal{D}}(A,P_{0})\succ M$ for all $A\in N_{A}$. Letting $\alpha\coloneqq\textnormal{tr}(VP_{0})>0$, we have that the set
[…]
is compact and contains the nonempty level set
[…]
for all $A\in N_{A}$. Taking the image of each of the above sets under the vecs operation gives condition (ii) of Proposition 6. All the conditions of Proposition 6 are thus satisfied for each $A_{0}\in\mathbb{A}_{\mathcal{D}}^{n}$, and we have $\phi_{\mathcal{D}}$ is continuous on $\mathbb{A}_{\mathcal{D}}^{n}$.
(b)—Continuity of $\phi_{\mathcal{D}}$ on $\mathbb{A}_{\mathcal{D}}^{n}$ implies closure of the sublevel sets of $\phi_{\mathcal{D}}$, and eq. 33 follows by definition of $\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon)$.
(c)—First, $M_{\mathcal{D}}(A,P)\succ 0$ implies $P\succ 0$ by Proposition 3. Moreover, for any $P\succ 0$ such that $M_{\mathcal{D}}(A,P)\succ 0$, we have $M_{\mathcal{D}}(A,P)\succeq\gamma M_{\mathcal{D}}(A,P)\succeq M$ with $P\coloneqq\gamma P$ and $\gamma\coloneqq\|M\|\times\|[M_{\mathcal{D}}(A,P)]^{-1}\|$, so feasibility of eq. 17 is equivalent to feasibility of
$M_{\mathcal{D}}(A,P)$ $\succ M,$ $P$ $\succeq 0$
and therefore $\bigcup_{\varepsilon>0}\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon)=\mathbb{A}_{\mathcal{D}}^{n}$. But $\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon)$ is monotonically decreasing, [^15] By “monotonically decreasing” we mean $\varepsilon\leq\varepsilon^{\prime}\Rightarrow\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon)\supseteq\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon^{\prime})$. so $\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon)\nearrow\bigcup_{\varepsilon>0}\mathbb{A}_{\mathcal{D}}^{n}(\varepsilon)=\mathbb{A}_{\mathcal{D}}^{n}$ as $\varepsilon\searrow 0$. ∎
Appendix E Proof of Theorems 4 and 5
Starting with Theorem 4:
Proof of Theorem 4.
Since $\mu_{\varepsilon}$ is nondecreasing and bounded from below by $\mu$, it suffices to show that for each $\delta>0$, there exists a $\overline{\varepsilon}>0$ such that $\mu_{\overline{\varepsilon}}-\mu<\delta$.
Let $\theta^{*}\in\Theta$ denote a point for which $\mu=f(\theta^{*})$. If $\theta^{*}\in\Theta_{++}$, we could simply choose $\overline{\varepsilon}>0$ large enough to put $\theta^{*}$ in $\mathcal{T}(\Phi_{\overline{\varepsilon}})$ and achieve $\mu_{\overline{\varepsilon}}-\mu=0<\delta$.
Instead, we assume $\theta^{*}\not\in\Theta_{++}$. By 1, there exists a sequence $\theta_{k}\in\Theta_{++},k\in\mathbb{I}_{>0}$ such that $\theta_{k}\rightarrow\theta$ as $k\rightarrow\infty$. Defining $\nu_{k}\coloneqq f(\theta_{k})$, we have $\nu_{k}\rightarrow\mu$ by continuity of $f$. Therefore, there exists some $k_{0}\in\mathbb{I}_{>0}$ such that $\nu_{k}-\mu<\delta$ for all $k\geq k_{0}$. For each $\theta_{k}\in\Theta_{++}$, there exists a unique $\phi_{k}=(\beta_{k},L^{\mathcal{I}_{\Sigma}}_{k},L^{\mathcal{I}_{\mathcal{A}}}_{k})\in\Phi$ such that $\theta_{k}=\mathcal{T}(\phi_{k})$ (by Lemma 4). Let $\overline{\varepsilon}$ be the minimum over all the diagonal elements of $L^{\mathcal{I}_{\Sigma}}_{k_{0}}$ and $L^{\mathcal{I}_{\mathcal{A}}}_{k_{0}}$. Then $(\beta_{k_{0}},L^{\mathcal{I}_{\Sigma}}_{k_{0}},L^{\mathcal{I}_{\mathcal{A}}}_{k_{0}})\in\Phi_{\overline{\varepsilon}}$ by construction, $\nu_{k_{0}}\geq\mu_{\overline{\varepsilon}}$ by optimality, and $\mu_{\overline{\varepsilon}}-\mu\leq\nu_{k_{0}}-\mu<\delta$. ∎
As in Appendix D, we use sensitivity results of [71] on optimization problems to prove Theorem 5. This time, however, we consider the continuity of the value function for parameterized NLPs on Banach spaces. Let $\mathcal{X}$, $\mathcal{Y}$, and $\mathcal{K}$ be Banach spaces and consider the parameterized NLP,
$V(y)\coloneqq\inf_{x\in\mathbb{X}(y)}F(x,y)$ (47)
where the set-valued function $\mathbb{X}:\mathcal{Y}\rightarrow\mathcal{P}(\mathcal{X})$ is defined by
[…]
for some $G:\mathcal{X}\times\mathcal{Y}\rightarrow\mathcal{K}$ and $K\subseteq\mathcal{K}$ is closed. Let $X^{0}(y)$ denote the (possibly empty) set of solutions to eq. 47. Define the graph of the set-valued function $\mathbb{X}(\cdot)$ by
[…]
Notice that $\mathbb{Z}$ is closed if $G$ is continuous and $K$ is closed.
Proposition 7 ([71, Prop. 4.4]).
Let $y_{0}\in\mathcal{Y}$ and assume:
  1. (i)
    $F$ and $G$ are continuous on $\mathcal{X}\times\mathcal{Y}$ and $K$ is closed;
  2. (ii)
    there exist $\alpha\in\real$ and a compact set $C\subseteq\mathcal{X}$ such that, for every $y$ in a neighborhood of $y_{0}$, the level set
[…]
    is nonempty and contained in $C$; and
  3. (iii)
    for any neighborhood $N_{x}$ of the solution set $X^{0}(y_{0})$, there exists a neighborhood $N_{y}$ of $y_{0}$ such that $N_{x}\cap\mathbb{X}(y)$ is nonempty for all $y\in N_{y}$;
then $V(y)$ is continuous and $X^{0}(y)$ is outer semicontinuous at $y=y_{0}$.
Proof of Theorem 5.
First, we must specify $\overline{\varepsilon}$. For each $\theta\in\Theta_{++}$, let
[…]
where the maximum is achieved since there is a finite number of diagonal elements of the Cholesky factors that must be lower bounded. Now we specify $\overline{\varepsilon}$ as the supremum of $\varepsilon(\theta)$ over all $\theta\in\Theta_{f\leq\alpha}\cap\Theta_{++}$,
[…]
so that, for any $\varepsilon\in(0,\overline{\varepsilon})$, $\Theta_{f\leq\alpha}\cap\mathcal{T}(\Phi_{\varepsilon})$ is nonempty and is contained in the compact set $C$.
(a)—Following the proof of [71, Prop. 4.4], we have (i) $F$ is continuous and (ii) the level set $\Theta_{f\leq\alpha}$ is nonempty and contained in the compact set $C$, which implies $\Theta_{f\leq\alpha}$ is a compact level set and therefore the minimum of $f$ over $\Theta_{f\leq\alpha}$ is achieved and equals the minimum over $\Theta$. Moreover, $\hat{\theta}_{0}$ must be nonempty.
(b)—Similarly to part (a), we have, for each $\varepsilon\in(0,\overline{\varepsilon})$, that the level set $\Theta_{f\leq\alpha}\cap\mathcal{T}(\Phi_{\varepsilon})$ is nonempty and contained in the compact set $C$, so $f$ achieves its minimum over $\mathcal{T}(\Phi_{\varepsilon})$ and $\hat{\theta}_{\varepsilon}$ is nonempty.
(c)—Consider the graph of the constraint function,
[…]
Consider a sequence $(\theta_{k},\varepsilon_{k})\in\mathbb{Z},k\in\mathbb{I}_{>0}$ that is convergent $(\theta_{k},\varepsilon_{k})\rightarrow(\theta,\varepsilon)$. Then $\varepsilon\geq 0$, otherwise the sequence would not converge. Moreover, $\theta\in\Theta$ since $\theta_{k}\in\mathcal{T}(\Phi_{\varepsilon_{k}})\subseteq\Theta$ for all $k\in\mathbb{I}_{>0}$ and $\Theta$ contains all its limit points. If $\varepsilon=0$, then $(\theta,\varepsilon)\in\mathbb{Z}$ trivially. On the other hand, if $\varepsilon>0$, then $\varepsilon(\theta_{k})$ converges to $\varepsilon(\theta)$ because $\mathcal{T}$ is continuous and the max can be taken over a finite number of elements of $\mathcal{T}^{-1}(\theta_{k})$. Moreover, $\varepsilon(\theta_{k})$ upper bounds $\varepsilon_{k}$ because $\theta_{k}\in\mathcal{T}(\Phi_{\varepsilon_{k}})$, so $\varepsilon(\theta)\geq\varepsilon$. Finally, we have $\theta\in\mathcal{T}(\Phi_{\varepsilon})$, $(\theta,\varepsilon)\in\mathbb{Z}$, and $\mathbb{Z}$ is closed.
Let $\varepsilon_{0}\geq 0$ and $N_{\theta}$ be a neighborhood of $\hat{\theta}_{\varepsilon_{0}}$. With
[…]
we have $N_{\theta}\cap\Theta$ and $N_{\theta}\cap\mathcal{T}(\Phi_{\varepsilon})$ are nonempty for all $\varepsilon\in(0,\varepsilon_{0}+\delta)$.
Finally, the requirements of Proposition 7 are satisfied for all $\varepsilon_{0}\in[0,\overline{\varepsilon})$, so $\mu_{\varepsilon}$ is continuous and $\hat{\theta}_{\varepsilon}$ is outer semicontinuous at $\varepsilon=\varepsilon_{0}$.
(d)—The last statement follows by the definition of outer semicontinuity and the fact that the $\limsup$ is nonempty. ∎